### v2

# Some Characterizations of TTC in Multiple-Object Reallocation Problems

Abstract.
This paper considers exchange of indivisible objects when agents are endowed with and desire bundles of objects. Agents are assumed to have lexicographic preferences over bundles. We show that the generalized Top Trading Cycles rule (TTC) is characterized by Pareto efficiency, balancedness, the weak endowment lower bound, and truncation-proofness (or drop strategy-proofness). In the classic Shapley-Scarf model, TTC is characterized by Pareto efficiency, individual rationality, and truncation-proofness. The proof is nonstandard and its novelty has independent methodological interest.
Keywords: exchange of indivisible objects; Top Trading Cycles; heuristic manipulation; endowment lower bound.
conference: ; ;
1. Introduction
In this paper, we consider reallocation problems when agents have multi-unit demands. [^1] One could also consider this problem as a barter economy with indivisibility. Here, we use the term problem to emphasize that our approach is in a centralized manner, rather than a decentralized exchange. There is a group of agents, each of whom is initially endowed with some set of heterogeneous and indivisible objects. Each agent has preferences over bundles of objects. A social planner can select an allocation rule to redistribute the objects among the agents in order to make agents better off, but monetary transfers are forbidden.
There are many examples of such problems. Employees at a firm may exchange tasks, shifts, or equipment (Yu and Zhang, 2020). University students exchange course schedules, while universities exchange students for semester-long exchange programs (Bloch et al., 2020). Teams in professional sports leagues trade players’ contract rights (Coreno and Balbuzanov, 2022). [^2] Of course, money is a feature of professional sports leagues. However, the presence of salary caps means that prices are not the main rationing instrument. Collectors trade collectibles, such as Pokémon cards (Caspari, 2020). The classic Shapley-Scarf model, in which each agent is endowed with a single object (called a house), is an important special case (Shapley and Scarf, 1974). [^3] Throughout this paper, we shall distinguish between single-object problems, in which agents are endowed with a single object, and multiple-object problems, in which agents are endowed with bundles.
Like other studies focusing on multiplicity, e.g., Altuntaş et al. (2023) and Biró and Csáji (2024), in this paper we assume that agents have lexicographic preferences over bundles of objects. That is, each agent has strict preferences over objects, and these preferences are extended to bundles as follows: for any two distinct bundles, $X$ and $Y$, bundle $X$ is preferred to $Y$ if and only if the best object that belongs to exactly one of the two bundles is in $X$. There are several examples of allocation problems with lexicographic preferences. For instance, students at a university may want to reschedule their courses by exchanging their assigned slots, and students may lexicographically prefer certain timings, e.g., a “party boy” prefers not to attend the class on Friday. Also, many single-object problems can be reinterpreted as a multiple-object problem by identifying each object with its set of attributes. For example, in kidney-exchange problems, each kidney is identified by its characteristics (e.g., compatibility, age of donor, etc.), which can be ranked lexicographically.
Aside from its suitability for applications, our focus on lexicographic preferences is also helpful from an implementation perspective: since each lexicographic preference relation defined on the set of bundles is completely specified by a strict preference relation on the set of single objects, the social planner only needs to elicit a minimal amount of information about the agents’ preferences. This is particularly useful if the set of objects is large, so that agents could not feasibly rank all bundles. Moreover, from a theoretical point of view, the lexicographic domain admits several positive results where larger preference domains yield impossibilities. For instance, the core can be empty under additive preferences (Konishi et al., 2001), while it is nonempty under (conditionally) lexicographic preferences (Sikdar et al., 2017; Fujita et al., 2018).
1.1. Overview of the paper
We are interested in allocation rules that satisfy appealing properties, such as efficiency and strategic robustness. Standard properties of efficiency and strategic robustness are Pareto efficiency and strategy-proofness, respectively. However, when agents consume more than one object, the combination of Pareto efficiency and strategy-proofness essentially results in serially dictatorial rules (Pápai, 2001; Klaus and Miyagawa, 2002; Monte and Tumennasan, 2015; Todo et al., 2014). Since such rules ignore property rights driven by endowments, agents who participate in those mechanisms may be even worse off after the reallocation. Thus, such rules may not be desirable, and this impossibility leads to a common trade-off between efficiency and strategic robustness. Based on this trade-off, in this paper we search for a plausible rule by relaxing strategy-proofness while keeping Pareto efficiency. As in Biró et al. (2022a) and Altuntaş et al. (2023), [^4] Biró et al.’s (2022a) model is slightly different to ours. In their model, each agent is endowed with a set of agent-specific homogeneous objects. since agents may only consider simple manipulation in practice, we consider two heuristics for misreporting preferences: drop strategies and truncation strategies. Loosely speaking, starting from her true preference relation over single objects, an agent may strategically misreport a new preference by dropping one object to the bottom, or by truncating the set of “acceptable objects.” [^5] Though preferences are lexicographic and, hence, monotone, we interpret an agent’s set of “acceptable objects” as the set of objects that are weakly preferred to some object in her endowment. A rule is drop strategy-proof if it is immune to manipulation via drop strategies, and is truncation-proof if it is immune to manipulation via truncation strategies. We believe that truncation strategies are perhaps the most obvious forms of manipulation, as they are “monotonic manipulations” in the sense of Troyan and Morrill (2020).
In addition to efficiency and strategic robustness, we also consider two properties that guarantee a lower bound on agents’ welfare, ensuring that they will not be worse off after the reallocation. The first one is balancedness, which says that a rule should assign the same number of objects to each agent as the initial endowment. The second one, which is our main property in this paper, is the weak endowment lower bound. It requires that no agent ever receives an object which is worse than all objects in her endowment. [^6] Here, by “weak” we mean that it is weaker than Altuntaş et al.’s (2023) strong endowment lower bound. However, our weak endowment lower bound is independent of individual rationality in multiple-object problems, whereas it coincides with individual rationality in single-object problems.
Though it is independent of individual rationality in multiple-object problems, there are practical reasons for considering the weak endowment lower bound as a desideratum. First, the weak endowment lower bound allows agents to veto some of the other agents’ objects, and the right to veto can be viewed as a minimal requirement. [^7] There are some settings in which agents should not expect reallocation to be Pareto-improving. For example, in professional sports leagues, some players at team $i$ may express their desire to be traded to another team. The resulting reallocation need not benefit team $i$, yet the team should not be forced to sign “unacceptable players.” Second, given the complexity of eliciting preferences over bundles, a social planner interested in individual rationality may seek to simplify the problem by focusing on only a few salient features. It turns out that, in the presence of other desirable properties, namely balancedness and truncation-proofness, satisfying the weak endowment lower bound is sufficient to guarantee full individual rationality (see Proposition 2.11).
Our main result is that the generalized Top Trading Cycles rule (TTC) [^8] For convenience, we still call our rule TTC. is characterized by Pareto efficiency, balancedness, the weak endowment lower bound, and truncation-proofness (or drop strategy-proofness) (Theorem 3.2). Our result supports the usefulness of the TTC algorithm in this more general and more complex environment. Since the combination of balancedness and the weak endowment lower bound is weaker than Altuntaş et al.’s (2023) strong endowment lower bound, their main characterization is a corollary of ours (Corollary 3.3). Moreover, we obtain a similar characterization for the Shapley-Scarf model, which is of independent interest. Specifically, TTC is characterized by Pareto efficiency, individual rationality, and truncation-proofness (Theorem 3.4).
We note that, unlike Altuntaş et al.’s (2023) proof, the proof of our characterization is not completely constructive. We borrow the notions of size and similarity from Sethuraman (2016) and Ekici (2023), respectively, and we give a proof by minimum counter-example. To the best of our knowledge, we are the first to successfully extend Sethuraman (2016) and Ekici’s (2023) approach from single-object problems to multiple-object problems. We believe that in future studies, working with this new approach can be quite fruitful.
1.2. Related literature
The paper closest to ours is Altuntaş et al. (2023). As mentioned previously, their main characterization result is a corollary of ours. For a relative larger domain, monotonic and responsive preferences, Pápai (2003) characterizes a class of rules, the Segmented Trading Cycle rules by strategy-proofness, non-bossiness, trade sovereignty, and strong individual rationality. Manjunath and Westkamp (2021) show that for trichotomous preferences, Pareto efficiency, strategy-proofness, and individual rationality are compatible.
Some papers focus on models that are slightly different from ours, e.g., the multi-unit Shapley-Scarf model, where each agent is endowed with a set of agent-specific homogeneous objects (Biró et al., 2022b, a), and the multiple-type Shapley-Scarf model, in which objects are partitioned by different types, and each agent owns one object of each type (Feng et al., 2024; Feng, 2023).
Apart from the axiomatic approach, some papers consider the computational complexity of manipulation in TTC (Fujita et al., 2018; Phan and Purcell, 2022). Those studies emphasize that weakening strategy-proofness for our model does make sense as finding a beneficial manipulation may be computationally difficult.
In our model, one difficulty is the trade-off between Pareto efficiency and strategy-proofness. In this paper, we keep Pareto efficiency, and weaken strategy-proofness to drop strategy-proofness and truncation-proofness. There are also many other relaxed forms of strategy-proofness, e.g., rank monotonicity (Chen and Zhao, 2021), partial strategy-proofness (Mennle and Seuken, 2021), truncation-invariance (Chen et al., 2024), and weak truncation robustness (Hashimoto et al., 2014). On the other hand, several studies take the opposite approach by keeping strategy-proofness while relaxing Pareto efficiency, e.g., Pápai (2007), Anno and Kurino (2016), and Nesterov (2017).
1.3. Organization
The rest of our paper is organized as follows. In Section 2, we introduce our model, rules and their properties, and we give a description of generalized TTC. We state our result in Section 3: we show that for lexicographic preferences, TTC is characterized by Pareto efficiency, balancedness, the weak endowment lower bound, and truncation-proofness (or drop strategy-proofness). In Subsection 3.1, we also provide a related characterization in the Shapley–Scarf model, and give the detailed proof for this simpler result, to help the reader to get the intuition of our proof strategy of our main result. In subsection 3.2, we provide some possibility and impossibility results for two more general domains. In Section 4, we conclude with some final remarks. Appendix A contains all proofs that were omitted from the main text, and we demonstrate the independence of our properties in Appendix B.1. Several other important examples are relegated to Appendix B.2.
2. Preliminaries
2.1. Model
Let $N=\left\{1,2,\dots,n\right\}$ be a finite set of $n\geq 2$ agents. Let $O$ be a finite set of heterogeneous and indivisible objects A bundle is a subset of $O$. Let $2^{O}$ denote the set of bundles. A generic element of $O$ is denoted by a lowercase letter, e.g., $x,y$, and a generic element of $2^{O}$ is typically denoted by an uppercase letter, e.g., $X,Y$.
An allocation is a function $\mu:N\to 2^{O}$ such that (i) for all $i,j\in N$, $i\neq j$ implies $\mu\left(i\right)\cap\mu\left(j\right)=\emptyset$, and (ii) $\bigcup_{i\in N}\mu\left(i\right)=O$. We shall represent each allocation $\mu$ as a profile $\left(\mu_{i}\right)_{i\in N}$ of disjoint bundles with $\bigcup_{i\in N}\mu_{i}=O$. For each $i\in N$, $\mu_{i}$ denotes agent $i$’s assignment at $\mu$. Let ${\mathcal{M}}$ denote the set of allocations.
Let $\omega=\left(\omega_{i}\right)_{i\in N}\in\mathcal{M}$ denote the initial allocation. For each $i\in N$, bundle $\omega_{i}$ is agent $i$’s endowment. We assume that each $\omega_{i}$ is nonempty. With some abuse of notation, we denote the initial owner of object $o$ by $\omega^{-1}\left(o\right)$.
Each agent $i$ has a strict preference relation $P_{i}$ over the set of all bundles. If agent $i$ prefers bundle $X$ to bundle $Y$, then we write $X\mathbin{P_{i}}Y$. Let $R_{i}$ denote the at least as good as relation associated with $P_{i}$. Then $X\mathbin{P_{i}}Y$ means that ($X\mathbin{P_{i}}Y$ or $X=Y$). [^9] Formally, $R_{i}$ is a complete, transitive, and antisymmetric binary relation on $2^{O}$, and $P_{i}$ is the asymmetric part of $R_{i}$. Let $P=\left(P_{i}\right)_{i\in N}$ denote a preference profile, where $P_{i}$ is agent $i$’s preference relation.
When there is no danger of confusion, we abuse notation and denote singletons $\left\{x\right\}$ by $x$. For example, $xP_{i}y$ means $\left\{x\right\}P_{i}\left\{y\right\}$, and $X\cup x$ means $X\cup\{x\}$.
Given a preference relation $P_{i}$ on $2^{O}$ and a nonempty bundle $X\in 2^{O}$, let $\top_{P_{i}}\left(X\right)$ denote the top-ranked object in $X$, i.e., $\top_{P_{i}}\left(X\right)=x$, where $x\in X$ and $x\mathbin{R_{i}}y$ for all $y\in X$. Similarly, let $\bot_{P_{i}}\left(X\right)$ denote the bottom-ranked object in $X$, i.e., $\bot_{P_{i}}\left(X\right)=x$, where $x\in X$ and $y\mathbin{R_{i}}x$ for all $y\in X$.
Furthermore, we assume that agents’ preferences over $2^{O}$ are lexicographic. That is, when facing two distinct bundles $X$ and $Y$, if agent $i$ prefers the top-ranked object in $X$ to that in $Y$, then $i$ also prefers $X$ to $Y$. If these top-ranked objects are the same, then $i$ compares the second top-ranked object in each bundle, and so on. Finally, if $Y$ is a subset of $X$, then $i$ prefers $X$ to $Y$. Formally, for any two distinct bundles $X,Y\in 2^{O}$, $X\mathbin{P_{i}}Y$ if
(1) there is an object $\text{there is an object }o\in X\setminus Y\text{ such that }\top_{P_{i}}\left((Y\setminus X)\cup o\right)=o.$
Note that the set of lexicographic preference relations over $2^{O}$ is in bijection with the set of strict linear orders over $O$. Therefore, we will identify lexicographic preference relations with their associated strict linear orders over single objects. In particular, we frequently represent lexicographic preferences as ordered list of objects, e.g., $P_{i}:x_{1},x_{2},\dots,x_{m}$ means that $x_{1}\mathbin{P_{i}}x_{2}\mathbin{P_{i}}\cdots\mathbin{P_{i}}x_{m}$, with the understanding that all remaining relations among bundles in $2^{O}$ can be deduced from (1). Let ${\mathcal{P}}$ denote the set of all lexicographic preference relations on $2^{O}$ (or, equivalently, their associated strict linear orders on $O$). Let ${\mathcal{P}}^{N}$ denote the set of all preference profiles.
Given a preference profile $P=\left(P_{i}\right)_{i\in N}$ and a relation $P^{\prime}_{i}$ for agent $i$, $\left(P^{\prime}_{i},P_{-i}\right)$ denotes the preference profile in which agent $i$’s preference relation is $P^{\prime}_{i}$ and, for each agent $j\in N\backslash\left\{i\right\}$, agent $j$’s preference relation is $P_{j}$.
An object reallocation problem (a problem in short) is a triple $(N,\omega,P)$. Since $(N,\omega)$ remains fixed throughout, we will simply denote a problem by $P$. Thus, the preference profile domain ${\mathcal{P}}^{N}$ also denotes the set of all problems (with lexicographic preferences).
2.2. Rules and their properties
A rule is a function $\varphi:{\mathcal{P}}^{N}\to{\mathcal{M}}$ that associates to each profile $P$ an allocation $\varphi\left(P\right)$. For each $i\in N$, $\varphi_{i}\left(P\right)$ denotes agent $i$’s assignment at $\varphi\left(P\right)$.
We now introduce and discuss some well-known properties for allocations and rules.
First, we consider one well-known efficiency criterion. Let $P\in{\mathcal{P}}^{N}$. An allocation $\overline{\mu}$ Pareto-dominates another allocation $\mu$ at $P$ if (i) for all $i\in N$, $\overline{\mu}_{i}R_{i}\mu_{i}$, and (ii) for some $i\in N$, $\overline{\mu}_{i}P_{i}\mu_{i}$. An allocation $\mu$ is Pareto efficient at $P$ if it is not Pareto-dominated at $P$ by any other allocation.
Definition 2.1 (Pareto efficiency).
A rule $\varphi$ is Pareto efficient if, for each $P\in{\mathcal{P}}^{N}$, $\varphi\left(P\right)$ is Pareto efficient at $P$.
The next property, strategy-proofness, is one incentive property that is most frequently used in the literature on mechanism design. It says that no agent can benefit from misrepresenting her preferences.
Definition 2.2 (Strategy-proofness).
A rule $\varphi$ is strategy-proof if, for each $P\in{\mathcal{P}}^{N}$, each $i\in N$, and each $P^{\prime}_{i}\in\mathcal{P}$, $\varphi_{i}\left(P\right)\mathbin{R_{i}}\varphi_{i}\left(P^{\prime}_{i},P_{-i}\right)$. In other words, at any $P$, no agent $i$ can manipulate $\varphi$ by misreporting any preference relation.
A strategy-proof rule $\varphi$ prohibits each agent $i$ from manipulating the rule at any $P$ by misreporting any preference relation in ${\mathcal{P}}$. A natural weakening of strategy-proofness is to prohibit each agent $i$ from manipulating the rule at any $P$ by reporting any preference relation in some subset ${\mathcal{S}}\left(P_{i}\right)$ of ${\mathcal{P}}$. The restricted set of manipulation strategies, ${\mathcal{S}}\left(P_{i}\right)$, could consist of misrepresentations which are, in some sense, more robust, or whose consequences are simpler to understand. To this end, we define two manipulation heuristics, drop strategies and truncation strategies, and we define the associated incentive properties, namely drop strategy-proofness and truncation-proofness. Let $P_{i}\in\mathcal{P}$.
We say that $P^{\prime}_{i}\in\mathcal{P}$ is a drop strategy for $P_{i}$ if there exists $x\in O\backslash\omega_{i}$ such that:
  1. (1)
    $\bot_{P^{\prime}_{i}}\left(O\right)=x$; and
  2. (2)
    for any $y,z\in O\setminus x$, $y\mathbin{P^{\prime}_{i}}z$ if and only if $y\mathbin{P_{i}}z$.
In this case, we say that $P^{\prime}_{i}$ is obtained from $P_{i}$ by dropping object $x$. Let ${\mathcal{D}}\left(P_{i}\right)$ denote the set of all drop strategies for $P_{i}$.
Definition 2.3 (Drop strategy-proofness).
A rule $\varphi$ is drop strategy-proof if, for each $P\in{\mathcal{P}}^{N}$, each $i\in N$, and each $P^{\prime}_{i}\in{\mathcal{D}}\left(P_{i}\right)$, $\varphi_{i}\left(P\right)\mathbin{R_{i}}\varphi_{i}\left(P^{\prime}_{i},P_{-i}\right)$. In other words, at any $P$, no agent $i$ can manipulate $\varphi$ by any drop strategy. [^10] Our definition of drop strategy-proofness is precisely the weakening of drop strategy-proofness and subset total drop strategy-proofness discussed in Footnote 31 of Altuntaş et al. (2023).
A truncation strategy for agent $i$ involves dropping a subset $X\subseteq O\backslash\omega_{i}$ to the bottom of her preference list; we require that $X$ is a “tail” of her preference list, i.e., it consists of all objects in $O\backslash\omega_{i}$ that are weakly worse than some object $x\in O\backslash\omega_{i}$. To formalize the definition, we first introduce subset drop strategies.
We say that $P^{\prime}_{i}$ is a subset drop strategy for $P_{i}$ if there exists $X\subseteq O\backslash\omega_{i}$ such that:
  1. (1)
    for all $x\in X$ and $y\in O\backslash X$, $y\mathbin{P^{\prime}_{i}}x$; and
  2. (2)
    for all $x,y\in X$, $x\mathbin{P^{\prime}_{i}}y$ if and only if $x\mathbin{P_{i}}y$; and
  3. (3)
    for all $x,y\in O\backslash X$, $x\mathbin{P^{\prime}_{i}}y$ if and only if $x\mathbin{P_{i}}y$.
In this case, we say that $P^{\prime}_{i}$ is obtained from $P_{i}$ by dropping the subset $X$. We say that $P^{\prime}_{i}$ is a truncation strategy for $P_{i}$ if, in addition to (1)-(3), the following holds:
  1. (4)
    $X=O\backslash\omega_{i}$ or there exists $x\in O\backslash\omega_{i}$ such that $X=\{o\in O\backslash\omega_{i}\mid x\mathbin{P_{i}}o\}$.
If $X=\{o\in O\backslash\omega_{i}\mid x\mathbin{P_{i}}o\}$, then $P^{\prime}_{i}$ is called the truncation of $P_{i}$ at object $x$. An important distinction between truncation strategies and drop strategies is that, unlike every drop strategy for $P_{i}$, every truncation strategy $P^{\prime}_{i}$ of $P_{i}$ agrees with $P_{i}$ on $O\backslash\omega_{i}$, i.e., for all $x,y\in O\backslash\omega_{i}$, $x\mathbin{P_{i}}y$ if and only if $x\mathbin{P^{\prime}_{i}}y$.
Definition 2.4 (Truncation proofness).
A rule $\varphi$ is truncation proof if, for each $P\in{\mathcal{P}}^{N}$, each $i\in N$, and each $P^{\prime}_{i}\in{\mathcal{T}}\left(P_{i}\right)$, $\varphi_{i}\left(P\right)\mathbin{R_{i}}\varphi_{i}\left(P^{\prime}_{i},P_{-i}\right)$. In other words, at any $P$, no agent $i$ can manipulate $\varphi$ by any truncation strategy.
Remark 1.
Truncation strategies are well-studied in the literature on matching theory (Roth and Vande Vate, 1991; Roth and Rothblum, 1999; Ehlers, 2008; Castillo and Dianat, 2016). Intuitively, truncating means by shifting the outside option up, thereby truncating the “tail” of acceptable objects. This notion can be easily extended to the Shapley-Scarf model by replacing the outside option with the agent’s single endowment (see the definition of truncation strategy in Subsection 3.1). However, the difficulty arises when an agent has more than one endowed object, as now the definitions of acceptable objects and unacceptable objects become unclear. Of course, for an agent $i$, if an object is better than her best endowed object $\top_{P_{i}}\left(\omega_{i}\right)$, then it is unarguably acceptable; and if it is worse than her worst endowed object $\bot_{P_{i}}\left(\omega_{i}\right)$, then it is unacceptable. But what about objects that are ranked between these two extremes of endowed objects? Motivated by drop strategies in Kojima and Pathak (2009), [^11] Drop strategies defined in Kojima and Pathak (2009) are a generalization of truncation strategies, and Kojima and Pathak (2009) show that drop strategies are exhaustive when searching for manipulations for agents with a capacity greater than one. The same is not true in our model. here we consider an extreme way to model truncation strategies: truncating a preference relation $P_{i}$ at an object $x$ corresponds to vetoing all objects worse than $x$ (by declaring them worse than her worst endowed object $\bot_{P_{i}}(\omega_{i})$). While other definitions are conceivable, ours is the weakest that induces the characterization result. $\diamond$
Example 2.5 (Drop strategies and truncation strategies).
Suppose $\omega_{i}=\left\{o_{i}^{1},o_{i}^{2}\right\}$ and $P_{i}:a,b,o_{i}^{1},c,d,o_{i}^{2},e$. Then:
  • $P_{i}^{1}:b,o_{i}^{1},c,d,o_{i}^{2},e,a$ is a drop strategy for $P_{i}$; it is obtained by dropping object $a$. Note that $P^{1}_{i}$ is not a truncation strategy for $P_{i}$.
  • $P_{i}^{2}:a,b,o_{i}^{1},c,o_{i}^{2},d,e$ is a truncation strategy for $P_{i}$; it is obtained from $P_{i}$ by subsequently dropping $d$ and $e$.
  • $P_{i}^{3}:a,o_{i}^{1},o_{i}^{2},b,c,d,e,$ is a truncation strategy for $P_{i}$; it is obtained from $P_{i}$ by subsequently dropping $b,c,d$, and $e$. $\diamond$
Our next property requires that each agent is assigned exactly as many objects as they initially own.
Definition 2.6 (Balancedness).
A rule $\varphi$ is balanced if, for each $P\in{\mathcal{P}}^{N}$ and each $i\in N$, $\left|\varphi_{i}\left(P\right)\right|=\left|\omega_{i}\right|$.
Balancedness is a constraint in many practical reallocation problems, such as shift reallocation among workers and course reallocation among university students. In the absence of such constraints, balancedness can be interpreted as a mild endowment monotonicity property, or a weak form of fairness. It guarantees that agents are rewarded with more objects (but not necessarily better objects) whenever they bring more objects to the market.
We now introduce two related participation requirements. The first property, individual rationality, is well-studied; it guarantees that no agent is harmed by the reallocation.
Definition 2.7 (Individual rationality).
A rule $\varphi$ satisfies individual rationality if, for each $P\in{\mathcal{P}}^{N}$ and each $i\in N$, $\varphi_{i}\left(P\right)\mathbin{R_{i}}\omega_{i}$.
The second property, the strong endowment lower bound, was introduced by Altuntaş et al. (2023). It guarantees that each agent is assigned a bundle that “pairwise dominates” her initial endowment, in the following sense.
Definition 2.8 (Strong endowment lower bound).
A rule $\varphi$ satisfies the strong endowment lower bound if, for each $P\in{\mathcal{P}}^{N}$ and each $i\in N$, there is a bijection $\sigma:\omega_{i}\to\varphi_{i}\left(P\right)$ such that for each $o\in\omega_{i}$, $\sigma(o)\mathbin{R_{i}}o$. Note that $\{\sigma(o)\mid o\in\omega_{i}\}=\varphi_{i}\left(P\right)$.
The strong endowment lower bound ensures that, for each agent $i\in N$, each object in $x\in\omega_{i}$ is exchanged for something weakly better than $x$. Our main property, the weak endowment lower bound, relaxes this significantly. It requires that each agent is only assigned objects that are weakly better than her worst endowed object.
Let $P\in{\mathcal{P}}^{N}$. For each agent $i$ and each preference relation $P_{i}$, let $w_{i}(P_{i})\coloneqq\bot_{P_{i}}\left(\omega_{i}\right)$ be agent $i$’s worst endowed object at $P_{i}$. An allocation $\mu$ satisfies the weak endowment lower bound at $P$ if, for each agent $i\in N$, and each of her assigned objects $o\in\mu_{i}$, $o\mathbin{R_{i}}w_{i}(P_{i})$. [^12] Biró et al. (2022a) study a model related to ours, in which each agent is endowed with a set of homogeneous objects. They consider a domain of preferences such that each agent prefers her initial endowment to any bundle containing at least one unacceptable object, i.e., a single object worse than her own. In that setting, the weak endowment lower bound is implied by individual rationality. Our property captures a similar idea by restricting rules instead of preferences.
Definition 2.9 (Weak endowment lower bound).
A rule $\varphi$ satisfies the weak endowment lower bound if, for each $P\in{\mathcal{P}}^{N}$, $\varphi\left(P\right)$ satisfies the weak endowment lower bound at $P$.
As we mentioned in the introduction, the weak endowment lower bound is economically relevant as it (i) gives agents the right to veto some undesirable exchanges; and (ii) helps the social planner to simplify the problem, and this simplicity is without loss of generality because, together with our other properties, the weak endowment lower bound also corresponds to individual rationality (see Proposition 2.11 below).
Remark 2.
The relation between the weak endowment lower bound, individual rationality, the strong endowment lower bound, and balancedness is as follows.
  • The strong endowment lower bound implies each of the other three properties, but the converse is false.
  • Even in the presence of balancedness, the weak endowment lower bound and individual rationality are logically independent.
The independence is demonstrated in Example B.1 in Appendix B.2. $\diamond$
We conclude this subsection by stating some useful relations between our properties.
Proposition 2.10.
If a rule $\varphi$ satisfies drop strategy-proofness and the weak endowment lower bound, then it is truncation-proof.
Example B.2 in Appendix B.2 demonstrates that truncation-proofness and the weak endowment lower bound do not jointly imply drop strategy-proofness. Example B.3 shows that truncation-proofness is not implied by drop strategy-proofness alone.
Proposition 2.11.
If a rule $\varphi$ satisfies balancedness, the weak endowment lower bound, and truncation-proofness, then it is individually rational.
The proofs are given in Appendix A.
2.3. Top Trading Cycles
A cycle is a sequence
$C=\left(i_{1},o_{i_{2}},i_{2},o_{i_{3}},\dots,i_{k-1},o_{i_{k}},i_{k},o_{i_{k+1}},i_{k+1}=i_{1}\right)$
of $k\geq 1$ distinct agents and objects such that each agent on the cycle points to her favorite object among some subset of $O$ and each object points to its owner. The set of agents on $C$ is denoted by $N\left(C\right)=\left\{i_{1},i_{2},\dots,i_{k}\right\}$. Similarly, the set of objects on $C$ is denoted by $O\left(C\right)$.
An allocation $\mu$ executes the cycle $C$ if $\mu$ assigns each agent in $N\left(C\right)$ the object to which she points in $C$, i.e., for each $i_{\ell}\in N\left(C\right)$, $o_{i_{\ell+1}}\in\mu_{i_{\ell}}$.
We consider a generalized version of the Top Trading Cycles (TTC) procedure. At each step, every agent points to her favorite remaining object, and every remaining object points to its owner. At least one cycle obtains.
Our proofs rely on a variant of TTC that executes only a single cycle at each step. If multiple cycles obtain, then we exectue only the smallest cycle, i.e., the cycle containing the agent identified by the smallest natural number. That is, we assign each agent on the smallest cycle the object to which she points, then clear all objects on the cycle. It is well-known that this procedure is equivalent to the variant of TTC that, at each step, executes all prevailing cycles.
Formally, for each profile $P\in{\mathcal{P}}^{N}$, TTC selects the allocation $\varphi^{\text{TTC}}\left(P\right)$ determined by the following procedure, which we call $\text{TTC}\left(P\right)$.
Algorithm: $\text{TTC}\left(P\right)$
Input::
A preference profile $P\in{\mathcal{P}}^{N}$.
Output::
An allocation $\varphi^{\text{TTC}}\left(P\right)$.
Step 0.:
Define $\mu^{0}\coloneqq\left(\emptyset\right)_{i\in N}$ and $O^{1}\coloneqq O$.
Step 1.:
Construct a bipartite directed graph with independent vertex sets $N$ and $O^{1}$. For each agent $i\in N$, there is a directed edge from $i$ to agent $i$’s most-preferred object in $O^{1}$, $\top_{P_{i}}\left(O^{1}\right)$. For each object $o\in O^{1}$, there is a directed edge from $o$ to its owner, $\omega^{-1}\left(o\right)$. At least one cycle exists. Let $C_{1}^{1},C_{2}^{1},\dots,C_{k}^{1}$ denote the cycles that obtain at Step 1, and assume that $\min N\left(C_{1}^{1}\right)<\min N\left(C_{2}^{1}\right)<\cdots<\min N\left(C_{k}^{1}\right)$. Assign each agent $i\in N\left(C_{1}^{1}\right)$ the object to which she points on $C_{1}^{1}$, namely $\top_{P_{i}}\left(O^{1}\right)$; that is, let $\mu^{1}=\left(\mu_{i}^{1}\right)_{i\in N}$ be the allocation such that (i) for all $i\in N\left(C_{1}^{1}\right)$, $\mu_{i}^{1}=\mu_{i}^{0}\cup\left\{\top_{P_{i}}\left(O^{1}\right)\right\}$, and (ii) for all $i\in N\backslash N\left(C_{1}^{1}\right)$, $\mu_{i}^{1}=\mu_{i}^{0}$. Clear all objects in $O\left(C_{1}^{1}\right)$; that is, let $O^{2}\coloneqq O^{1}\backslash O\left(C_{1}^{1}\right)$. If $O^{2}\neq\emptyset$, then proceed to Step 2; otherwise, proceed to Termination.
Step $t\geq 2$.:
Construct a bipartite directed graph with independent vertex sets $N$ and $O^{t}$. For each agent $i\in N$, there is a directed edge from $i$ to agent $i$’s most-preferred object in $O^{t}$, $\top_{P_{i}}\left(O^{t}\right)$. For each object $o\in O^{t}$, there is a directed edge from $o$ to its owner, $\omega^{-1}\left(o\right)$. At least one cycle exists. Let $C_{1}^{t},C_{2}^{t},\dots,C_{k}^{t}$ denote the cycles that obtain at Step $t$, and assume that $\min N\left(C_{1}^{t}\right)<\min N\left(C_{2}^{t}\right)<\cdots<\min N\left(C_{k}^{t}\right)$. Assign each agent $i\in N\left(C_{1}^{t}\right)$ the object to which she points on $C_{1}^{t}$, namely $\top_{P_{i}}\left(O^{t}\right)$; that is, let $\mu^{t}=\left(\mu_{i}^{t}\right)_{i\in N}$ be the allocation such that (i) for all $i\in N\left(C_{1}^{t}\right)$, $\mu_{i}^{t}=\mu_{i}^{t-1}\cup\left\{\top_{P_{i}}\left(O^{t}\right)\right\}$, and (ii) for all $i\in N\backslash N\left(C_{1}^{t}\right)$, $\mu_{i}^{t}=\mu_{i}^{t-1}$. Clear all objects in $O\left(C_{1}^{t}\right)$; that is, let $O^{t+1}\coloneqq O^{t}\backslash O\left(C_{1}^{t}\right)$. If $O^{t+1}\neq\emptyset$, then proceed to Step $t+1$; otherwise, proceed to Termination.
Termination.:
Because $O$ is finite, and $\left|O^{1}\right|>\left|O^{2}\right|>\cdots>\left|O^{t}\right|$, the procedure terminates at some step $T$. Return the allocation $\varphi^{\text{TTC}}\left(P\right)\coloneqq\mu^{T}$.
In what follows, we focus on the rule that always selects the TTC allocation $\varphi^{\text{TTC}}\left(P\right)$ for each profile $P\in\mathcal{P}^{N}$. We call such a rule the generalized Top Trading Cycles rule, or TTC in short.
3. A characterization of TTC
We first show that TTC satisfies all of our properties.
Proposition 3.1.
TTC satisfies Pareto efficiency, balancedness, the weak endowment lower bound, drop strategy-proofness, and truncation-proofness.
The proof of Proposition 3.1 follows immediately from Altuntaş et al. (2023, Theorem 1) and Proposition 2.10. We now state our main result, a characterization of TTC.
Theorem 3.2.
A rule satisfies
  • Pareto efficiency,
  • balancedness,
  • weak endowment lower bound, and
  • truncation-proofness (or drop strategy-proofness)
if and only if it is TTC.
The proof of uniqueness is given in Appendix A. Before stating the intuition for our proof strategy, we make two observations.
First, it is unexpected that individual rationality does not appear in Theorem 3.2, and we do not use it for our proof. Second, by Remark 2, we see that Altuntaş et al. (2023, Theorem 1) is a corollary of Theorem 3.2.
Corollary 3.3 (Altuntaş et al., 2023).
A rule satisfies Pareto efficiency, the strong endowment lower bound, and drop strategy-proofness if and only if it is TTC.
Second, since the Shapley-Scarf model, in which each agent is endowed with a single object and has the need for a single object, is a special case of our model, our characterization of TTC is also valid for the Shapley-Scarf model. To emphasize this point, in what follows we firstly formally redefine the Shapley-Scarf model.
3.1. The Shapley-Scarf model
Recall that $N=\{1,\ldots,n\}$. Now, let $O=\{o_{1},\ldots,o_{n}\}$, and for each $i\in N$, $|\omega_{i}|=1$. Thus, it is without loss of generality to assume that $\omega_{i}=o_{i}$ for each $i\in N$. Moreover, in this model, each agent consumes exactly one object. Thus, an allocation is a bijection $\mu:N\to O$ (or, equivalently, a profile $\mu=(\mu_{i})_{i\in N}$ of distinct objects). [^13] Though our general model defines allocations more broadly, i.e., as functions $\overline{\mu}:N\to 2^{O}$ satisfying the feasibility constraints, under lexicographic preferences on $2^{O}$ any individually rational allocation must give exactly one object to each agent. Hence, the restriction to bijections $\mu:N\to O$ is not a substantive one. Therefore, in the Shapley-Scarf model, all allocations are balanced.
Since each agent only owns one object, it is easy to see that the weak endowment lower bound, individual rationality, and the strong endowment lower bound, are all equivalent. Finally, the definitions of drop strategy-proofness and truncation-proofness are exactly the same as before. Nevertheless, in this model, it is helpful to reformulate the definitions of truncation strategies and truncation-proofness as follows.
Let $P_{i},P^{\prime}_{i}\in{\mathcal{P}}$. We say that $P^{\prime}_{i}$ is a truncation strategy for $P_{i}$ if
  1. (1)
    $\{o\in O\mid o\mathbin{R^{\prime}_{i}}o_{i}\}\subsetneq\{o\in O\mid o\mathbin{R_{i}}o_{i}\}$, and
  2. (2)
    For each $o,o^{\prime}\in O\setminus o_{i}$, $o\mathbin{P^{\prime}_{i}}o^{\prime}$ if and only if $o\mathbin{P_{i}}o^{\prime}$.
Intuitively, $P^{\prime}_{i}$ is obtained from $P_{i}$ by “pushing up” agent $i$’s own object $o_{i}$, whilst leaving the relative rankings of all other objects unchanged. Let ${\mathcal{T}}\left(P_{i}\right)$ denote the set of all truncation strategies of $P_{i}$. Note that the definition from Section 2 is equivalent to this definition when each agent is endowed with one object.
We next show that our TTC characterization is also valid for the Shapley-Scarf model. [^14] Altuntaş et al. (2023, Corollary 3) show that for the Shapley-Scarf model, TTC is characterized by Pareto efficiency, individual rationality, and drop strategy-proofness. In view of Proposition 2.10, we generalize their result by showing that drop strategy-proofness can be replaced with truncation-proofness.
Theorem 3.4.
In the Shapley-Scarf model, a rule satisfies
  • Pareto efficiency,
  • individual rationality, and
  • truncation-proofness
if and only if it is TTC.
To help with the exposition, we give a proof of Theorem 3.4 below. Though our proof is remarkably simple, it shares the same ideas as the proof of Theorem 3.2, and it should provide intuition for the proof of the general result (Theorem 3.2).
3.1.1. Proof of Theorem 3.4
It is known that TTC satisfies all properties. Therefore, it suffices to prove the uniqueness. Toward contradiction, suppose that $\varphi$ satisfies the properties but $\varphi\neq\varphi^{TTC}$.
We borrow the clever notions of size and similarity from Sethuraman (2016) and Ekici (2023), respectively.
For each $P\in\mathcal{P}^{N}$, and each $i\in N$, let $s_{i}(P_{i})$ be the number of objects that are weakly better than $o_{i}$ at $P_{i}$, i.e., $s_{i}(P_{i})=|\{o\in O\mid o\mathbin{R_{i}}o_{i}\}|$. Then the size of $P$ is $s(P)=\sum_{i\in N}s_{i}(P_{i})$.
For each $P\in\mathcal{P}^{N}$, and each $t\in\mathbb{N}$, let $\mathcal{C}_{t}(P)$ be the set of cycles that obtain at step $t$ of $\text{TTC}\left(P\right)$. Similarly, let $C_{t}(P)\in\mathcal{C}_{t}(P)$ be the cycle that is executed at step $t$ of $\text{TTC}\left(P\right)$. The similarity between $\varphi$ and $\varphi^{TTC}$ is a function $\rho:\mathcal{P}^{N}\to\mathbb{N}$ such that, for all $P\in\mathcal{P}$, $\rho(P)$ is defined as follows.
  • Step 1. If $\varphi(P)$ does not execute $C_{1}(P)$, then $\rho(P)=1$. Suppose $\varphi(P)$ executes $C_{1}(P)$. If $\text{TTC}(P)$ terminates at Step 1, then $\rho(P)=|O|$; otherwise, proceed to Step $2$.
  • Step $t$ ($\geq 2$). If $\varphi(P)$ does not execute $C_{t}(P)$, then $\rho(P)=t$. Suppose $\varphi(P)$ executes $C_{t}(P)$. If $\text{TTC}(P)$ terminates at Step $t$, then $\rho(P)=|O|$; otherwise, proceed to Step $t+1$.
Note that, for all $P\in\mathcal{P}^{N}$, (i) $\rho(P)\leq|O|$, and (ii) $\rho(P)=|O|$ if and only if $\varphi(P)=\varphi^{TTC}(P)$.
Suppose that $\min_{P\in\mathcal{P}^{N}}\rho\left(P\right)=t$. Then $\varphi\neq\varphi^{TTC}(P)$ implies that $t<|O|$. Among all profiles in $\{P\in\mathcal{P}^{N}\mid\rho(P)=t\}$, let $P$ be one whose size is smallest. Hence, for any other profile $\tilde{P}\neq P$, we have
  • (1)
    $\rho(P)<\rho(\tilde{P})$, or
  • (2)
    $\rho(P)=\rho(\tilde{P})$ and $s(P)\leq s(\tilde{P})$.
Since $\rho(P)=t$, $\varphi(P)$ executes the cycles $C_{1}(P),C_{2}(P),\dots,C_{t-1}(P)$ but not $C_{t}(P)$. Let $\mu\coloneqq\varphi^{TTC}(P)$, $C\coloneqq C_{t}(P)$, and $O^{t}\coloneqq O\backslash\bigcup_{\tau=1}^{t-1}O\left(C_{\tau}\left(P\right)\right)$. Thus, there is an agent $i\in N(C)$ such that, although agent $i$ points to $\mu_{i}$ on $C$, she does not receive $\mu_{i}$ at $\varphi(P)$, i.e., $\mu_{i}\neq\varphi_{i}(P)$. Note that, by the definition of $\text{TTC}(P)$, $\mu_{i}$ is agent $i$’s most-preferred object in $O^{t}$ at $P_{i}$, i.e., $\mu_{i}=\top_{P_{i}}\left(O^{t}\right)$. Thus,
(2) $\mu_{i}\mathbin{P_{i}}\varphi_{i}(P).$
If $|N(C)|=1$, then $C=(i,o_{i},i)$ and $\mu_{i}=o_{i}\mathbin{P_{i}}\varphi_{i}(P)$, which violates individual rationality. Thus, $|N(C)|\geq 2$.
Next, we show that $\varphi_{i}(P)=o_{i}$. Toward contradiction, suppose that $\varphi_{i}(P)\neq o_{i}$. By (2) and individual rationality, $\mu_{i}\mathbin{P_{i}}\varphi_{i}(P)\mathbin{P_{i}}o_{i}$.
Let $P^{\prime}_{i}$ be the truncation of $P_{i}$ at $\mu_{i}$, i.e., $P^{\prime}_{i}:\dots,\mu_{i},o_{i},\dots$. [^15] A similar argument using drop strategy-proofness applies if $P^{\prime}_{i}$ is obtained from $P_{i}$ by dropping object $\varphi_{i}(P)$. Let $P^{\prime}\coloneqq(P^{\prime}_{i},P_{-i})$. Then $s_{i}(P^{\prime}_{i})<s_{i}(P_{i})$ and $s(P^{\prime})<s(P)$.
By the definition of TTC, for each step $\tau\in\{1,\dots,t\}$, the procedures $\text{TTC}(P)$ and $\text{TTC}(P^{\prime})$ generate and execute the same cycles, i.e., for each $\tau$, $C_{\tau}(P)=C_{\tau}(P^{\prime})$. Moreover, the choice of $P$ implies that $\rho(P^{\prime})\geq\rho(P)=t$. Thus, $\varphi(P^{\prime})$ executes all cycles in $\{C_{1}(P),\dots,C_{t-1}(P)\}$. Thus, $\varphi_{i}(P^{\prime})\in O^{t}$. Moreover, since $\varphi$ is truncation-proof, we know that $i$ cannot be better of by misreporting $P^{\prime}_{i}$ at $P$. Together with (2), we conclude that $\mu_{i}\neq\varphi_{i}(P^{\prime})$. Consequently, $\varphi_{(}P^{\prime})$ does not execute $C_{t}(P^{\prime})$, which means that $\rho(P^{\prime})=t$. But then we have $\rho(P^{\prime})=\rho(P)$ and $s(P^{\prime})<s(P)$, which contradicts the choice of $P$! Thus, we conclude that $\varphi_{i}(P)=o_{i}$.
Let $j$ be the agent who points to $o_{i}$ on $C$, i.e., $j\in N(C)$ is such that $\mu_{j}=o_{i}$. Then $\varphi_{i}(P)=o_{i}=\mu_{j}$ implies that $\varphi_{j}(P)\neq\mu_{j}$. Thus, $\mu_{j}\mathbin{P_{j}}\varphi_{j}(P).$ A similar argument shows that $\varphi_{j}(P)=o_{j}$.
By repeating the same argument for each agent in $N(C)$, we can show that each agent in $N(C)$ is assigned her endowment at $\varphi(P)$, i.e., for each $i^{\prime}\in N(C)$, $\varphi_{i^{\prime}}(P)=o_{i^{\prime}}$. But then $\varphi$ is not Pareto efficient, because the agents in $N(C)$ can benefit by executing $C$. This contradiction completes the proof of Theorem 3.4.
3.2. Additive and responsive preferences
In this subsection, we consider two more general preference domains: (strictly) additive preferences and responsive preferences. A preference relation $P_{i}$ is (strictly) additive if it is represented by an additive utility function $u_{i}:2^{O}\to\mathbb{R}_{++}$ such that, (1) for all $X\in 2^{O}$, $u_{i}(X)=\sum_{x\in X}u_{i}(x)$, and (2) for any two distinct bundles $X,Y\in 2^{O}$, $u_{i}(X)\neq u_{i}(Y)$. Let $\mathcal{A}$ denote the set of additive preference relations on $2^{O}$. A preference relation $P_{i}$ on $2^{O}$ is responsive if for any bundle $X\in 2^{O}$ and any two distinct objects $y,z\in O\setminus X$, $y\mathbin{P_{i}}z$ if and only if $(X\cup y)\mathbin{P_{i}}(X\cup z)$. Let $\mathcal{R}$ denote the set of responsive preference relations on $2^{O}$. Note that $\mathcal{A}\subsetneq\mathcal{R}$.
Given a responsive preference relation $P_{i}\in\mathcal{R}$, let $\succ^{P_{i}}$ denote the induced linear order on $O$ such that, for all $x,y\in O$ $x\mathbin{P_{i}}y$ if and only if $x\succ^{P_{i}}y$. Given a profile $P=(P_{i})_{i\in N}\in\mathcal{R}^{N}$, we denote $\succ^{P}=(\succ^{P_{i}})_{i\in N}$. We restrict attention to the class of rules $\varphi$ that depend only on the profile of induced linear orders associated with a given preference profile, i.e., such that for all $P,P^{\prime}\in\mathcal{R}^{N}$, $\succ^{P}=\succ^{P^{\prime}}$ implies $\varphi(P)=\varphi(P^{\prime})$. Equivalently, we assume that agents report only their induced linear orders $\succ^{P}$, while they evaluate allocations based on their actual preferences $P$.
Assuming that agents report only their induced linear orders, drop strategies and truncation strategies are defined as before, i.e., agent $i$ can misreport her induced linear order by dropping objects from her true induced linear order $\succ^{P_{i}}$. The rest of the properties are defined as in Section 2.2.161616Formally, any property that was required to hold for any $P\in\mathcal{P}^{N}$ is now required to hold for any $P\in\mathcal{A}^{N}$ on the additive domain, and for any $P\in\mathcal{R}^{N}$ on the responsive domain.
It is easy to see that TTC satisfies balancedness and the weak endowment lower bound on $\mathcal{R}^{N}$. Though TTC is not Pareto efficient on $\mathcal{A}^{N}$, it satisfies individual-good efficiency (ig-efficiency) on all of $\mathcal{R}^{N}$, which ensures that there is no single-object exchange that improves the welfare of all participants (see, e.g., Aziz et al. (2019), Biró et al. (2022a), Caspari (2020), Coreno and Balbuzanov (2022)).
Definition 3.5 (ig-efficiency).
A rule $\varphi$ is ig-efficient if, for each $P\in{\mathcal{R}}^{N}$, there is no cycle $C=\left(i_{1},o_{2},i_{2},o_{3},\dots,i_{k},o_{k+1},i_{k+1}=i_{1}\right)$ of agents and objects such that, for all $\ell\in\left\{1,\dots,k\right\}$, $o_{\ell}\in\varphi_{i}\left(P\right)$ and $o_{\ell+1}P_{i_{\ell}}o_{\ell}$.
On the lexicographic preference domain, Pareto efficiency is equivalent to ig-efficiency. It also means that our TTC characterization on the lexicographic preference domain is valid if we use ig-efficiency.
Corollary 3.6.
On the lexicographic domain, a rule satisfies ig-efficiency, balancedness, the weak endowment lower bound, and truncation-proofness (or drop strategy-proofness) if and only if it is TTC.
The following example illustrates that TTC is not drop strategy-proof, even if agents have additive preferences.
Example 3.7.
Let $N=\{1,2,3\}$ and $\omega=\left(\left\{a,b\right\},\left\{c,d\right\},\left\{e,f\right\}\right)$. Let $P_{2},P_{3}\in\mathcal{P}$ be two lexicographic preference relations such that
$P_{2}:e,f,b,\dots;\quad P_{3}:a,c,d,\dots.$
Let $P_{1}$ be an additive preference relation such that $\succ^{P_{1}}:c,f,e,d,a,b$ and $\{d,f\}\mathbin{P_{1}}\{c,b\}$. Then $\varphi^{\text{TTC}}\left(P\right)=\left(\left\{c,b\right\},\left\{e,f\right\},\left\{a,d\right\}\right)$.
Now assume that agent $1$ drops object $c$ from her induced linear order, i.e., she reports $P^{\prime}_{1}$ such that $\succ^{P_{1}}:f,e,d,a,b,c.$ Then $\varphi^{\text{TTC}}\left(P^{\prime}_{1},P_{-1}\right)=\left(\left\{d,f\right\},\left\{e,b\right\},\left\{a,c\right\}\right)$. Because $\varphi_{1}^{\text{TTC}}\left(P^{\prime}_{1},P_{-1}\right)=\{d,f\}\mathbin{P_{1}}\{c,b\}=\varphi_{1}^{\text{TTC}}\left(P\right)$, we see that TTC is not drop strategy-proof. $\diamond$
Example 3.7 demonstrates that even if we allow only a single agent, say $i$, to report an additive preference relation which is not lexicographic, then there is a problem $P$ with $P_{i}\in\mathcal{A}\backslash\mathcal{P}$ such that $i$ can manipulate by a drop strategy. [^17] If there are only two agents, then TTC is drop strategy-proof as each agent can only drop the opponent’s endowed objects.
However, one can show that TTC is truncation-proof. The positive results are summarized in the following proposition.
Proposition 3.8.
On the domain of responsive preferences, TTC satisfies ig-efficiency, balancedness, the weak endowment lower bound, individual rationality, and truncation-proofness.
4. Final remarks
We conclude the paper with two final remarks.
4.1. Efficiency
First, one may wonder if we can weaken Pareto efficiency to other efficiency notions, such as pairwise efficiency (Ekici, 2023). [^18] A rule satisfies pairwise efficiency if for each profile $P$, there is no pair of agents $i,j$ such that $\varphi_{i}(P)\mathbin{P_{j}}\varphi_{j}(P)$ and $\varphi_{j}(P)\mathbin{P_{i}}\varphi_{i}(P)$. Ekici (2023) originally refers to it as pair efficiency. The following example shows that, even for the Shapley–Scarf model, the answer is negative.
Example 4.1 (Pareto efficiency cannot be weakened to pairwise efficiency).
For simplicity, consider the Shapley–Scarf model with three agents, $N=\left\{1,2,3\right\}$, and suppose $\omega=\left(o_{1},o_{2},o_{3}\right)$. Let $P^{*}\in{\mathcal{P}}^{N}$ be such that (endowment is underlined)
$P_{1}^{*}:o_{2},\underline{o_{1}},o_{3},\quad P_{2}^{*}:o_{3},\underline{o_{2}},o_{1},\quad P_{3}^{*}:o_{1},\underline{o_{3}},o_{2}.$
Then the initial allocation $\omega$ is pairwise efficient at $P^{*}$; but it is not Pareto efficient at $P^{*}$ because it is Pareto-dominated by $\varphi^{\text{TTC}}\left(P^{*}\right)=\left(o_{2},o_{3},o_{1}\right)$.
Now, consider the rule $\varphi$ defined as follows.
$\varphi\left(P\right)=\begin{cases}\omega,&\text{if }P=P^{*}\\ \varphi^{\text{TTC}}\left(P\right),&\text{otherwise}.\end{cases}$
Clearly, $\varphi$ is pairwise efficient and individually rational. It is straightforward to show that $\varphi$ is drop strategy-proof and truncation-proof. $\diamond$
4.2. Two larger domains
Second, based on Proposition 3.8 and Corollary 3.6, one may wonder whether our TTC characterization can be extended to more general (and maybe more economically interesting) preference domains, such as additive preferences. The following example shows that the answer is no.
Example 4.2 (Another rule).
Let $N=\{1,2\}$ and $\omega=(\{a,b\},\{c,d\})$. Let $\mu\coloneqq(\omega_{2},\omega_{1})$. Let $\varphi:\mathcal{A}^{N}\to\mathcal{M}$ be defined as follows.
$\varphi\left(P\right)=\begin{cases}\mu,&\text{if }\varphi^{TTC}(P)=\omega\text{, and }\mu\text{ Pareto dominates }\omega;\\ \varphi^{TTC}(P),&\text{otherwise.}\end{cases}$
One can show that $\varphi$ is truncation-proof.
It is easy to see that $\varphi$ satisfies ig-efficiency, balancedness, and the weak endowment lower bound. Thus, we see that TTC is not the unique rule satisfying our properties on $\mathcal{A}^{N}$.
Also, note that such a rule $\varphi$ does not exist on $\mathcal{P}^{N}$: when agents’ preferences are lexicographic, $\varphi^{TTC}(P)=\omega$ is Pareto efficient, and hence “$\mu$ Pareto dominates $\omega$” is always false. $\diamond$
However, there is another more general domain, that shares the spirit of lexicographic preferences, the conditionally lexicographic preferences. The next question is whether our TTC characterization applies to this domain as well. The answer is partially, yes. Once we adjust the definition of the weak endowment lower bound and drop strategy-proofness for conditionally lexicographic preferences, we can show that still, TTC is the only rule that satisfies Pareto efficiency, the (conditional) weak endowment lower bound, and the adjusted drop strategy-proofness. [^19] For conditionally lexicographic preferences, agents do not have a meaningful linear order over objects. Thus, the original definition of drop strategy-proofness is not well-defined. We provide formal definitions of conditionally lexicographic preferences, the (conditional) weak endowment lower bound, and adjusted drop strategy-proofness in the full version of the paper. The details are available upon request.
Theorem 4.3.
On the domain of conditionally lexicographic preferences, a rule satisfies Pareto efficiency, balancedness, the (conditional) weak endowment lower bound, and adjusted drop strategy-proofness, if and only if it is TTC.
Example 4.2 and Theorem 4.3 together points to an interesting issue: Although our TTC characterization is not valid for more general domains with substitution, such as additive preferences, it is valid for more general domains with complementarity. We believe this observation could be useful in the future, as there are not as many studies on complementary preferences.
Appendix A Omitted proofs
Proof of Proposition 2.10
Suppose that $\varphi$ satisfies drop strategy-proofness and the weak endowment lower bound. Let $P^{\prime}_{i}$ be obtained from $P_{i}$ by dropping some subset $X\subseteq O\backslash\omega_{i}$. Suppose that $X=\left\{x_{1},x_{2},\dots,x_{k}\right\}$, where $x_{1}P_{i}x_{2}P_{i}\cdots P_{i}x_{k}$. Then $P^{\prime}_{i}$ is obtained from a sequence of $k$ drop strategies. That is, $P^{\prime}_{i}=P_{i}^{k}$, where $P_{i}^{0}=P_{i}$ and $P_{i}^{1},\dots,P_{i}^{k}$ are such that, for each $\ell\in\left\{1,\dots,k\right\}$, $P_{i}^{\ell}$ is obtained from $P_{i}^{\ell-1}$ by dropping object $x_{\ell}$.
Claim 1.
For each $\ell\in\left\{1,\dots,k\right\}$, $\varphi_{i}\left(P_{i}^{\ell-1},P_{-i}\right)R_{i}\varphi_{i}\left(P_{i}^{\ell},P_{-i}\right)$.
Proof of Claim 1.
The proof is by induction on $\ell$. Clearly, $\varphi_{i}\left(P\right)=\varphi_{i}\left(P_{i}^{0},P_{-i}\right)R_{i}\varphi_{i}\left(P_{i}^{1},P_{-i}\right)$ by drop strategy-proofness. For the inductive step, suppose that $\ell\in\left\{1,\dots,k-1\right\}$ is such that
$\varphi_{i}\left(P\right)=\varphi_{i}\left(P_{i}^{0},P_{-i}\right)R_{i}\varphi_{i}\left(P_{i}^{1},P_{-i}\right)R_{i}\cdots R_{i}\varphi_{i}\left(P_{i}^{\ell},P_{-i}\right).$
It suffices to show that $\varphi_{i}\left(P_{i}^{\ell},P_{-i}\right)R_{i}\varphi_{i}\left(P_{i}^{\ell+1},P_{-i}\right)$. By drop strategy-proofness, we have $\varphi_{i}\left(P_{i}^{\ell},P_{-i}\right)R_{i}^{\ell}\varphi_{i}\left(P_{i}^{\ell+1},P_{-i}\right)$. By the weak endowment lower bound, $\varphi_{i}\left(P_{i}^{\ell},P_{-i}\right)\subseteq O\backslash\left\{x_{1},\dots,x_{\ell}\right\}$ and $\varphi_{i}\left(P_{i}^{\ell+1},P_{-i}\right)\subseteq O\backslash\left\{x_{1},\dots,x_{\ell+1}\right\}\subseteq O\backslash\left\{x_{1},\dots,x_{\ell}\right\}$. Because $P_{i}^{\ell}$ agrees with $P_{i}$ on $O\backslash\left\{x_{1},\dots,x_{\ell}\right\}$, we have $\varphi_{i}\left(P_{i}^{\ell},P_{-i}\right)R_{i}\varphi_{i}\left(P_{i}^{\ell+1},P_{-i}\right)$, as desired. ∎
It follows from Claim 1 that
$\varphi_{i}\left(P\right)=\varphi_{i}\left(P_{i}^{0},P_{-i}\right)R_{i}\varphi_{i}\left(P_{i}^{1},P_{-i}\right)R_{i}\cdots R_{i}\varphi_{i}\left(P_{i}^{k},P_{-i}\right).$
Therefore, $\varphi_{i}\left(P\right)R_{i}\varphi_{i}\left(P_{i}^{k},P_{-i}\right)$. $\blacksquare$
Proof of Proposition 2.11
Suppose that $\varphi$ satisfies the stated properties. Let $P\in\mathcal{P}^{N}$ and $i\in N$. Let $P^{\prime}_{i}\in{\mathcal{T}}\left(P_{i}\right)$ be a truncation strategy such that all non-endowed objects are dropped, i.e., for each $o\in O\setminus\omega_{i}$, $w_{i}(P_{i})\mathbin{P^{\prime}_{i}}o$. By the weak endowment lower bound, $\varphi_{i}\left(P^{\prime}_{i},P_{-i}\right)\subseteq\omega_{i}$. Because $\varphi$ is balanced, we must have $\left|\varphi_{i}\left(P^{\prime}_{i},P_{-i}\right)\right|=\left|\omega_{i}\right|$ and, hence, $\varphi_{i}\left(P^{\prime}_{i},P_{-i}\right)=\omega_{i}$. It follows from truncation-proofness that $\varphi_{i}\left(P\right)R_{i}\varphi_{i}\left(P^{\prime}_{i},P_{-i}\right)=\omega_{i}$. $\blacksquare$
Proof of Theorem 3.2
Toward contradiction, suppose that $\varphi$ satisfies the properties but $\varphi\neq\varphi^{TTC}$.
For each $P\in\mathcal{P}^{N}$, and each $t\in\mathbb{N}$, let $\mathcal{C}_{t}(P)$ be the set of cycles that obtain at step $t$ of $\text{TTC}\left(P\right)$. Similarly, let $C_{t}(P)\in\mathcal{C}_{t}(P)$ be the cycle that is executed at step $t$ of $\text{TTC}\left(P\right)$.
Denote the size of a profile $P$ by $s\left(P\right)=\sum_{i\in N}\left|\left\{o\in O\mid oR_{i}\bot_{P_{i}}\left(\omega_{i}\right)\right\}\right|$. The similarity $\rho$ is defined exactly as in the proof of Theorem 3.4.
Let $P$ be a profile such that for any other profile $\tilde{P}\neq P$, either
  • (1)
    $\rho(P)<\rho(\tilde{P})$, or
  • (2)
    $\rho(P)=\rho(\tilde{P})$ and $s(P)\leq s(\tilde{P})$.
Let $t\coloneqq\rho(P)$. Since $\rho(P)=t$, $\varphi(P)$ executes the cycles $C_{1}(P),C_{2}(P),\dots,C_{t-1}(P)$ but not $C_{t}(P)$. Let $C\coloneqq C_{t}(P)$ and $O^{t}\coloneqq O\backslash\bigcup_{\tau=1}^{t-1}O\left(C_{\tau}\left(P\right)\right)$. Suppose that
$C=\left(i_{1},o_{i_{2}},i_{2},o_{i_{3}},\dots,i_{k-1},o_{i_{k}},i_{k},o_{i_{k+1}},i_{k+1}=i_{1}\right).$
Because $\varphi(P)$ does not execute $C$, there is an agent $i_{\ell}\in N(C)$ such that, although agent $i_{\ell}$ points to $o_{i_{\ell+1}}$ on $C$, she does not receive $o_{i_{\ell+1}}$ at $\varphi(P)$, i.e., $o_{i_{\ell+1}}\notin\varphi_{i}(P)$. Note that, by the definition of $\text{TTC}(P)$, $o_{i_{\ell}}$ is agent $i_{\ell}$’s most-preferred object in $O^{t}$ at $P_{i_{\ell}}$, i.e., $o_{i_{\ell}}=\top_{P_{i_{\ell}}}\left(O^{t}\right)$. Without loss of generality, let $i_{\ell}=i_{k}$. Thus, $o_{i_{1}}\notin\varphi_{i_{k}}(P)$.
Let $X_{i_{k}}$ be the set of all objects in $O\backslash O^{t}$ that are assigned to agent $i_{k}$, i.e., $X_{i_{k}}=\varphi_{i_{k}}(P)\backslash O^{t}$. We next show that, apart from the objects in $X_{i}$, agent $i_{k}$ only receives the remainder of her endowment.
Claim 2.
$\varphi_{i_{k}}(P)\cap O^{t}=\omega_{i_{k}}\cap O^{t}$.
Proof of Claim 2.
Suppose otherwise. By balancedness and the fact that $\left|\varphi_{i_{k}}\left(P\right)\backslash O^{t}\right|=\left|\omega_{i_{k}}\backslash O^{t}\right|$, we must have $\left|\varphi_{i_{k}}\left(P\right)\cap O^{t}\right|=\left|\omega_{i_{k}}\cap O^{t}\right|$. Consequently, $\varphi_{i_{k}}\left(P\right)\cap O^{t}\neq\omega_{i_{k}}\cap O^{t}$ implies there is an object $o^{\prime}\in\varphi_{i_{k}}(P)\cap O^{t}$ with $o^{\prime}\not\in\omega_{i_{k}}$.
By the definition of $\text{TTC}(P)$, we know that $o_{i_{1}}$ is agent $i_{k}$’s most-preferred object in $O^{t}$, i.e., for each $o\in O^{t}$, $o_{i_{1}}\mathbin{R_{i_{k}}}o$. Because $o^{\prime}\in\varphi_{i_{k}}(P)\cap O^{t}$ and $o_{i_{1}}\notin\varphi_{i_{k}}(P)$, we have $o_{i_{1}}\mathbin{P_{i_{k}}}o^{\prime}$. By the weak endowment lower bound,
$o^{\prime}\mathbin{P_{i_{k}}}w_{i_{k}}(P_{i_{k}}).$
Let $P^{\prime}_{i_{k}}$ be the truncation of $P_{i_{k}}$ at $o_{i_{1}}$. That is, $P^{\prime}_{i_{k}}$ is obtained from $P_{i_{k}}$ by dropping the set $\{o\in O\backslash\omega_{i_{k}}\mid o_{i_{1}}\mathbin{P_{i_{k}}}o\}$. [^20] If we use drop strategy-proofness, we can consider another preference $P^{\prime}_{i_{k}}$ obtained from $P_{i_{k}}$ by dropping object $o^{\prime}$. Let $P^{\prime}\coloneqq(P^{\prime}_{i_{k}},P_{-i_{k}})$.
By the definition of $\text{TTC}(P)$, we see that until step $t$, all top trading cycles that are executed via TTC at $P$ and $P^{\prime}$ are the same, i.e., $C_{\tau}(P)=C_{\tau}(P^{\prime})$ for $\tau=1,\ldots,t$. By the selection of $P$, we know that $\rho(P)=t\leq\rho(P^{\prime})$. Thus, all cycles in $\{C_{1}(P),\dots,C_{t-1}(P)\}$ are executed at $\varphi(P^{\prime})$. Therefore,
$X_{i_{k}}\subseteq\varphi_{i_{k}}(P^{\prime}).$
Since $\varphi$ is truncation-proof, we know that $i_{k}$ cannot be better off by misreporting $P^{\prime}_{i_{k}}$ at $P$. Together with $X_{i_{k}}\subseteq\varphi_{i_{k}}(P^{\prime})$ and $o_{i_{k}}\notin\varphi_{i_{k}}(P)$, we conclude that $i_{k}$ cannot receive $o_{i_{1}}$ at $\varphi(P^{\prime})$, i.e., $o_{i_{1}}\notin\varphi_{i_{k}}(P^{\prime})$. Therefore, $C$ is not executed at $\varphi(P^{\prime})$, which means that $\rho(P^{\prime})\leq t$. As a result, we find that $\rho(P^{\prime})=t$. However, since $P^{\prime}_{i_{k}}$ is obtained from $P_{i_{k}}$ by dropping the subset $\{o\in O\backslash\omega_{i_{k}}\mid o_{i_{1}}\mathbin{P_{i_{k}}}o\}$, and $o_{i_{1}}\mathbin{P_{i_{k}}}o^{\prime}\mathbin{P_{i_{k}}}w_{i_{k}}(P_{i_{k}})$, we see that $s(P^{\prime})<s(P)$. This contradicts the choice of $P$. ∎
Next, we show that at least two agents are involved in $C$.
Claim 3.
$|N(C)|>1$.
Proof.
Toward contradiction, suppose that $|N(C)|=1$, that is to say, $i_{1}=i_{k}$ and $C=(i_{1},o_{i_{1}},i_{1})$. Then $o_{i_{1}}\in\omega_{i_{k}}$ and Claim 2 implies that $o_{i_{1}}\in\varphi_{i_{k}}(P)$, a contradiction. ∎
By Claim 2 and Claim 3, agent $i_{k-1}$ points to $o_{i_{k}}$ on $C$, yet she does not receive $o_{i_{k}}$. An argument similar to the proof of Claim 2 implies that $\varphi_{i_{k-1}}(P^{\prime})\cap O^{t}=\omega_{i_{k-1}}\cap O^{t}$.
Proceeding by induction, we conclude that for each agent $i_{\ell}$ that is involved in $C$, we have $\varphi_{i_{\ell}}(P)\cap O^{t}=\omega_{i_{\ell}}\cap O^{t}$.
However, this means that $\varphi$ is not ig-efficient (or Pareto efficient), as agents in $N(C)$ can benefit by execution of $C$. This contradiction completes the proof of Theorem 3.2. $\blacksquare$
Proof of Proposition 3.8
It is clear that TTC satisfies ig-efficiency, balancedness, the weak endowment lower bound, and individual rationality. It remains to show that it is truncation-proof.
As we mentioned previously, for TTC we only need to consider the induced linear orders over $O$. Fix a profile of induced linear orders $\succ^{P}$. Let $i\in N$, $x\in O\setminus\omega_{i}$, and $\succ^{\prime}_{i}$ be a truncation strategy for $\succ_{i}^{P_{i}}$ at $x$. Let
$\succ^{\prime}\coloneqq(\succ^{\prime}_{i},\succ^{P}_{-i}).$
Let $U\coloneqq\{o\in O\mid o\mathbin{\succ^{P_{i}}}x\}$ be the upper contour set of $x$ at $\succ^{P_{i}}$. There are two cases.
Case one. $\varphi^{TTC}_{i}(\succ^{P})\subseteq U$. Then, by the definition of TTC, we know that all top trading cycles that are obtained at $\succ^{P}$ and at $\succ^{\prime}$ are exactly the same. As a result, $\varphi^{TTC}_{i}(\succ^{P})=\varphi^{TTC}_{i}(\succ^{\prime})$. So, agent $i$ cannot be better off by $\succ_{i}^{\prime}$.
Case two. $\varphi^{TTC}_{i}(\succ^{P})\not\subseteq U$. So agent $i$ receives at least one object in $O\setminus U$ at $\varphi^{TTC}(\succ^{P})$. Assume that at $\varphi^{TTC}(\succ^{P})$, the earliest step that $i$ receives an object in $O\setminus U$ is $t$.
Let $O_{t}\subseteq O$ be the set of objects that are remaining at step $t$ of $\varphi^{TTC}(\succ^{P})$. Since the ranking over $U$ at $\succ^{P_{i}}$ and $\succ^{\prime}_{i}$ are the same, [^21] Formally, $\succ^{P_{i}}\cap(U\times U)=\succ^{\prime}_{i}\cap(U\times U)$. by the definition of TTC, we know that top trading cycles that are obtained from step 1 to step $t-1$ at $\varphi^{TTC}(\succ^{P})$ and $\varphi^{TTC}(\succ^{\prime})$ are the same. Thus, at step $t$ of $\varphi^{TTC}(\succ^{\prime})$, the set of remaining objects is also $O_{t}$. Moreover, we have
(3) $\varphi^{TTC}_{i}(\succ^{P})\setminus O_{t}=\varphi^{TTC}_{i}(\succ^{\prime})\setminus O_{t}.$
By the definition of $\succ^{\prime}_{i}$, we know that at $\succ^{\prime}_{i}$, all objects in $\{o\in O\setminus\omega_{i}\mid x\mathbin{P_{i}}o\}$ are ranked lower than $i$’s worst endowed object ($i$’s worst endowed object is the same at $\succ^{P_{i}}$ and $\succ^{\prime}_{i}$). Thus, by the definition of TTC, from step $t$ of $\varphi^{TTC}(\succ^{\prime})$, agent $i$ only points (and hence receives) his endowed objects. That is, $\varphi^{TTC}_{i}(\succ^{\prime})\cap O_{t}=\omega_{i}\cap O_{t}$.
Again, by the definition of TTC, from step $t$ of $\varphi^{TTC}(\succ^{P})$, agent $i$ only points (and hence receives) some objects that are weakly better (with respect to $\succ^{P_{i}}$) than her endowed objects. In other words, there is a bijection $\sigma:\omega_{i}\cap O_{t}\to\varphi^{TTC}_{i}(\succ^{P})\cap O_{t}$ such that for each $o\in\omega_{i}\cap O_{t}$, $\sigma(o)\mathbin{\succeq^{P_{i}}}o$. Together with $\varphi^{TTC}_{i}(\succ^{\prime})\cap O_{t}=\omega_{i}\cap O_{t}$, we conclude that
(4) for $\text{for each }o\in\varphi^{TTC}_{i}(\succ^{\prime})\cap O_{t},\sigma(o)\mathbin{\succeq^{P_{i}}}o,\text{ where }\sigma(o)\in\varphi^{TTC}_{i}(\succ^{P})\cap O_{t}.$
Finally, recall that $i$’s original preference $P_{i}$ is additive, hence, it can be represented by an utility $u_{i}$ such that for any two objects $a,b,\in O$, $u_{i}(a)>u_{i}(b)$ if and only if $a\succ^{P_{i}}b$. Also, since $P_{i}$ is strict, there is no indifference between any two objects. Thereofre, by (3) and (4), we find that $u_{i}(\varphi^{TTC}_{i}(\succ^{P}))\geq u_{i}(\varphi^{TTC}_{i}(\succ^{\prime}))$. So, agent $i$ cannot be better off by $\succ_{i}^{\prime}$. $\blacksquare$
Appendix B Examples
B.1. Independence of Properties
For each of the examples introduced to establish independence below we indicate the property of Theorems 3.2 and 3.4, it fails (while it satisfies all remaining properties).
Pareto efficiency: the no-trade rule that always selects the initial allocation satisfies all properties except for Pareto efficiency.
The weak endowment lower bound: the serial dictatorships subject to balancedness [^22] That is, for each dictator, the number of objects she can pick up equals to the number of objects endowed to her. satisfy all properties except for the weak endowment lower bound.
Balancedness: Let $N=\{1,2\}$, $O=\{x,y,z\}$ $\omega_{1}=\{x\}$, and $\omega_{2}=\{y,z\}$. For each profile $P$, let $\mu\coloneqq(O\setminus w_{2}(P_{2}),w_{2}(P_{2}))$.
Let $\varphi$ be defined as follows.
$\varphi\left(P\right)=\begin{cases}\mu,&\text{if }\varphi^{TTC}(P)=\omega\\ (\{y,z\},x),&\text{otherwise.}\end{cases}$
It is easy to see that $\varphi$ satisfies all properties except for balancedness.
Truncation-proofness / Drop strategy-proofness: the following example is from Altuntaş et al. (2023). Let $N=\{1,2,3,4\}$, and for each $i\in N$, $\omega_{i}=o_{i}$. Let $P^{*}$ be defined as follows (endowment is underlined).
$P^{*}_{1}:o_{2},o_{4},\underline{o_{1}},o_{3};$
$P^{*}_{2}:o_{1},o_{3},\underline{o_{2}},o_{4};$
$P^{*}_{3}:o_{2},o_{4},\underline{o_{3}},o_{1};$
$P^{*}_{4}:o_{1},o_{3},\underline{o_{4}},o_{2}.$
Let $\varphi$ be defined as follows.
$\varphi\left(P\right)=\begin{cases}(o_{4},o_{3},o_{2},o_{1}),&\text{if }P=P^{*}\\ \varphi^{TTC}(P),&\text{otherwise.}\end{cases}$
Altuntaş et al. (2023) show that $\varphi$ satisfies Pareto efficiency and the strong endowment lower bound (hence balancedness and the weak endowment lower bound), but violates drop strategy-proofness / truncation proofness. To see this, note that agent $1$ receives $o_{4}$ at $\varphi\left(P^{*}\right)$, and she can receive her best object $o_{2}$ by misreporting a drop strategy $P^{\prime}_{1}:o_{2},\underline{o_{1}},o_{3},o_{4}$ or a truncation strategy $P^{\prime\prime}_{1}:o_{2},\underline{o_{1}},o_{4},o_{3}$.
B.2. Other examples
Example B.1 (Remark 2).
Suppose $N=\left\{1,2\right\}$ and $\omega=\left(\left\{a,b,c\right\},\left\{x,y,z\right\}\right)$.
Consider the following profile $(P_{1},P_{2})$ ($\omega_{1},\omega_{2}$ are underlined, respectively)
$P_{1}:x,\underline{a},y,\underline{b},z,\underline{c}\quad\text{and}\quad P_{2}:a,\underline{x},b,\underline{y},c,\underline{z}.$
Consider the allocation $\mu=\left(\left\{x,z,c\right\},\left\{a,b,y\right\}\right)$. Clearly, $\mu$ satisfies individual rationality, balancedness, and the weak endowment lower bound. However, $\mu$ violates the strong endowment lower bound because $\mu_{1}$ does not pairwise dominate $\omega_{1}$. In particular, the second-best object in $\omega_{1}$, $b$, is better than the second-best object in $\mu_{1}$, $z$.
Now consider the allocation $\nu=\left(\left\{b,z,c\right\},\left\{a,x,y\right\}\right)$. Then $\nu$ satisfies balancedness and the weak endowment lower bound. However, $\nu$ is not individually rational because $\omega_{1}P_{1}\nu_{1}$.
Finally, consider
$P^{\prime}_{1}:x,\underline{a},\underline{b},\underline{c},y,z.$
We see that at $(P^{\prime}_{1},P_{2})$, $\mu$ satisfies individual rationality, balancedness, but violates the weak endowment lower bound as $z\in\mu_{1}$ is worse than agent $1$’s worst endowed object $c\in\omega_{1}$.  $\diamond$
Example B.2 (A rule that satisfies truncation-proofness, the weak endowment lower bound and individual rationality, but not drop strategy-proofness.).
Suppose $N=\left\{1,2,3\right\}$, and $\omega=\left(o_{1},o_{2},o_{3}\right)$. Let $P^{*}\in{\mathcal{P}}^{N}$ be such that
$P_{1}^{*}:o_{2},o_{3},o_{1},\quad P_{2}^{*}:o_{3},o_{1},o_{2},\quad P_{3}^{*}:o_{1},o_{2},o_{3}.$
Note that for each $i\in N$, $i$ only has two truncation strategies, i.e.,
[…]
Consider the rule $\varphi$ defined for all $P\in{\mathcal{P}}^{N}$ by
$\varphi\left(P\right)=\begin{cases}\omega,&\text{if }P\in{\mathcal{T}}\left(P_{1}^{*}\right)\times{\mathcal{T}}\left(P_{2}^{*}\right)\times{\mathcal{T}}\left(P_{3}^{*}\right)\\ \varphi^{\text{TTC}}\left(P\right),&\text{otherwise}.\end{cases}$
That is, $\varphi\left(P\right)=\omega$ whenever $P$ is such that, for all $i\in N$, $P_{i}$ is a truncation of $P_{i}^{*}$; otherwise, $\varphi\left(P\right)=\varphi^{\text{TTC}}\left(P\right)$. By the definition of $\varphi$, it is easy to see that $\varphi$ is individually rational. It is straightforward to show that $\varphi$ is truncation-proof. To see that $\varphi$ is not drop strategy-proof, consider
$P^{\prime}_{1}:o_{3},o_{1},o_{2}.$
Then $P^{\prime}_{1}$ is obtained from $P_{1}^{*}$ by dropping object $o_{2}\in O\backslash\omega_{1}$. Because
$\varphi_{1}\left(P^{\prime}_{1},P_{-1}^{*}\right)=\varphi_{1}^{\text{TTC}}\left(P^{\prime}_{1},P_{-1}^{*}\right)=o_{3}P_{1}^{*}o_{1}=\varphi_{1}\left(P^{*}\right),$
agent 1 can manipulate at $P^{*}$ by misreporting $P^{\prime}_{1}\in{\mathcal{D}}\left(P_{1}^{*}\right)$. Note that $P^{\prime}_{1}$ is not a truncation of $P_{1}^{*}$.  $\diamond$
Example B.3 (Drop strategy-proofness does not imply truncation-proofness.).
Suppose $N=\left\{1,2,3\right\}$ and $\omega=\left(o_{1},o_{2},o_{3}\right)$. Let $\varphi$ be such that, for all $P\in{\mathcal{P}}^{N}$,
$\varphi\left(P\right)$ =$=\begin{cases}\left(o_{1},o_{2},o_{3}\right),&\text{if }o_{1}\text{ is agent 1's second-ranked object at }P_{1}\\ \left(o_{2},o_{1},o_{3}\right),&\text{if }o_{2}\text{ is agent 1's second-ranked object at }P_{1}\\ \left(o_{3},o_{2},o_{1}\right),&\text{if }o_{3}\text{ is agent 1's second-ranked object at }P_{1}.\end{cases}$
Then $\varphi$ is drop strategy-proof. Indeed, if $P^{\prime}_{1}$ is obtained from $P_{1}$ by dropping any single object, then agent 1 receives her worst object at $\left(P^{\prime}_{1},P_{-1}\right)$, i.e., $\varphi_{1}\left(P^{\prime}_{1},P_{-1}\right)=\bot_{P_{1}}\left(O\right)$.
Consider
$P_{1}:o_{2},o_{3},o_{1}.$
Note that $\varphi_{1}(P)=o_{3}$ for any $P_{-1}$. Next, consider a truncation strategy by dropping $\{o_{2},o_{3}\}$, i.e.,
$P^{\prime}_{1}:o_{1},o_{2},o_{3}.$
We see that $\varphi_{1}\left(P^{\prime}_{1},P_{-1}\right)=o_{2}$. Since $\varphi_{1}\left(P^{\prime}_{1},P_{-1}\right)=o_{2}\mathbin{P_{1}}o_{3}=\varphi_{1}\left(P\right)$, we find that $\varphi$ is not truncation-proof.  $\diamond$