### v3

# Some Characterizations of TTC in Multiple-Object Reallocation Problems

Abstract.
This paper considers exchange of indivisible objects when agents are endowed with and can consume any bundles. We focus on efficient allocation rules that satisfy a novel participation requirement, the weak endowment lower bound, and which defend against simple manipulation heuristics: drop strategies and truncation strategies. Based on these properties, we obtain characterizations of a generalized version of Top Trading Cycles (TTC) on several domains. On the lexicographic and conditionally lexicographic domains, TTC is characterized by Pareto efficiency, balancedness, the weak endowment lower bound, and truncation proofness (or drop strategy-proofness). On the domain of responsive preferences, similar characterizations are obtained by restricting attention to rules that are “individual-good-based” and weakening Pareto efficiency to individual-good efficiency. For the Shapley-Scarf model, TTC is characterized by Pareto efficiency, individual rationality, and truncation-proofness. The lexicographic and conditionally lexicographic domains are maximal domains on which Pareto efficiency coincides with individual-good efficiency.
Keywords: exchange of indivisible objects; Top Trading Cycles; heuristic manipulation; endowment lower bound.
conference: ; ;
1. Introduction
This paper considers reallocation problems when agents have multi-unit demands. [^1] One could also consider this problem as a barter economy with indivisibility. Here, we use the term problem to emphasize that our approach is in a centralized manner, rather than a decentralized exchange. There is a group of agents, each of whom is initially endowed with some set of heterogeneous and indivisible objects. Each agent has preferences over bundles of objects. A social planner can redistribute the objects to make agents better off, but monetary transfers are forbidden.
There are many examples of such problems. Employees at a firm may exchange tasks, shifts, or equipment (Yu and Zhang, 2020). University students exchange course schedules, while universities exchange students for semester-long exchange programs (Bloch et al., 2020). Teams in professional sports leagues trade players’ contract rights (Coreno and Balbuzanov, 2022). Our model is an extension of the classic Shapley-Scarf model (Shapley and Scarf, 1974), and it is extensively studied in the literature, e.g., by Pápai (2007) and Altuntaş et al. (2023). [^2] Throughout this paper, we shall distinguish between single-object problems, in which agents are endowed with a single object, and multi-object problems, in which agents are endowed with bundles.
We are interested in allocation rules that elicit a minimal amount of relevant information about the agents’ preferences over bundles (e.g., the associated rankings over individual objects) and which satisfy desirable properties, such as efficiency, strategic robustness, and a participation guarantee which ensures that agents are not harmed by the reallocation. Given the incompatibility of Pareto efficiency, strategy-proofness, and individual rationality (Sönmez, 1999), we consider rules that satisfy relaxed versions of strategy-proofness. Our main incentive properties are truncation-proofness and drop strategy-proofness which, loosely speaking, ensure that agents cannot manipulate by vetoing certain objects.
We provide characterizations of generalized versions of Top Trading Cycles (TTC) on three domains of preferences. The first two are the lexicographic and responsive domains, both of which capture substitutability between the objects. The responsive domain is widely studied in economics and, in particular, matching problems, where it plays an important role in stability (Roth, 1985). The lexicographic domain is a narrower domain which fits some important practical examples and is technically tractable. The third domain is the conditionally lexicographic domain, which has been recently studied among computer science researchers (Booth et al., 2010). Conditionally lexicographic preferences are flexible enough to capture complementarity among the objects, yet their compact representation makes them simple to elicit. Our results highlight the usefulness of TTC in these more general environments.
1.1. Overview of the paper
We are interested in allocation rules that are simple to implement in the sense that they elicit a minimal amount of relevant information about the agents’ preferences over bundles. When objects are substitutable, we focus on individual-good-based rules, i.e., rules that elicit only ordinal rankings over individual objects. This approach is common in the theoretical literature (e.g., Aziz et al. (2019)) as well as in practice. For example, in the National Resident Matching Program that matches doctors to hospitals in the United States, each hospital reports only its ranking over individual doctors, even though they may have complicated preferences over sets of doctors (see, e.g., Milgrom (2009), (2011)). The restriction to individual-good-based rules is particularly useful when the set of objects is large, so that agents could not feasibly rank all bundles.
In addition to this simplicity requirement, we seek rules that satisfy desirable properties, such as efficiency, strategic robustness, and a participation guarantee which ensures that agents are not harmed by the reallocation. Ideal properties, such as Pareto efficiency, strategy-proofness, and individual rationality, are incompatible for general multi-object problems (Konishi et al. (2001), Todo et al. (2014)). [^3] Ma (1994) shows that these properties characterize the celebrated Top Trading Cycles rule in the Shapley-Scarf model (see also Sethuraman (2016)). Based on this trade-off, we search for a plausible rule by relaxing these requirements. Our main result is, loosely speaking, that the generalized Top Trading Cycles rule (TTC) is characterized by Pareto efficiency (or individual-good efficiency), balancedness, the weak endowment lower bound, and truncation-proofness (or drop strategy-proofness).
Our first property, balancedness, says that a rule should assign the same number of objects to each agent as the initial endowment. Balancedness is an inviolable constraint in many practical reallocation problems. For example, in the context of shift reallocation among workers, balancedness is often imposed so that staff can meet the necessary training requirements or because it is stipulated in employment contracts (Manjunath and Westkamp, 2021). Balancedness is also a requirement in student exchange programs such as the Erasmus Student Exchange Program in Europe and “tuition exchange programs” in the United States (Dur and Ünver, 2019). In the absence of such constraints, balancedness plays a mostly technical role. To the extent that all balanced allocations can be obtained from the initial allocation by executing a series of simple “single-object exchanges,” the restriction to balanced rules vastly simplifies the planner’s problem. Furthermore, balancedness implies a form of “endowment monotonicity,” as it guarantees that agents are rewarded with more objects whenever they bring more objects to the market.
Our participation guarantee, the weak endowment lower bound, requires that no agent ever receives an object which is worse than every object in her endowment. [^4] Here, by “weak” we mean that it is weaker than Altuntaş et al.’s (2023) strong endowment lower bound. The weak endowment lower bound is important at a technical level, as it imposes discipline directly on the set of objects that an agent can receive in any assigned bundle. In contrast, individual rationality allows for an agent to be assigned any object, provided that the entire bundle of assigned objects is sufficiently desirable. Since the two properties coincide for single-unit problems, the weak endowment lower bound can be viewed as one possible extension of individual rationality to multi-object problems. [^5] As we see in the proofs of Theorems 3.2 and 3.5, the weak endowment lower bound plays a role that is analogous to that of individual rationality in characterizations of TTC for single-object problems. Aside from the technical reasons, there are a couple of practical reasons for considering the weak endowment lower bound as an alternative desideratum. First, in contrast with individual rationality, which is based on the agents’ actual preferences over bundles, the weak endowment lower bound pertains only to the agents’ rankings over individual objects. Thus, a social planner may seek to simplify the complex problem of achieving full individual rationality by focusing on only a few of its salient features. Second, the weak endowment lower bound allows agents to explicitly veto some of the other agents’ objects, and the right to veto can be viewed as a minimal requirement. [^6] There are some settings in which agents should not expect reallocation to be Pareto-improving. For example, in professional sports leagues, some players at team $i$ may express their desire to be traded to another team. The resulting reallocation need not benefit team $i$, yet the team should not be forced to sign “unacceptable players.” Proposition 2.12 demonstrates that, in the presence of our other properties, the weak endowment lower bound implies individual rationality.
Finally, motivated by evidence that agents may only consider manipulations that are similar to their true preferences (Mennle et al., 2015), we consider incentive properties that defend against two simple manipulation heuristics. Drop strategy-proofness defends against drop strategies, whereby an agent misrepresents her ranking over individual objects by dropping one object to the bottom. [^7] Drop strategy-proofness was introduced by Altuntaş et al. (2023) to give a characterization of generalized Top Trading Cycles on the lexicographic domain (Corollary 3.4 of the present paper). Truncation-proofness defends against truncation strategies, whereby an agent truncates her list of “acceptable objects” but truthfully reports her ranking over the objects she deems acceptable. [^8] Though we consider only monotone preferences (i.e., all objects are goods), we interpret an agent’s set of “acceptable objects” as the set of objects that are weakly preferred to some object in her endowment. Together with our participation guarantees, both incentive properties ensure that agents cannot benefit by vetoing some of the other agents’ objects. Truncation strategies are among the most obvious forms of manipulation; they are “monotonic manipulations” in the sense of Troyan and Morrill (2020) and, in many cases, they are the only manipulations that are profitable (Roth and Rothblum (1999); Ehlers (2008); Kojima and Pathak (2009), Kojima (2013)).
As mentioned previously, we are primarily interested in individual-good-based rules. One setting where the focus on individual-good-based rules is not a substantive restriction is when agents have lexicographic preferences over bundles of objects. That is, each agent’s preferences over bundles are completely determined by her preferences over individual objects as follows: for any two distinct bundles, $X$ and $Y$, bundle $X$ is preferred to $Y$ if and only if the best object that belongs to exactly one of the two bundles is in $X$999 For example, students at a university may want to reschedule their courses by exchanging their assigned slots, and students may lexicographically prefer certain timings, e.g., a “party boy” strongly prefers not to attend the class on Friday.
On the lexicographic domain, the generalized Top Trading Cycles rule (TTC) is characterized by Pareto efficiency, balancedness, the weak endowment lower bound, and truncation-proofness (or drop strategy-proofness) (Theorems 3.2 and 3.3). These results generalize a recent innovation of Altuntaş et al. (2023). Our results also yield a new characterization for the Shapley-Scarf model, sharpening results due to Ma (1994). Specifically, TTC is characterized by Pareto efficiency, individual rationality, and truncation-proofness (Theorem 3.5).
On the more general domain of responsive preferences, the restriction to individual-good-based rules is a substantive one. With only limited information about the agents’ preferences, it is impossible to achieve full Pareto efficiency in conjunction with our other properties. However, TTC satisfies a weaker efficiency notion, individual-good efficiency, which requires that agents cannot Pareto-improve upon the allocation by executing some single-object exchange. Though TTC is not drop strategy-proof in this setting, it is truncation-proof, and our main characterization extends straightforwardly to this more general domain. In particular, TTC is the unique individual-good-based rule that satisfies individual-good efficiency, balancedness, the weak endowment lower bound, and truncation-proofness (Theorem 4.4). Moreover, on this domain, we obtain another characterization by replacing the weak endowment lower bound with individual rationality (Corollary 4.6).
We also consider a richer reporting language that allows agents to express conditionally lexicographic preferences over bundles of objects. That is, for any disjoint bundles $X$ and $Y$ such that $X$ is nonempty, there is an object $o\in X$ such that agent $i$ lexicographically prefers adding $o$ to $Y$ rather than adding any other bundle $Z\subseteq X\setminus\{o\}$. Conditionally lexicographic preferences have a compact representation in terms of “lexicographic preference trees,” which makes them appealing from an implementation perspective. They are more flexible than purely lexicographic preferences, and they are economically interesting as they can capture some complementarities. The Augmented Top Trading Cycles rule (ATTC), first introduced by Fujita et al. (2018), is the natural extension of TTC to the conditionally lexicographic domain. We show that ATTC is characterized by Pareto efficiency, balancedness, and generalized versions of the weak endowment lower bound and drop strategy-proofness (Theorem 4.13).
Finally, we show that the lexicographic and conditionally lexicographic domains are quite special. The lexicographic domain is a maximal domain of responsive preferences on which the two efficiency notions, Pareto efficiency and individual-good efficiency, are equivalent. More generally, the conditionally lexicographic domain is a maximal domain of monotone preferences on which the two notions coincide.
1.2. Related literature
Lexicographic preferences are common in other studies focusing on multi-unit demand (e.g., Altuntaş et al. (2023); Biró and Csáji (2024)). From a theoretical point of view, the lexicographic domain admits several positive results where larger preference domains yield impossibilities. For instance, the core can be empty under additive preferences (Konishi et al., 2001), while it is always nonempty under (conditionally) lexicographic preferences (Sikdar et al., 2017; Fujita et al., 2018).
The paper closest to ours is Altuntaş et al. (2023). They showed that TTC is drop strategy-proof on the lexicographic domain, and our proof of Proposition 4.12 borrows their ingenuity. As mentioned previously, our characterization on the lexicographic domain generalizes their result. Unlike Altuntaş et al.’s (2023) proof, the proof of our characterization is not completely constructive. We borrow the notions of size and similarity from Sethuraman (2016) and Ekici (2023), respectively, and we give a proof by minimum counter-example. We note, however, that our “minimality condition” differs from the conditions used by Sethuraman (2016) and Ekici (2023).
Biró et al. (2022) study the closely related multi-unit Shapley-Scarf model, in which each agent is endowed with a set of agent-specific homogeneous objects. They first characterize the Circulation Top Trading Cycles (cTTC) rule on the lexicographic domain, and then extend their characterization to the responsive domain via the “individual-good-based” property. We use similar ideas to extend our main characterization from the lexicographic domain to the responsive domain.
In our model, one difficulty is the trade-off between Pareto efficiency, strategy-proofness, and individual rationality. When agents consume more than one object, the combination of Pareto efficiency and strategy-proofness essentially results in serially dictatorial rules—or slight generalizations of them (Pápai, 2001; Klaus and Miyagawa, 2002; Monte and Tumennasan, 2015; Todo et al., 2014). Several studies insist on strategy-proofness while relaxing Pareto efficiency. For example, Pápai (2003) shows that, on the domain of responsive preferences, the Segmented Trading Cycles rules are characterized by strategy-proofness, non-bossiness, trade sovereignty, and strong individual rationality (see also Pápai (2007), Anno and Kurino (2016), and Nesterov (2017)). In our paper, we keep Pareto efficiency (or individual-good efficiency) but weaken strategy-proofness to truncation-proofness and drop strategy-proofness. There are also many other relaxed forms of strategy-proofness, e.g., rank monotonicity (Chen and Zhao, 2021), partial strategy-proofness (Mennle and Seuken, 2021), truncation-invariance (Chen et al., 2024), weak truncation robustness (Hashimoto et al., 2014), and convex strategy-proofness (Balbuzanov, 2016).
Several papers focus on models that are only slightly different from ours. Feng et al. (2024); Feng (2023) consider the multiple-type Shapley-Scarf model, in which objects are partitioned by different types, and each agent owns one object of each type. Manjunath and Westkamp (2021) show that, for reallocation problems with trichotomous preferences, Pareto efficiency, strategy-proofness, and individual rationality are compatible. Echenique et al. (2024) studies the existence of stable allocations by extending the TTC algorithm.
Aside from the axiomatic approach, some papers consider the computational complexity of manipulation in TTC (Fujita et al., 2018; Phan and Purcell, 2022). Those studies emphasize that relaxing strategy-proofness for our model does make sense as finding a beneficial manipulation is computationally difficult in general.
1.3. Organization
The rest of our paper is organized as follows. In Section 2 we introduce our model and some properties of allocation rules, and we give a description of generalized TTC. Our main results for the lexicographic domain are stated in Section 3. In Subsection 3.1, we state and prove a corollary for the Shapley–Scarf model. The proof of Theorem 3.5 provides intuition for the general proof strategy. In Section 4.1 we introduce the domain of responsive preferences, and we give characterizations of TTC on this more general domain. Section 4.2 introduces the domain of conditionally lexicographic preferences. We discuss the extension of our properties to this domain, and give a characterization of ATTC (Theorem 4.13). In Section 5, we study the maximal domains on which our two efficiency properties are equivalent. Section 6 concludes with some final remarks. Appendix A contains all proofs that were omitted from the main text, and we demonstrate the independence of our properties in Appendix B.1. Several other important examples are relegated to Appendix B.2.
2. Preliminaries
2.1. Model
Let $N=\left\{1,2,\dots,n\right\}$ be a finite set of $n\geq 2$ agents. Let $O$ be a finite set of heterogeneous and indivisible objects such that $|O|\geq n$. A bundle is a subset of $O$. Let $2^{O}$ denote the set of bundles. A generic element of $O$ is denoted by a lowercase letter, e.g., $x,y,z$, and a generic element of $2^{O}$ is typically denoted by an uppercase letter, e.g., $X,Y,Z$.
An allocation is a function $\mu:N\to 2^{O}$ such that (i) for all $i\in N$, $\mu(i)\neq\emptyset$, (ii) for all $i,j\in N$, $i\neq j$ implies $\mu\left(i\right)\cap\mu\left(j\right)=\emptyset$, and (iii) $\bigcup_{i\in N}\mu\left(i\right)=O$. We shall represent each allocation $\mu$ as a profile $\left(\mu_{i}\right)_{i\in N}$ of nonempty and pairwise disjoint bundles with $\bigcup_{i\in N}\mu_{i}=O$. For each $i\in N$, $\mu_{i}$ denotes agent $i$’s assignment at $\mu$. Let ${\mathcal{M}}$ denote the set of allocations.
Let $\omega=\left(\omega_{i}\right)_{i\in N}\in\mathcal{M}$ denote the initial allocation. For each $i\in N$, bundle $\omega_{i}$ is agent $i$’s endowment. With some abuse of notation, we denote the initial owner of object $o$ by $\omega^{-1}\left(o\right)$.
Each agent $i$ has a strict preference relation $P_{i}$ over the set of all bundles. If agent $i$ prefers bundle $X$ to bundle $Y$, then we write $X\mathbin{P_{i}}Y$. Let $R_{i}$ denote the at least as good as relation associated with $P_{i}$. That is, $X\mathbin{R_{i}}Y$ means that ($X\mathbin{P_{i}}Y$ or $X=Y$). [^10] Formally, $R_{i}$ is a complete, transitive, and antisymmetric binary relation on $2^{O}$, and $P_{i}$ is the asymmetric part of $R_{i}$. Let $P=\left(P_{i}\right)_{i\in N}$ denote a preference profile, where $P_{i}$ is agent $i$’s preference relation.
When there is no danger of confusion, we abuse notation and denote singletons $\left\{x\right\}$ by $x$. For example, $xP_{i}y$ means $\left\{x\right\}P_{i}\left\{y\right\}$, and $X\cup x$ means $X\cup\{x\}$.
Given a preference relation $P_{i}$ on $2^{O}$ and a nonempty bundle $X\in 2^{O}$, let $\top_{P_{i}}\left(X\right)$ denote the top-ranked object in $X$, i.e., $\top_{P_{i}}\left(X\right)=x$, where $x\in X$ and $x\mathbin{R_{i}}y$ for all $y\in X$. Similarly, let $\bot_{P_{i}}\left(X\right)$ denote the bottom-ranked object in $X$, i.e., $\bot_{P_{i}}\left(X\right)=x$, where $x\in X$ and $y\mathbin{R_{i}}x$ for all $y\in X$.
Given a preference profile $P=\left(P_{i}\right)_{i\in N}$ and a relation $P^{\prime}_{i}$ for agent $i$, $\left(P^{\prime}_{i},P_{-i}\right)$ denotes the preference profile in which agent $i$’s preference relation is $P^{\prime}_{i}$ and, for each agent $j\in N\backslash\left\{i\right\}$, agent $j$’s preference relation is $P_{j}$.
An object reallocation problem (a problem in short) is a triple $(N,\omega,P)$. Since $(N,\omega)$ remains fixed throughout, we will simply denote a problem by $P$. We will assume that each agent’s preference relation belongs to some domain $\mathcal{P}$, i.e., $\mathcal{P}$ is a subset of all strict preference relations on $2^{O}$. Thus, the preference profile domain $\mathcal{P}^{N}$ also denotes the set of all problems.
Let $\mathcal{P}$ denote a domain of preferences. A rule (on $\mathcal{P}$) is a function $\varphi:{\mathcal{P}}^{N}\to{\mathcal{M}}$ that associates to each profile $P$ an allocation $\varphi\left(P\right)$. For each $i\in N$, $\varphi_{i}\left(P\right)$ denotes agent $i$’s assignment at $\varphi\left(P\right)$.
2.2. Lexicographic preferences
In what follows, we assume that agents’ preferences over $2^{O}$ are lexicographic. That is, when facing two distinct bundles $X$ and $Y$, if agent $i$ prefers the top-ranked object in $X$ to that in $Y$, then $i$ also prefers $X$ to $Y$. If these top-ranked objects are the same, then $i$ compares the second top-ranked object in each bundle, and so on. Finally, if $Y$ is a subset of $X$, then $i$ prefers $X$ to $Y$. Formally, for any two distinct bundles $X,Y\in 2^{O}$, $X\mathbin{P_{i}}Y$ if
(1) there is an object $\text{there is an object }o\in X\setminus Y\text{ such that }\top_{P_{i}}\left((Y\setminus X)\cup o\right)=o.$
Let $\mathcal{L}$ denote the set of all lexicographic preference relations on $2^{O}$.
Note that the set of lexicographic preference relations over $2^{O}$ is in bijection with the set of linear orders over $O$. Therefore, we will identify lexicographic preference relations with their associated linear orders, e.g., $P_{i}:x_{1},x_{2},\dots,x_{m}$ means that $P_{i}\in\mathcal{L}$ and $x_{1}\mathbin{P_{i}}x_{2}\mathbin{P_{i}}\cdots\mathbin{P_{i}}x_{m}$, with the understanding that all remaining relations among bundles in $2^{O}$ can be deduced from (1).
In Section 4, we relax the assumption of lexicographic preferences. There we consider two more general domains. First we consider the domain $\mathcal{R}$ of responsive preferences. The responsive domain captures substitutability between the objects, and it is well-studied in the literature on matching problems (Roth, 1985). We also consider the conditionally lexicographic domain $\mathcal{CL}$, which is a more general domain exhibiting complementarity that has been recently studied by computer science researchers (Booth et al., 2010). For now we note that $\mathcal{R}\cap\mathcal{CL}=\mathcal{L}$.
2.3. Rules and their properties
In what follows, $\mathcal{P}\coloneqq\mathcal{L}$ denotes the lexicographic preference domain. We now introduce and discuss our main properties for allocations and rules defined on $\mathcal{P}$ [^11] As we shall see in Section 4, many of these properties extend to the responsive domain $\mathcal{R}$ without any modification. However, some of the properties must be adapted in order to deal with the complementarities that arise on $\mathcal{CL}$.
First, we consider one well-known efficiency criterion. Let $P\in{\mathcal{P}}^{N}$. An allocation $\overline{\mu}$ Pareto-dominates another allocation $\mu$ at $P$ if (i) for all $i\in N$, $\overline{\mu}_{i}R_{i}\mu_{i}$, and (ii) for some $i\in N$, $\overline{\mu}_{i}P_{i}\mu_{i}$. An allocation $\mu$ is Pareto efficient at $P$ if it is not Pareto-dominated at $P$ by any other allocation.
Definition 2.1 (Pareto efficiency).
A rule $\varphi$ is Pareto efficient if, for each $P\in\mathcal{P}^{N}$, $\varphi(P)$ is Pareto efficient at $P$.
Next, we introduce a weak version of efficiency, called individual-good efficiency (or ig-efficiency in short), which precludes the existence of a single-object exchange that improves the welfare of all of its participants. Given an allocation $\mu$ and a preference profile $P$, a Pareto-improving single-object exchange at $\mu$ is a cycle $C=\left(i_{1},o_{i_{2}},i_{i_{2}},o_{i_{3}},\dots,i_{k},o_{i_{k+1}},i_{k+1}=i_{1}\right)$ of agents and objects such that, for all $\ell\in\left\{1,\dots,k\right\}$,
(2) $i_{\ell}\in N,\quad o_{\ell}\in\mu_{i_{\ell}},\quad\text{and}\quad(\mu_{i_{\ell}}\cup o_{i_{\ell+1}})\setminus o_{i_{\ell}}\mathbin{P_{i_{\ell}}}\mu_{i_{\ell}}.$
An allocation $\mu$ is ig-efficient at $P$ if there does not exist a Pareto-improving single-object exchange at $\mu$ [^12] The terminology “ig-efficiency” is borrowed from Biró et al. (2022). Similar properties are studied in Aziz et al. (2019), Caspari (2020), and Coreno and Balbuzanov (2022)).
Definition 2.2 (ig-efficiency).
A rule $\varphi$ is ig-efficient if, for each $P\in{\mathcal{P}}^{N}$, $\varphi(P)$ is ig-efficient at $P$.
Clearly, Pareto efficiency implies ig-efficiency on an arbitrary domain of preferences. On the lexicographic domain, Pareto efficiency and ig-efficiency are equivalent (Aziz et al., 2019). (We generalize this result in Proposition 5.2 of Section 5.)
Our next property requires that each agent is always assigned exactly as many objects as they initially own. An allocation $\mu$ is called balanced if, for each $i\in N$, $|\mu_{i}|=|\omega_{i}|$.
Definition 2.3 (Balancedness).
A rule $\varphi$ is balanced if, for each $P\in\mathcal{P}^{N}$, $\varphi(P)$ is balanced.
Balancedness is an inviolable constraint in many practical reallocation problems, such as shift reallocation among workers (Manjunath and Westkamp, 2021) and “tuition exchange programs” in the United States (Dur and Ünver, 2019). In the absence of such constraints, the restriction to balanced rules helps to simplify the planner’s problem, since all balanced allocations can be obtained from the initial allocation by executing a series of simple “single-object exchanges.” [^13] To see this, let $\mu$ be a balanced allocation. Let $I\coloneqq\{i\in N\mid\mu_{i}\neq\omega_{i}\}$ and $X\coloneqq\bigcup_{i\in N}(\mu_{i}\setminus\omega_{i})$. Construct a digraph on $I\cup X$ as follows. Each agent $i\in I$ points to her best object in $X\cap(\mu_{i}\setminus\omega_{i})$, and each object $x\in X$ points to its owner. There exists a cycle, say $C$. Execute the corresponding exchange. Letting $O(C)$ denote the objects on $C$, redefine $X$ as $X\setminus O(C)$ and repeat., [^14] Note that these exchanges need not be Pareto-improving. Furthermore, balancedness implies a form of “endowment monotonicity,” as it guarantees that agents are rewarded with more objects (but not necessarily better objects) whenever they bring more objects to the market.
We now introduce two related participation requirements. The first property, individual rationality, is well-studied; it guarantees that no agent is harmed by the reallocation.
Definition 2.4 (Individual rationality).
A rule $\varphi$ satisfies individual rationality if, for each $P\in\mathcal{P}^{N}$ and each $i\in N$, $\varphi_{i}(P)\mathbin{R_{i}}\omega_{i}$.
The second property, the strong endowment lower bound, was introduced by Altuntaş et al. (2023). It guarantees that each agent is assigned a bundle that “pairwise dominates” her initial endowment, in the following sense.
Definition 2.5 (Strong endowment lower bound).
A rule $\varphi$ satisfies the strong endowment lower bound if, for each $P\in\mathcal{P}^{N}$ and each $i\in N$, there is a bijection $\sigma:\omega_{i}~{}\to~{}\varphi_{i}(P)$ such that for each $o\in\omega_{i}$, $\sigma(o)\mathbin{R_{i}}o$. Note that $\{\sigma(o)\mid o\in\omega_{i}\}=\varphi_{i}(P)$.
The strong endowment lower bound ensures that, for each agent $i\in N$, each object $x\in\omega_{i}$ is exchanged for something weakly better than $x$. Our main property, the weak endowment lower bound, relaxes this significantly. It requires that each agent is only assigned objects that are weakly better than her worst endowed object.
Let $P\in{\mathcal{P}}^{N}$. For each agent $i$ and each preference relation $P_{i}$, let $w_{i}(P_{i})\coloneqq\bot_{P_{i}}\left(\omega_{i}\right)$ be agent $i$’s worst endowed object at $P_{i}$. An allocation $\mu$ satisfies the weak endowment lower bound at $P$ if, for each agent $i\in N$, and each of her assigned objects $o\in\mu_{i}$, $o\mathbin{R_{i}}w_{i}(P_{i})$.
Definition 2.6 (Weak endowment lower bound).
A rule $\varphi$ satisfies the weak endowment lower bound if, for each $P\in\mathcal{P}^{N}$, $\varphi(P)$ satisfies the weak endowment lower bound at $P$.
As we mentioned in the introduction, the weak endowment lower bound is economically relevant as it (i) gives agents the right to veto some objects, and (ii) helps the social planner to simplify the problem—and this simplicity is without loss of generality because, together with our other properties, the weak endowment lower bound also corresponds to individual rationality (see Proposition 2.12 below).
Remark 1.
The relation between the weak endowment lower bound, individual rationality, the strong endowment lower bound, and balancedness is as follows.
  • The strong endowment lower bound implies each of the other three properties, but the converse is false.
  • Even in the presence of balancedness, the weak endowment lower bound and individual rationality are logically independent.
The independence is demonstrated in Example B.1 in Appendix B.2. $\diamond$
The next property, strategy-proofness, is one incentive property that is most frequently used in the literature on mechanism design. It says that no agent can benefit from misrepresenting her preferences.
Definition 2.7 (Strategy-proofness).
A rule $\varphi$ is strategy-proof if, for each $P\in\mathcal{P}^{N}$, each $i\in N$, and each $P^{\prime}_{i}\in\mathcal{P}$, $\varphi_{i}(P)\mathbin{R_{i}}\varphi_{i}(P^{\prime}_{i},P_{-i})$. In other words, at any $P$, no agent $i$ can manipulate $\varphi$ by misreporting any preference relation.
2.3.1. Heuristic manipulations
A strategy-proof rule $\varphi$ prohibits each agent $i$ from manipulating the rule at any problem $P$ by misreporting any preference relation in $\mathcal{P}$. A natural weakening of strategy-proofness is to prohibit each agent $i$ from manipulating the rule at any problem $P$ by reporting any preference relation in some subset $\mathcal{S}(P_{i})$ of $\mathcal{P}$. The restricted set of manipulation strategies, $\mathcal{S}(P_{i})$, could consist of misrepresentations which are, in some sense, more robust, or whose consequences are simpler to understand.
To this end, we define three simple manipulation heuristics: subset drop strategies, drop strategies, , and truncation strategies. When agents only need to report their preferences over individual objects (such as when they have lexicographic preferences or when a rule is “individual-good-based” as in Section 4.1), these heuristics are “intuitively appealing and simple for agents to implement” (Castillo and Dianat, 2016). We also define the associated incentive properties, namely drop strategy-proofness and truncation-proofness. Let $P_{i}\in\mathcal{P}$ $(=\mathcal{L})$.
Drop strategies and truncation strategies are both special cases of subset drop strategies. We say that $P^{\prime}_{i}\in\mathcal{P}$ is a subset drop strategy for $P_{i}$ if there exists $X\subseteq O\backslash\omega_{i}$ such that:
  1. (1)
    for all $x\in X$ and $y\in O\backslash X$, $y\mathbin{P^{\prime}_{i}}x$; and
  2. (2)
    for all $x,y\in X$, $x\mathbin{P^{\prime}_{i}}y$ if and only if $x\mathbin{P_{i}}y$; and
  3. (3)
    for all $x,y\in O\backslash X$, $x\mathbin{P^{\prime}_{i}}y$ if and only if $x\mathbin{P_{i}}y$.
In this case, we say that $P^{\prime}_{i}$ is obtained from $P_{i}$ by dropping the subset $X$.
We say that $P^{\prime}_{i}\in\mathcal{P}$ is a drop strategy for $P_{i}$ if it is obtained from $P_{i}$ by dropping a singleton subset $\{x\}$ of $O\setminus\omega_{i}$. In this case, we say that $P^{\prime}_{i}$ is obtained from $P_{i}$ by dropping object $x$. Let $\mathcal{D}(P_{i})$ denote the set of all drop strategies for $P_{i}$.
Finally, $P^{\prime}_{i}\in\mathcal{P}$ is a truncation strategy for $P_{i}$ if it is obtained from $P_{i}$ by dropping a “tail subset” $X$ of $O\setminus\omega_{i}$. [^15] By a “tail subset” of $O\setminus\omega_{i}$ we mean a set $X\subseteq O\setminus\omega_{i}$ such that if $x\in X$, $y\in O\setminus\omega_{i}$ and $x\mathbin{R_{i}}y$, then $y\in X$. Formally, in addition to (1)-(3) above, the following holds:
  1. (4)
    $X=O\backslash\omega_{i}$ or there exists $x\in O\backslash\omega_{i}$ such that $X=\{o\in O\backslash\omega_{i}\mid x\mathbin{P_{i}}o\}$.
If $X=\{o\in O\backslash\omega_{i}\mid x\mathbin{P_{i}}o\}$, then $P^{\prime}_{i}$ is called a truncation of $P_{i}$ at object $x$. Let $\mathcal{T}(P_{i})$ denote the set of all truncation strategies for $P_{i}$.
An important distinction between truncation strategies and drop strategies is that, unlike every drop strategy for $P_{i}$, every truncation strategy $P^{\prime}_{i}$ of $P_{i}$ agrees with $P_{i}$ on $O\backslash\omega_{i}$, i.e., for all $x,y\in O\backslash\omega_{i}$, $x\mathbin{P_{i}}y$ if and only if $x\mathbin{P^{\prime}_{i}}y$. In particular, an agent cannot push an arbitrary object to the top of her preference list by repeatedly applying truncation strategies.
A rule is drop strategy-proof if it is not manipulable via drop strategies, and truncation-proof if it is not manipulable via truncation strategies.
Definition 2.8 (Drop strategy-proofness).
A rule $\varphi$ is drop strategy-proof if, for each $P\in\mathcal{P}^{N}$, each $i\in N$, and each $P^{\prime}_{i}\in\mathcal{D}(P_{i})$, $\varphi_{i}(P)\mathbin{R_{i}}\varphi_{i}(P^{\prime}_{i},P_{-i})$. In other words, at any $P$, no agent $i$ can manipulate $\varphi$ by reporting any drop strategy. [^16] For the case of lexicographic preferences, our definition of drop strategy-proofness is precisely the weakening of drop strategy-proofness and subset total drop strategy-proofness discussed in Footnote 31 of Altuntaş et al. (2023).
Definition 2.9 (Truncation-proofness).
A rule $\varphi$ is truncation-proof if, for each $P\in{\mathcal{P}}^{N}$, each $i\in N$, and each $P^{\prime}_{i}\in{\mathcal{T}}\left(P_{i}\right)$, $\varphi_{i}\left(P\right)\mathbin{R_{i}}\varphi_{i}\left(P^{\prime}_{i},P_{-i}\right)$. In other words, at any $P$, no agent $i$ can manipulate $\varphi$ by reporting any truncation strategy.
Remark 2.
Truncation strategies are well-studied in the literature on matching theory (Roth and Vande Vate, 1991; Roth and Rothblum, 1999; Ehlers, 2008; Castillo and Dianat, 2016). Intuitively, truncating means by shifting the outside option up, thereby truncating the “tail” of acceptable objects. This notion can be easily extended to the Shapley-Scarf model by replacing the outside option with the agent’s single endowment (see the definition of truncation strategy in Subsection 3.1). However, the difficulty arises when an agent has more than one endowed object, as now the definitions of acceptable objects and unacceptable objects become unclear. Of course, for an agent $i$, if an object is better than her best endowed object $\top_{P_{i}}\left(\omega_{i}\right)$, then it is certainly acceptable; and if it is worse than her worst endowed object $\bot_{P_{i}}\left(\omega_{i}\right)$, then it is unacceptable. But what about objects that are ranked between these two extremes of endowed objects? Motivated by drop strategies in Kojima and Pathak (2009), [^17] Drop strategies defined in Kojima and Pathak (2009) are a generalization of truncation strategies, and Kojima and Pathak (2009) show that drop strategies are exhaustive when searching for manipulations for agents with a capacity greater than one. The same is not true in our model. here we consider an extreme way to model truncation strategies: truncating a preference relation $P_{i}$ at an object $x$ corresponds to vetoing all objects worse than $x$ (by declaring them worse than her worst endowed object $\bot_{P_{i}}(\omega_{i})$). $\diamond$
Example 2.10 (Drop strategies and truncation strategies).
Suppose $\omega_{i}=\{x,y\}$ and $P_{i}:a,b,\underline{x},c,d,\underline{y},e$ (endowments are underlined). Then:
  • $P^{1}_{i}:b,\underline{x},c,d,\underline{y},e,a$ is a drop strategy for $P_{i}$; it is obtained by dropping object $a$. Note that $P^{1}_{i}$ is not a truncation strategy for $P_{i}$.
  • $P^{2}_{i}:a,b,\underline{x},c,\underline{y},d,e$ is a truncation strategy for $P_{i}$; it is obtained from $P_{i}$ by dropping the subset $\{d,e\}$ (or subsequently dropping $d$ then $e$).
  • $P^{3}_{i}:a,\underline{x},\underline{y},b,c,d,e,$ is a truncation strategy for $P_{i}$; it is obtained from $P_{i}$ by dropping the subset $\{b,c,d,e\}$ (or subsequently dropping $b,c,d$, then $e$). $\diamond$
We conclude this subsection by stating some useful relations between our properties on the lexicographic domain.
Proposition 2.11.
On the lexicographic domain, if a rule $\varphi$ satisfies drop strategy-proofness and the weak endowment lower bound, then it is truncation-proof.
Example B.2 in Appendix B.2 demonstrates that truncation-proofness and the weak endowment lower bound do not jointly imply drop strategy-proofness. Example B.3 shows that truncation-proofness is not implied by drop strategy-proofness alone.
Proposition 2.12.
On the lexicographic domain, if a rule $\varphi$ satisfies balancedness, the weak endowment lower bound, and truncation-proofness, then it is individually rational.
The proofs are given in Appendix A.
2.4. Top Trading Cycles
A cycle is a sequence
$C=\left(i_{1},o_{i_{2}},i_{2},o_{i_{3}},\dots,i_{k-1},o_{i_{k}},i_{k},o_{i_{k+1}},i_{k+1}=i_{1}\right)$
of $k\geq 1$ distinct agents and objects such that each agent on the cycle points to her favorite object among some subset of $O$ and each object points to its owner. The set of agents on $C$ is denoted by $N\left(C\right)=\left\{i_{1},i_{2},\dots,i_{k}\right\}$. Similarly, the set of objects on $C$ is denoted by $O\left(C\right)$.
An allocation $\mu$ executes the cycle $C$ if $\mu$ assigns each agent in $N\left(C\right)$ the object to which she points in $C$, i.e., for each $i_{\ell}\in N\left(C\right)$, $o_{i_{\ell+1}}\in\mu_{i_{\ell}}$.
We consider a generalized version of the Top Trading Cycles (TTC) procedure. At each step, every agent points to her favorite remaining object, and every remaining object points to its owner. At least one cycle obtains.
Our proofs rely on a variant of TTC that executes only a single cycle at each step. If multiple cycles obtain, then we execute only the smallest cycle, i.e., the cycle containing the agent identified by the smallest natural number. That is, we assign each agent on the smallest cycle the object to which she points, then clear all objects on the cycle. It is well-known that this procedure is equivalent to the variant of TTC that, at each step, executes all prevailing cycles.
Formally, for each profile $P\in{\mathcal{P}}^{N}$, TTC selects the allocation $\varphi^{\text{TTC}}\left(P\right)$ determined by the following procedure, which we call $\text{TTC}\left(P\right)$.
Algorithm: $\text{TTC}\left(P\right)$
Input:: A preference profile $P\in{\mathcal{P}}^{N}$. Output:: An allocation $\varphi^{\text{TTC}}\left(P\right)$. Step 0.: Define $\mu^{0}\coloneqq\left(\emptyset\right)_{i\in N}$ and $O^{1}\coloneqq O$. Step $t\geq 1$.: Construct a bipartite directed graph with independent vertex sets $N$ and $O^{t}$. For each agent $i\in N$, there is a directed edge from $i$ to agent $i$’s most-preferred object in $O^{t}$, $\top_{P_{i}}\left(O^{t}\right)$. For each object $o\in O^{t}$, there is a directed edge from $o$ to its owner, $\omega^{-1}\left(o\right)$. At least one cycle exists. Let $\mathcal{C}_{t}(P)$ denote the set of cycles that obtain at step $t$, and let $C_{t}(P)$ be the smallest cycle in $\mathcal{C}_{t}(P)$, i.e., $\min N(C_{t}(P))\leq\min N(C)$ for all $C\in\mathcal{C}_{t}(P)$. Assign each agent $i\in N(C_{t}(P))$ the object to which she points on $C_{t}(P)$, namely $\top_{P_{i}}\left(O^{t}\right)$; that is, let $\mu^{t}=\left(\mu_{i}^{t}\right)_{i\in N}$ be the allocation such that (i) for all $i\in N(C_{t}(P))$, $\mu_{i}^{t}=\mu_{i}^{t-1}\cup\left\{\top_{P_{i}}\left(O^{t}\right)\right\}$, and (ii) for all $i\in N\backslash N(C_{t}(P))$, $\mu_{i}^{t}=\mu_{i}^{t-1}$. Let $O^{t+1}\coloneqq O^{t}\backslash O(C_{t}(P))$ denote the set of objects remaining at Step $t+1$. If $O^{t+1}\neq\emptyset$, then proceed to Step $t+1$; otherwise, proceed to Termination. Termination.: Because $O$ is finite, and $\left|O^{1}\right|>\left|O^{2}\right|>\cdots>\left|O^{t}\right|$, the procedure terminates at some step $T$. Return the allocation $\varphi^{\text{TTC}}\left(P\right)\coloneqq\mu^{T}$.
In what follows, we focus on the rule that always selects the TTC allocation $\varphi^{\text{TTC}}\left(P\right)$ for each profile $P\in\mathcal{P}^{N}$. We call such a rule the generalized Top Trading Cycles rule, or TTC in short.
3. A characterization of TTC
Throughout this section we assume that agents have lexicographic preferences, i.e., $\mathcal{P}=\mathcal{L}$. We first show that TTC satisfies all of our properties.
Proposition 3.1.
On the lexicographic domain, TTC satisfies Pareto efficiency, balancedness, the weak endowment lower bound, drop strategy-proofness, and truncation-proofness.
The proof of Proposition 3.1 follows immediately from Altuntaş et al. (2023, Theorem 1) and Proposition 2.11.181818Altuntaş et al. (2023) show that TTC satisfies a stronger version of drop strategy-proofness which, in addition, requires that no agent can manipulate by dropping an object in her own endowment. We now state our main result, a characterization of TTC.
Theorem 3.2.
On the lexicographic domain, a rule satisfies
  • Pareto efficiency,
  • balancedness,
  • the weak endowment lower bound, and
  • truncation-proofness
if and only if it is TTC.
Our second characterization is an immediate corollary of Theorem 3.2 and Propositions 2.11 and 3.1.191919In footnote 37 we explain how to give a direct proof of this result by instead invoking drop strategy-proofness in the proof of Theorem 3.2.
Theorem 3.3.
On the lexicographic domain, a rule satisfies Pareto efficiency, balancedness, the weak endowment lower bound, and drop strategy-proofness if and only if it is TTC.
The proof of uniqueness is given in Appendix A. Before stating the intuition for our proof strategy, we make three observations.
First, it is perhaps surprising that individual rationality does not appear in Theorem 3.2, and we do not use it for our proof. Second, together with Remark 1, Theorem 3.2 implies the following corollary.
Corollary 3.4 (Altuntaş et al., 2023).
On the lexicographic domain, a rule satisfies Pareto efficiency, the strong endowment lower bound, and drop strategy-proofness if and only if it is TTC.
Third, our characterization of TTC is also valid for the the Shapley-Scarf model, the special case of our model in which each agent is endowed with—and has the need for—a single object. Given the significance of this result, we provide a detailed statement and proof in what follows. We begin by formally reintroducing the Shapley-Scarf model.
3.1. The Shapley-Scarf model
Recall that $N=\{1,\ldots,n\}$. Now, let $O=\{o_{1},\ldots,o_{n}\}$, and for each $i\in N$, $|\omega_{i}|=1$. Thus, it is without loss of generality to assume that $\omega_{i}=o_{i}$ for each $i\in N$. Moreover, in this model, each agent consumes exactly one object. Thus, an allocation is a bijection $\mu:N\to O$ (or, equivalently, a profile $\mu=(\mu_{i})_{i\in N}$ of distinct objects). Therefore, in the Shapley-Scarf model, all allocations are balanced.
Since each agent receives one object, it is easy to see that the weak endowment lower bound, individual rationality, and the strong endowment lower bound, are all equivalent. Although the definitions of drop strategy-proofness and truncation-proofness are exactly the same as before, it is helpful to reformulate the definitions of truncation strategies and truncation-proofness as follows.
Let $P_{i},P^{\prime}_{i}\in{\mathcal{P}}$. We say that $P^{\prime}_{i}$ is a truncation strategy for $P_{i}$ if
  1. (1)
    $\{o\in O\mid o\mathbin{R^{\prime}_{i}}o_{i}\}\subsetneq\{o\in O\mid o\mathbin{R_{i}}o_{i}\}$, and
  2. (2)
    For each $o,o^{\prime}\in O\setminus o_{i}$, $o\mathbin{P^{\prime}_{i}}o^{\prime}$ if and only if $o\mathbin{P_{i}}o^{\prime}$.
Intuitively, $P^{\prime}_{i}$ is obtained from $P_{i}$ by “pushing up” agent $i$’s own object $o_{i}$, whilst leaving the relative rankings of all other objects unchanged. Let ${\mathcal{T}}\left(P_{i}\right)$ denote the set of all truncation strategies of $P_{i}$. Note that the definition from Section 2 is equivalent to this definition when each agent is endowed with one object.
Theorem 3.2 implies the following corollary for the Shapley-Scarf model. [^20] Altuntaş et al. (2023, Corollary 3) show that for the Shapley-Scarf model, TTC is characterized by Pareto efficiency, individual rationality, and drop strategy-proofness. In view of Proposition 2.11, we generalize their result by showing that drop strategy-proofness can be replaced with truncation-proofness.
Theorem 3.5.
In the Shapley-Scarf model, a rule satisfies
  • Pareto efficiency,
  • individual rationality, and
  • truncation-proofness
if and only if it is TTC.
Because of the interest in characterizations of TTC for the Shapley-Scarf model, we give a proof of Theorem 3.5 below. Our proof of remarkably simple, and it shares the same ideas as the proof of Theorem 3.2; it should therefore serve as intuition for the proof of the general result.
3.1.1. Proof of Theorem 3.5
It is known that TTC satisfies all properties. Therefore, it suffices to prove the uniqueness. Toward contradiction, suppose that $\varphi$ satisfies the properties but $\varphi\neq\varphi^{\text{TTC}}$.
We borrow the clever notions of size and similarity from Sethuraman (2016) and Ekici (2023), respectively.
For each $P\in\mathcal{P}^{N}$, and each $i\in N$, let $s_{i}(P_{i})$ be the number of objects that are weakly better than $o_{i}$ at $P_{i}$, i.e., $s_{i}(P_{i})=|\{o\in O\mid o\mathbin{R_{i}}o_{i}\}|$. Then the size of $P$ is $s(P)=\sum_{i\in N}s_{i}(P_{i})$.
For each $P\in\mathcal{P}^{N}$, and each $t\in\mathbb{N}$, let $\mathcal{C}_{t}(P)$ be the set of cycles that obtain at step $t$ of $\text{TTC}\left(P\right)$. Similarly, let $C_{t}(P)\in\mathcal{C}_{t}(P)$ be the cycle that is executed at step $t$ of $\text{TTC}\left(P\right)$. The similarity between $\varphi$ and $\varphi^{\text{TTC}}$ is a function $\rho:\mathcal{P}^{N}\to\mathbb{N}$ such that, for all $P\in\mathcal{P}$, $\rho(P)$ is defined as follows.
  • Step 1. If $\varphi(P)$ does not execute $C_{1}(P)$, then $\rho(P)=1$. Suppose $\varphi(P)$ executes $C_{1}(P)$. If $\text{TTC}(P)$ terminates at Step 1, then $\rho(P)=|O|$; otherwise, proceed to Step $2$.
  • Step $t$ ($\geq 2$). If $\varphi(P)$ does not execute $C_{t}(P)$, then $\rho(P)=t$. Suppose $\varphi(P)$ executes $C_{t}(P)$. If $\text{TTC}(P)$ terminates at Step $t$, then $\rho(P)=|O|$; otherwise, proceed to Step $t+1$.
Note that, for all $P\in\mathcal{P}^{N}$, (i) $\rho(P)\leq|O|$, and (ii) $\rho(P)=|O|$ if and only if $\varphi(P)=\varphi^{\text{TTC}}(P)$.
Suppose that $\min_{P\in\mathcal{P}^{N}}\rho\left(P\right)=t$. Then $\varphi\neq\varphi^{\text{TTC}}(P)$ implies that $t<|O|$. Among all profiles in $\{P\in\mathcal{P}^{N}\mid\rho(P)=t\}$, let $P$ be one whose size is smallest. Hence, for any other profile $\tilde{P}\neq P$, we have
  • (1)
    $\rho(P)<\rho(\tilde{P})$, or
  • (2)
    $\rho(P)=\rho(\tilde{P})$ and $s(P)\leq s(\tilde{P})$.
Since $\rho(P)=t$, $\varphi(P)$ executes the cycles $C_{1}(P),C_{2}(P),\dots,C_{t-1}(P)$ but not $C_{t}(P)$. Let $\mu\coloneqq\varphi^{\text{TTC}}(P)$, $C\coloneqq C_{t}(P)$, and $O^{t}\coloneqq O\backslash\bigcup_{\tau=1}^{t-1}O\left(C_{\tau}\left(P\right)\right)$. Thus, there is an agent $i\in N(C)$ such that, although agent $i$ points to $\mu_{i}$ on $C$, she does not receive $\mu_{i}$ at $\varphi(P)$, i.e., $\mu_{i}\neq\varphi_{i}(P)$. Note that, by the definition of $\text{TTC}(P)$, $\mu_{i}$ is agent $i$’s most-preferred object in $O^{t}$ at $P_{i}$, i.e., $\mu_{i}=\top_{P_{i}}\left(O^{t}\right)$. Thus,
(3) $\mu_{i}\mathbin{P_{i}}\varphi_{i}(P).$
If $|N(C)|=1$, then $C=(i,o_{i},i)$ and $\mu_{i}=o_{i}\mathbin{P_{i}}\varphi_{i}(P)$, which violates individual rationality. Thus, $|N(C)|\geq 2$.
Next, we show that $\varphi_{i}(P)=o_{i}$. Toward contradiction, suppose that $\varphi_{i}(P)\neq o_{i}$. By (3) and individual rationality, $\mu_{i}\mathbin{P_{i}}\varphi_{i}(P)\mathbin{P_{i}}o_{i}$.
Let $P^{\prime}_{i}$ be the truncation of $P_{i}$ at $\mu_{i}$, i.e., $P^{\prime}_{i}:\dots,\mu_{i},o_{i},\dots$. [^21] A similar argument using drop strategy-proofness applies if $P^{\prime}_{i}$ is obtained from $P_{i}$ by dropping object $\varphi_{i}(P)$. Let $P^{\prime}\coloneqq(P^{\prime}_{i},P_{-i})$. Then $s_{i}(P^{\prime}_{i})<s_{i}(P_{i})$ and $s(P^{\prime})<s(P)$.
By the definition of TTC, for each step $\tau\in\{1,\dots,t\}$, the procedures $\text{TTC}(P)$ and $\text{TTC}(P^{\prime})$ generate and execute the same cycles, i.e., for each $\tau$, $C_{\tau}(P)=C_{\tau}(P^{\prime})$. Moreover, the choice of $P$ implies that $\rho(P^{\prime})\geq\rho(P)=t$. Thus, $\varphi(P^{\prime})$ executes all cycles in $\{C_{1}(P),\dots,C_{t-1}(P)\}$. Thus, $\varphi_{i}(P^{\prime})\in O^{t}$. Moreover, since $\varphi$ is truncation-proof, we know that $i$ cannot be better of by misreporting $P^{\prime}_{i}$ at $P$. Together with (3), we conclude that $\mu_{i}\neq\varphi_{i}(P^{\prime})$. Consequently, $\varphi(P^{\prime})$ does not execute $C_{t}(P^{\prime})$, which means that $\rho(P^{\prime})=t$. But then we have $\rho(P^{\prime})=\rho(P)$ and $s(P^{\prime})<s(P)$, which contradicts the choice of $P$! Thus, we conclude that $\varphi_{i}(P)=o_{i}$.
Let $j$ be the agent who points to $o_{i}$ on $C$, i.e., $j\in N(C)$ is such that $\mu_{j}=o_{i}$. Then $\varphi_{i}(P)=o_{i}=\mu_{j}$ implies that $\varphi_{j}(P)\neq\mu_{j}$. Thus, $\mu_{j}\mathbin{P_{j}}\varphi_{j}(P).$ A similar argument shows that $\varphi_{j}(P)=o_{j}$.
By repeating the same argument for each agent in $N(C)$, we can show that each agent in $N(C)$ is assigned her endowment at $\varphi(P)$, i.e., for each $i^{\prime}\in N(C)$, $\varphi_{i^{\prime}}(P)=o_{i^{\prime}}$. But then $\varphi$ is not Pareto efficient, because the agents in $N(C)$ can benefit by executing $C$. This contradiction completes the proof of Theorem 3.5.
4. Other preference domains
In this section, we consider the extension of TTC and our properties to more general preference domains, namely, the sets $\mathcal{R}$ of responsive preferences, and $\mathcal{CL}$ of conditionally lexicographic preferences. Note that all domains are sub-domains of the strict preference domain.
4.1. Responsive preferences
A preference relation $P_{i}$ on $2^{O}$ is responsive if for any bundle $X\in 2^{O}$ and any two distinct objects $y,z\in O\setminus X$, $y\mathbin{P_{i}}z$ if and only if $(X\cup y)\mathbin{P_{i}}(X\cup z)$. Let $\mathcal{R}$ denote the set of responsive preference relations on $2^{O}$. Note that $\mathcal{L}\subseteq\mathcal{R}$, with strict inclusion whenever $|O|\geq 3$.
Given a responsive preference relation $P_{i}\in\mathcal{R}$, let $\succ^{P_{i}}$ denote the induced (strict) linear order on $O$ such that, for all $x,y\in O$, $x\mathbin{P_{i}}y$ if and only if $x\succ^{P_{i}}y$. We denote the reflexive closure of $\succ^{P_{i}}$ by $\succeq^{P_{i}}$, i.e., for all $x,y\in O$, $x\mathbin{R_{i}}y$ if and only if $x\succeq^{P_{i}}y$. Given a profile $P=(P_{i})_{i\in N}\in\mathcal{R}^{N}$, we denote $\succ^{P}=(\succ^{P_{i}})_{i\in N}$. We represent a strict linear order over $O$ as an ordered list of objects, e.g., $\succ^{P_{i}}:x_{1},x_{2},\dots,x_{m}$ means that $x_{1}\succ^{P_{i}}x_{2}\succ^{P_{i}}\cdots\succ^{P_{i}}x_{m}$.
4.1.1. Rules and their properties
On the responsive domain, we restrict attention to the class of rules that are individual-good-based in the sense that they depend only on the profile of induced linear orders associated with a given preference profile (see, e.g., Biró et al. 2022). Formally, for each rule $\varphi$ on $\mathcal{R}$, we assume that [^22] This restriction is automatically satisfied for rules on $\mathcal{L}$, since each lexicographic $P_{i}$ is completely determined by its induced linear order $\succ^{P_{i}}$. For rules $\varphi$ on $\mathcal{R}$, this restriction is substantive, but it is often justified on the basis of “simplicity.”
(4) $\text{for all }P,P^{\prime}\in\mathcal{R}^{N},\succ^{P}=\succ^{P^{\prime}}\text{ implies }\varphi(P)=\varphi(P^{\prime}).$
The generalized Top Trading Cycles rule of Section 2.4, which extends to all $P\in\mathcal{R}^{N}$ via the procedure $\text{TTC}(P)$, is one such individual-good-based rule.
One could equivalently assume that agents report only their induced linear orders $\succ^{P}$, which is common in practice. [^23] The restriction to individual-good-based rules is relaxed in Section 4.2. There, we introduce a more general domain exhibiting complementarity, the conditionally lexicographic preferences, and we consider a more general class of rules which are not based solely on the agents’ preferences over individual objects. Note, however, that in our properties below all allocations are evaluated based on the agents’ actual preferences $P$.
Letting $\mathcal{P}\coloneqq\mathcal{R}$, the definitions of Pareto efficiency, balancedness, individual rationality, the strong endowment lower bound, the weak endowment lower bound, and strategy-proofness from Section 2.3 carry over exactly to this setting.
At a technical level, drop strategies, subset drop strategies, and truncation strategies are defined exactly as in Section 2.3. However, their interpretation is slightly different in this more general setting. First note that if $P_{i}\in\mathcal{L}$ is lexicographic, then there is a unique $P^{\prime}_{i}\in\mathcal{L}$ obtained from $P_{i}$ by dropping an object $x$ (or by dropping a subset $X$). Similarly, if $P_{i}\in\mathcal{R}\setminus\mathcal{L}$, then there is a unique linear order over individual objects $\succ^{\prime}_{i}$ obtained from $\succ^{P_{i}}$ by dropping $x$ (or $X$). However, there are many $P^{\prime}_{i}\in\mathcal{R}$ such that $\succ^{P^{\prime}_{i}}=\succ^{\prime}_{i}$, and any such $P^{\prime}_{i}$ is a (subset) drop strategy for $P_{i}$. Thus, drop strategies, subset drop strategies, and truncation strategies are all defined with respect to the agents’ induced linear orders.
Example 4.1 (Drop strategies and truncation strategies).
As in Example 2.10, suppose $\omega_{i}=\{x,y\}$ and that $P_{i}\in\mathcal{R}$ is such that $\succ^{P_{i}}:a,b,\underline{x},c,d,\underline{y},e$ (endowments are underlined). Then:
  • any $P^{1}_{i}\in\mathcal{R}$ such that $\succ^{P^{1}_{i}}:b,\underline{x},c,d,\underline{y},e,a$ is a drop strategy for $P_{i}$; it is obtained by dropping object $a$. Note that $P^{1}_{i}$ is not a truncation strategy for $P_{i}$.
  • any $P^{2}_{i}\in\mathcal{R}$ such that $\succ^{P^{2}_{i}}:a,b,\underline{x},c,\underline{y},d,e$ is a truncation strategy for $P_{i}$; it is obtained from $P_{i}$ by dropping the subset $\{d,e\}$ (or subsequently dropping $d$ then $e$).
  • any $P^{3}_{i}\in\mathcal{R}$ such that $\succ^{P^{3}_{i}}:a,\underline{x},\underline{y},b,c,d,e,$ is a truncation strategy for $P_{i}$; it is obtained from $P_{i}$ by dropping the subset $\{b,c,d,e\}$ (or subsequently dropping $b,c,d$, then $e$). $\diamond$
With the slightly modified interpretation in mind, the definitions of drop strategy-proofness and truncation-proofness are identical to those stated in Definitions 2.8 and 2.9.
4.1.2. Characterizations of TTC
It is easy to see that TTC satisfies the strong endowment lower bound. Therefore, TTC satisfies balancedness and the weak endowment lower bound. Though TTC is not Pareto efficient on $\mathcal{R}$, it satisfies the weaker ig-efficiency.
The following example illustrates that if even one agent has responsive preferences which are not lexicographic, then that agent may have an incentive to manipulate via a drop strategy. Thus, TTC is not drop strategy-proof on the responsive domain $\mathcal{R}$ [^24] If there are only two agents, then TTC is drop strategy-proof as each agent can only drop the opponent’s endowed objects.
Example 4.2.
Let $N=\{1,2,3\}$ and $\omega=\left(\left\{a,b\right\},\left\{c,d\right\},\left\{e,f\right\}\right)$. Let $P_{2},P_{3}\in\mathcal{L}$ be two lexicographic preference relations such that
$P_{2}:e,f,b,\dots;\quad P_{3}:a,c,d,\dots.$
Let $P_{1}$ be an additive preference relation such that $\succ^{P_{1}}:c,f,e,d,a,b$ and $\{d,f\}\mathbin{P_{1}}\{c,b\}$. [^25] A strict preference relation $P_{i}$ is additive if it is represented by an additive utility function, i.e., a function $u_{i}:2^{O}\to\mathbb{R}_{+}$ such that for all $X\in 2^{O}$, $u_{i}(X)=\sum_{x\in X}u_{i}(x)$. Let $\mathcal{A}$ denote the domain of additive preferences. Note that $\mathcal{A}\subseteq\mathcal{R}$, with strict inclusion whenever $|O|\geq 4$. Then $\varphi^{\text{TTC}}\left(P\right)=\left(\left\{c,b\right\},\left\{e,f\right\},\left\{a,d\right\}\right)$.
Now assume that agent $1$ drops object $c$ from her induced linear order, i.e., she reports $P^{\prime}_{1}$ such that $\succ^{P^{\prime}_{1}}:f,e,d,a,b,c.$ Then $\varphi^{\text{TTC}}\left(P^{\prime}_{1},P_{-1}\right)=\left(\left\{d,f\right\},\left\{e,b\right\},\left\{a,c\right\}\right)$. Because $\varphi_{1}^{\text{TTC}}\left(P^{\prime}_{1},P_{-1}\right)=\{d,f\}\mathbin{P_{1}}\{c,b\}=\varphi_{1}^{\text{TTC}}\left(P\right)$, we see that TTC is not drop strategy-proof on $\mathcal{R}$ [^26] In fact, this example demonstrates that TTC is not drop strategy-proof on $\mathcal{A}$ either. $\diamond$
However, we show that TTC is truncation-proof on the responsive domain $\mathcal{R}$. This highlights another important distinction between drop strategies and truncation strategies. The positive results are summarized in the following proposition.
Proposition 4.3.
On the responsive domain, TTC satisfies ig-efficiency, the strong endowment lower bound (hence, balancedness and the weak endowment lower bound), and truncation-proofness.
Let $\varphi$ be any individual-good-based rule on $\mathcal{R}$ that satisfies ig-efficiency, balancedness, the weak endowment lower bound, and truncation-proofness. An argument similar to the proof of Theorem 3.2 demonstrates that $\varphi$ must coincide with TTC on $\mathcal{L}^{N}$. That is, $\varphi(P)=\varphi^{\text{TTC}}(P)$ for all $P\in\mathcal{L}^{N}$. Now consider any $P^{\prime}\in\mathcal{R}^{N}$. Let $P\in\mathcal{L}^{N}$ be the uniquely corresponding lexicographic preference profile such that $\succ^{P}=\succ^{P^{\prime}}$. Because $\varphi$ and TTC are “individual-good-based,” it holds that
$\varphi(P^{\prime})=\varphi(P)=\varphi^{\text{TTC}}(P)=\varphi^{\text{TTC}}(P^{\prime}).$
Thus, we have shown that our characterization extends to all of $\mathcal{R}^{N}$. [^27] We owe a large debt to Biró et al. (2022), who used a similar approach to extend their characterization of cTTC for the multi-unit Shapley-Scarf model from the lexicographic domain to the responsive domain.
Theorem 4.4.
On the responsive domain, an individual-good-based rule satisfies
  • ig-efficiency,
  • balancedness,
  • the weak endowment lower bound, and
  • truncation-proofness
if and only if it is TTC.
It turns out that the weak endowment lower bound can be replaced with individual rationality in the statement of Theorem 4.4. To see how, we first state a useful lemma.
Lemma 4.5.
On the responsive domain, if an individual-good-based rule is balanced and individually rational, then it satisfies the weak endowment lower bound.
The following corollary is an immediate consequence of Theorem 4.4, Lemma 4.5, and the fact that TTC is individually rational.
Corollary 4.6.
On the responsive domain, an individual-good-based rule satisfies
  • ig-efficiency,
  • balancedness,
  • individual rationality, and
  • truncation-proofness
if and only if it is TTC.
4.2. Conditionally lexicographic preferences
We now introduce a generalization of (purely) lexicographic preferences, called conditionally lexicographic preferences. Loosely speaking, a preference relation $P_{i}$ on $2^{O}$ is conditionally lexicographic if, given any bundle $Y\subsetneq O$, there is an object $o\in O\setminus Y$ such that agent $i$ lexicographically prefers adding $o$ to $Y$ rather than adding any other bundle $X\subseteq O\setminus(Y\cup o)$. Such preferences are widely studied in Artificial Intelligence. [^28] See Domshlak et al. (2011) and Pigozzi et al. (2016) for surveys.
Formally, conditionally lexicographic preferences are modeled by lexicographic preference trees (LP trees) over the set $O$ of objects.
Definition 4.7 (Lexicographic preference tree (LP tree)).
An LP tree over $O$ is an arborescence $\tau_{i}$ such that
  • every vertex $v$ is labeled with an object $o\left(v\right)\in O$.
  • every object appears exactly once on any path from the root of $\tau_{i}$ to a leaf of $\tau_{i}$.
  • every internal (non-leaf) vertex has two outgoing edges, labeled by $o\left(v\right)$ and $\neg o\left(v\right)$ respectively.
We shall refer to edges labeled with $o\left(v\right)$ and $\neg o\left(v\right)$ as “in edges” and “not-in edges”, respectively.
Figures 2 and 2 provide a graphical representation of two LP trees, $\tau_{i}$ and $\tau^{*}_{i}$, over the set $O=\{a,b,c,d\}$ of objects.
Given $X\subseteq O$, let $\tau_{i}\left(X\right)$ be a directed path from the root of $\tau_{i}$ to a leaf of $\tau_{i}$ containing only edges consistent with $X$, i.e., edges $\left(v,v^{\prime}\right)$ labeled with $o\left(v\right)$ if $o\left(v\right)\in X$ and $\neg o\left(v\right)$ if $o\left(v\right)\notin X$. Note that such a path $\tau_{i}\left(X\right)$ is unique. Given bundles $A$ and $B$ with $A\neq B$, let $\tau_{i}\left(A,B\right)$ denote the first vertex $v$ visited by both $\tau_{i}\left(A\right)$ and $\tau_{i}\left(B\right)$ and such that $o\left(v\right)\in A\triangle B$, i.e., $o\left(v\right)$ belongs to exactly one of $A$ and $B$. Equivalently, $\tau_{i}\left(A,B\right)$ is the last vertex which is common to both $\tau_{i}\left(A\right)$ and $\tau_{i}\left(B\right)$.
Definition 4.8 (Conditionally lexicographic preferences).
The preference relation $P_{\tau_{i}}$ associated with an LP tree $\tau_{i}$ is defined as follows:
  1. (1)
    for all $A,B\subseteq O$ with $A\neq B$, $\left[AP_{\tau_{i}}B\iff o\left(\tau_{i}\left(A,B\right)\right)\in A\backslash B\right]$.
  2. (2)
    for all $A,B\subseteq O$, $\left[AR_{\tau_{i}}B\iff\left(A=B\text{ or }AP_{\tau_{i}}B\right)\right]$.
Figure 1. An LP tree $\tau_{i}$
Figure 2. An LP tree $\tau^{*}_{i}$
Notes: The paths $\tau_{i}(\{a,b\})$ and $\tau^{*}_{i}(\{a,b\})$ are highlighted in red, while the paths $\tau_{i}(\{a,c\})$ and $\tau^{*}_{i}(\{a,c\})$ are highlighted in blue. The last vertex in $\tau_{i}$ common to both $\tau_{i}(\{a,b\})$ and $\tau_{i}(\{a,c\})$, namely $\tau_{i}(\{a,b\},\{a,c\})$, is highlighted in blue. The last vertex in $\tau^{*}_{i}$ common to both $\tau^{*}_{i}(\{a,b\})$ and $\tau^{*}_{i}(\{a,c\})$, namely $\tau^{*}_{i}(\{a,b\},\{a,c\})$, is highlighted in red. Note that $o(\tau_{i}(\{a,b\},\{a,c\})\in\{a,c\}\setminus\{a,b\}$ implies that $\{a,c\}\mathbin{P_{\tau_{i}}}\{a,b\}$, while $o(\tau^{*}_{i}(\{a,b\},\{a,c\})\in\{a,b\}\setminus\{a,c\}$ implies that $\{a,b\}\mathbin{P_{\tau^{*}_{i}}}\{a,c\}$. Because any two directed paths $(v_{1},v_{2},v_{3},v_{4})$ and $(v^{\prime}_{1},v^{\prime}_{2},v^{\prime}_{3},v^{\prime}_{4})$ in $\tau^{*}_{i}$ from the root to a leaf are labeled with the same objects, the relation $P_{\tau^{*}_{i}}$ is (purely) lexicographic. $\diamond$
A preference relation $P_{i}$ on $2^{O}$ is called conditionally lexicographic if there exists an LP tree $\tau_{i}$ such that $P_{i}=P_{\tau_{i}}$. Let $\mathcal{CL}$ denote the set of all conditionally lexicographic preferences, and $\mathcal{CL}^{N}$ the set of all conditionally lexicographic preference profiles. There is a one-to-one correspondence between LP trees $\tau_{i}$ and conditionally lexicographic preference relations $P_{\tau_{i}}$. We therefore identify the set $\mathcal{CL}$ with the set of LP trees over $O$. In particular, we may refer to LP trees $\tau_{i}$ as preference relations, and write statements such as $\tau_{i}\in\mathcal{CL}$ as well as $P_{i}\in\mathcal{CL}$. Given $P_{i}\in\mathcal{CL}$, let $\tau_{P_{i}}$ denote the LP tree associated with $P_{i}$.
It is easy to see that $\mathcal{L}\subseteq\mathcal{CL}$, with strict inclusion whenever $|O|\geq 3$. Indeed, every $P_{i}\in\mathcal{L}$ is associated with an LP tree $\tau_{P_{i}}$ such that all vertices with the same depth are labelled with the same object, i.e., for any two directed paths $(v_{1},v_{2},\dots,v_{k})$ and $(v^{\prime}_{1},v^{\prime}_{2},\dots,v^{\prime}_{k})$ starting from the root of $\tau_{P_{i}}$, it holds that $o(v_{\ell})=o(v^{\prime}_{\ell})$ for all $\ell\in\{1,\dots,k\}$. Moreover, $\mathcal{CL}\cap\mathcal{R}=\mathcal{L}$.
The following characterization of conditionally lexicographic preferences is useful. It says that a preference relation is conditionally lexicographic if and only if, for any two disjoint bundles $X,Y\in 2^{O}$ with $X\neq\emptyset$, there is an object in $X$, denoted $\top_{P_{i}}(X\mid Y)$, which is “lexicographically best” among $X$ conditional on receiving $Y$. Its proof is straightforward and it is omitted.
Proposition 4.9.
A preference relation $P_{i}$ on $2^{O}$ is conditionally lexicographic if and only if the following property holds: for all disjoint subsets $X,Y\in 2^{O}$ such that $X\neq\emptyset$, there is a unique object in $X$, denoted $x^{*}\coloneqq\top_{P_{i}}(X\mid Y)$, such that
(5) for all $\text{for all }Z\subseteq X\setminus x^{*},\quad(Y\cup x^{*})\mathbin{P_{i}}(Y\cup Z).$
That is, $\top_{P_{i}}(X\mid Y)$ is agent $i$’s (lexicographically) best object in $X$ conditional on receiving $Y$ [^29] We note that $\top_{P_{i}}(X\mid Y)=o(\tau_{P_{i}}(Y,Y\cup X))$.
4.2.1. Rules and their properties
Let $\mathcal{P}\coloneqq\mathcal{CL}$ in what follows. In this subsection, a rule is a function $\varphi:\mathcal{P}^{N}\to\mathcal{M}$ which is not necessarily “individual-good-based,” i.e., (4) need not hold.
Most of the properties, namely Pareto efficiency, individual rationality, and balancedness, are defined exactly as in Section 2.3. However, the definitions of the weak endowment lower bound and drop strategy-proofness must be adapted to the present setting.
To define the weak endowment lower bound, we first need some notation. Given an LP tree $\tau_{i}\in\mathcal{CL}$ and a vertex $v$ of $\tau_{i}$, let $a\left(v\right)$ denote the set of objects labeling the ancestors of $v$, i.e., if $\left(v_{1},v_{2},\dots,v_{k}=v\right)$ is the unique path from the root of $\tau_{i}$ to $v$, then $a\left(v\right)=\left\{o\left(v_{1}\right),o\left(v_{2}\right),\dots,o\left(v_{k}\right)\right\}$. Given a preference relation $P_{i}\in\mathcal{CL}$ and a bundle $X\in 2^{O}$, let $w_{P_{i}}\left(\omega_{i}\mid X\right)$ denote the vertex corresponding to the worst object in $\omega_{i}$ conditional on receiving $X$. That is, $w_{P_{i}}\left(\omega_{i}\mid X\right)$ is the unique vertex on $\tau_{P_{i}}\left(X\right)$ such that $o\left(w_{P_{i}}\left(\omega_{i}\mid X\right)\right)\in\omega_{i}$ and $\omega_{i}\subseteq a\left(w_{P_{i}}\left(\omega_{i}\mid X\right)\right)$. This construction is illustrated graphically in figures 4 and 4.
An allocation $\mu$ satisfies the weak endowment lower bound at $P\in\mathcal{CL}^{N}$ if, for all $i\in N$, $\mu_{i}\subseteq a\left(w_{P_{i}}\left(\omega_{i}\mid\mu_{i}\right)\right)$. That is, every object in $\mu_{i}$ labels a vertex $v$ on $\tau_{P_{i}}\left(\mu_{i}\right)$ that weakly precedes vertex $w_{P_{i}}\left(\omega_{i}\mid\mu_{i}\right)$. Intuitively, no agent $i$ is assigned an object which is “conditionally worse” than the “conditionally worst” endowed object in $\omega_{i}$ (conditional on receiving $\mu_{i}$). Note that this definition coincides with the original definition when agents have (purely) lexicographic preferences.
Definition 4.10 (Weak endowment lower bound).
A rule $\varphi$ satisfies the weak endowment lower bound if, for each profile $P\in\mathcal{CL}^{N}$, the allocation $\varphi\left(P\right)$ satisfies the weak endowment lower bound at $P$.
Figure 3. The LP tree $\tau_{i}$
Figure 4. The LP tree $\tau^{*}_{i}$
Notes: Recall the LP trees $\tau_{i}$ and $\tau^{*}_{i}$ from Figures 2 and 2. Suppose agent $i$’s endowment is $\omega_{i}=\{b,c\}$. Then the vertices in $\{w_{P_{\tau_{i}}}(\omega_{i}\mid X)\mid X\in 2^{O}\}$ and $\{w_{P_{\tau^{*}_{i}}}(\omega_{i}\mid X)\mid X\in 2^{O}\}$ are highlighted in blue. Note that there is exactly one such vertex on any path from the root to a leaf. Under $\tau_{i}$, the weak endowment lower bound requires that agent $i$ is not assigned any bundle in $\{\{a,d\},\{a,b,d\},\{c,d\},\{d\}\}$; under $\tau^{*}_{i}$, it requires that agent $i$ does not receive any bundle containing object $d$. $\diamond$
The definition of drop strategies extends naturally to this setting. Let $P_{i}\in\mathcal{CL}$. We say that $P^{\prime}_{i}\in\mathcal{CL}$ is a drop strategy for $P_{i}$ if there exists $x\in O\setminus\omega_{i}$ such that:
  1. (1)
    for all nonempty bundles $Y\subseteq O\setminus\{x\}$, $Y\mathbin{P^{\prime}_{i}}x$; and
  2. (2)
    for any $Y,Z\subseteq O\setminus\{x\}$, $Y\mathbin{P^{\prime}_{i}}Z$ if and only if $Y\mathbin{P_{i}}Z$.
In this case, we say that $P^{\prime}_{i}$ is obtained from $P_{i}$ by dropping object $x$ [^30] As is the case for purely lexicographic preferences (but not responsive preferences), there is exactly one $P^{\prime}_{i}$ obtained from $P_{i}$ by dropping object $x$. Note that this definition coincides with the original definition from Section 2.3 whenever the agent’s true preference relation $P_{i}$ is purely lexicographic. Let $\mathcal{D}(P_{i})$ denote the set of all drop strategies for $P_{i}$.
The drop strategy $P^{\prime}_{i}$ obtained from $P_{i}$ by dropping object $x$ can be equivalently defined in terms of its LP tree. Let $\tau_{P_{i}}^{-x}$ be the LP tree over $O\setminus\{x\}$ that represents the restriction of $P_{i}$ to subsets of $O\setminus\{x\}$. [^31] We can construct $\tau_{P_{i}}^{-x}$ as follows. For each vertex $v$ of $\tau_{P_{i}}$, let $T_{v}$ denote the maximal subtree of $\tau_{P_{i}}$ consisting of a vertex $v$ of $\tau_{P_{i}}$ together with all of its successors, and let $v^{\prime}$ be the child of $v$ whose incoming edge $(v,v^{\prime})$ is labeled with $\neg o(v)$. For each vertex $v$ of $\tau_{P_{i}}$ such that $o(v)=x$, simply replace $T_{v}$ with the subtree $T_{v^{\prime}}$ (or the empty tree if $v$ is a leaf of $\tau_{P_{i}}$). Then $\tau_{P^{\prime}_{i}}$ is the LP tree over $O$ obtained by appending each leaf of $\tau_{P_{i}}^{-x}$ with two child nodes, each labeled with object $x$, and with appropriately labeled incoming edges (i.e., an “in edge” and “not-in edge”). This construction is illustrated in Figures 6 and 6.
Figure 5. The LP tree $\tau_{i}$
Figure 6. The LP tree $\tau^{\prime}_{i}$ obtained by dropping $b$
Notes: Recall the LP tree $\tau_{i}$ from Figures 2 and 4. Suppose agent $i$ drops object $b$ to the bottom of $\tau_{i}$. The LP tree $\tau^{\prime}_{i}$ is obtained from $\tau_{i}$ by (i) replacing each maximal subtree whose root is an internal vertex labeled with $b$ with the blue subtree that follows the corresponding $\neg b$ edge, then (ii) appending each blue leaf of the resulting tree with two child nodes, each labeled with object $b$, and with appropriately labeled incoming edges. Note that $\tau^{\prime}_{i}$ coincides with the drop strategy obtained by dropping $x$ from the LP tree $\tau^{*}_{i}$ from Figures 2 and 4. $\diamond$
Definition 4.11 (Drop strategy-proofness).
A rule $\varphi$ is drop strategy-proof if, for each profile $P\in\mathcal{CL}^{N}$, each $i\in N$, and each $P^{\prime}_{i}\in\mathcal{D}(P_{i})$, $\varphi_{i}(P)\mathbin{R_{i}}\varphi_{i}(P^{\prime}_{i},P_{-i})$.
4.2.2. Augmented Top Trading Cycles
We now introduce the Augmented Top Trading Cycles (ATTC) rule, which was first proposed by Fujita et al. (2018). This rule coincides with TTC when restricted to the lexicographic domain.
At each step of the ATTC procedure, every agent points to her favorite object remaining, conditional on the objects already assigned to her, and every remaining object points to its owner. At least one cycle obtains. As in the standard case, our proofs rely on the variant of ATTC that executes only the smallest cycle at each step.
Formally, for each profile $P\in{\mathcal{P}}^{N}$, ATTC selects the allocation $\varphi^{\text{ATTC}}\left(P\right)$ determined by the following procedure, which we call $\text{ATTC}\left(P\right)$.
Algorithm: $\text{ATTC}\left(P\right)$
Input:: A preference profile $P\in\mathcal{CL}^{N}$. Output:: An allocation $\varphi^{\text{TTC}}\left(P\right)$. Step 0.: Define $\mu^{0}\coloneqq\left(\emptyset\right)_{i\in N}$ and $O^{1}\coloneqq O$. Step $t\geq 1$.: Construct a bipartite directed graph with independent vertex sets $N$ and $O^{t}$. For each agent $i\in N$, there is a directed edge from $i$ to agent $i$’s most-preferred object in $O^{t}$ conditional on receiving $\mu^{t-1}_{i}$, namely $\top_{P_{i}}(O^{t}\mid\mu^{t-1}_{i})$. For each object $o\in O^{t}$, there is a directed edge from $o$ to its owner, $\omega^{-1}\left(o\right)$. At least one cycle exists. Let $\mathcal{C}_{t}(P)$ denote the set of cycles that obtain at step $t$, and let $C_{t}(P)$ be the smallest cycle in $\mathcal{C}_{t}(P)$, i.e., $\min N(C_{t}(P))\leq\min N(C)$ for all $C\in\mathcal{C}_{t}(P)$. Assign each agent $i\in N(C_{t}(P))$ the object to which she points on $C_{t}(P)$; that is, let $\mu^{t}=\left(\mu_{i}^{t}\right)_{i\in N}$ be the allocation such that (i) for all $i\in N(C_{t}(P))$, $\mu_{i}^{t}=\mu_{i}^{t-1}\cup\left\{\top_{P_{i}}\left(O^{t}\mid\mu^{t-1}_{i}\right)\right\}$, and (ii) for all $i\in N\backslash N(C_{t}(P))$, $\mu_{i}^{t}=\mu_{i}^{t-1}$. Let $O^{t+1}\coloneqq O^{t}\backslash O(C_{t}(P))$ denote the set of objects remaining at Step $t+1$. If $O^{t+1}\neq\emptyset$, then proceed to Step $t+1$; otherwise, proceed to Termination. Termination.: Because $O$ is finite, and $\left|O^{1}\right|>\left|O^{2}\right|>\cdots>\left|O^{t}\right|$, the procedure terminates at some step $T$. Return the allocation $\varphi^{\text{TTC}}\left(P\right)\coloneqq\mu^{T}$.
4.2.3. A characterization of ATTC
We first show that ATTC is drop strategy-proof on the conditionally lexicographic domain. This extends the result of Altuntaş et al. (2023) for the lexicographic domain.
Proposition 4.12.
On the conditionally lexicographic domain, ATTC is drop strategy-proof. [^32] Recall that we define a weak notion of drop strategy-proofness that defends only against manipulations in which agents drop objects they do not own. One can easily show that ATTC satisfies a stronger notion that also defends against manipulations in which agents drop objects they do own. Our weak version suffices for our characterization result.
Our characterization of TTC on the lexicographic domain (with drop strategy-proofness) extends naturally to the conditionally lexicographic domain. This settles an open question posed by Fujita et al. (2018, p. 531).
Theorem 4.13.
On the conditionally lexicographic domain, a rule satisfies
  • Pareto efficiency,
  • balancedness,
  • the weak endowment lower bound, and
  • drop strategy-proofness
if and only if it is ATTC.
5. Maximal domain results
As mentioned in Section 2.3, our two notions of efficiency, namely ig-efficiency and Pareto efficiency, are equivalent on the lexicographic domain. Moreover, within the class of responsive preferences, the lexicographic domain is a maximal domain on which the two properties coincide. The following is an immediate corollary of Proposition 3 of Altuntaş et al. (2023) and the fact that TTC is ig-efficient on $\mathcal{R}$.
Proposition 5.1.
Within the domain of responsive preferences, the lexicographic domain is a maximal domain on which ig-efficiency and Pareto efficiency are equivalent. [^33] That is, if $\mathcal{L}\subsetneq\mathcal{P}\subseteq\mathcal{R}$, then there is a set $N$ of agents, a profile $P\in\mathcal{P}^{N}$, and an allocation $\mu\in\mathcal{M}$ such that $\mu$ is ig-efficient at $P$ but not Pareto efficient at $P$.
We now show that the equivalence between our two efficiency notions extends beyond the class of responsive preferences to the conditionally lexicographic domain.
Proposition 5.2.
On the conditionally lexicographic domain, an allocation is ig-efficient if and only if it is Pareto efficient.
One implication of Proposition 5.2 is that, on the conditionally lexicographic domain, any iterative procedure that identifies and executes Pareto-improving single-object exchanges in a greedy fashion until no such exchange remains will terminate with a Pareto efficient allocation. ATTC is one of many such iterative procedures.
Much like the purely lexicographic domain, the conditionally lexicographic domain is special. Within the class of monotone preferences, the conditionally lexicographic domain is a maximal domain on which the two efficiency properties coincide. A (strict) preference relation $P_{i}$ on $2^{O}$ is monotone if, for any bundles $X,Y\in 2^{O}$, $Y\subsetneq X$ implies $X\mathbin{P_{i}}Y$. Let $\mathcal{\overline{P}}$ denote the set of all monotone preference relations.
Proposition 5.3.
Within the domain of monotone preferences, the conditionally lexicographic domain is a maximal domain on which ig-efficiency and Pareto efficiency are equivalent. [^34] That is, if $\mathcal{CL}\subsetneq\mathcal{P}\subseteq\mathcal{\overline{P}}$, then there is a set $N$ of agents, a profile $P\in\mathcal{P}^{N}$, and an allocation $\mu\in\mathcal{M}$ such that $\mu$ is ig-efficient at $P$ but not Pareto efficient at $P$, [^35] It is straightforward to strengthen Proposition 5.1 if we broaden our definition of allocations to allow agents to be assigned empty bundles. In this case, the conditionally lexicographic domain is a maximal domain among all strict preferences on which the two notions coincide.
6. Final remarks
We conclude the paper with two final remarks.
6.1. Efficiency
First, one may wonder if we can weaken Pareto efficiency to other efficiency notions, such as pairwise efficiency (Ekici, 2023). [^36] A rule satisfies pairwise efficiency if for each profile $P$, there is no pair of agents $i,j$ such that $\varphi_{i}(P)\mathbin{P_{j}}\varphi_{j}(P)$ and $\varphi_{j}(P)\mathbin{P_{i}}\varphi_{i}(P)$. Ekici (2023) originally refers to it as pair efficiency. The following example shows that, even for the Shapley–Scarf model, the answer is negative.
Example 6.1 (Pareto efficiency cannot be weakened to pairwise efficiency).
For simplicity, consider the Shapley–Scarf model with three agents, $N=\left\{1,2,3\right\}$, and suppose $\omega=\left(o_{1},o_{2},o_{3}\right)$. Let $P^{*}\in{\mathcal{P}}^{N}$ be such that (endowment is underlined)
$P_{1}^{*}:o_{2},\underline{o_{1}},o_{3},\quad P_{2}^{*}:o_{3},\underline{o_{2}},o_{1},\quad P_{3}^{*}:o_{1},\underline{o_{3}},o_{2}.$
Then the initial allocation $\omega$ is pairwise efficient at $P^{*}$; but it is not Pareto efficient at $P^{*}$ because it is Pareto-dominated by $\varphi^{\text{TTC}}\left(P^{*}\right)=\left(o_{2},o_{3},o_{1}\right)$.
Now, consider the rule $\varphi$ defined as follows.
$\varphi\left(P\right)=\begin{cases}\omega,&\text{if }P=P^{*}\\ \varphi^{\text{TTC}}\left(P\right),&\text{otherwise}.\end{cases}$
Clearly, $\varphi$ is pairwise efficient and individually rational. It is straightforward to show that $\varphi$ is drop strategy-proof and truncation-proof. $\diamond$
6.2. Beyond individual-good-based rules
Second, one may wonder whether our characterization of TTC on the responsive domain (Corollary 4.4) extends beyond the class of individual-good-based rules. The following example shows that the answer is no.
Example 6.2 (Another rule).
Let $N=\{1,2\}$ and $\omega=(\{a,b\},\{c,d\})$. Let $\mu\coloneqq(\omega_{2},\omega_{1})$. Let $\varphi:\mathcal{A}^{N}\to\mathcal{M}$ be defined as follows.
$\varphi\left(P\right)=\begin{cases}\mu,&\text{if }\varphi^{\text{TTC}}(P)=\omega\text{, and }\mu\text{ Pareto dominates }\omega\text{ at }P;\\ \varphi^{\text{TTC}}(P),&\text{otherwise.}\end{cases}$
It is easy to see that $\varphi$ satisfies ig-efficiency, balancedness, and the weak endowment lower bound. One can also show that $\varphi$ is truncation-proof. Thus, we see that TTC is not the unique rule satisfying our properties on $\mathcal{A}^{N}$.
Also, note that such a rule $\varphi$ coincides with $\varphi^{\text{TTC}}$ on $\mathcal{L}^{N}$: when agents’ preferences are lexicographic, $\varphi^{\text{TTC}}(P)=\omega$ is Pareto efficient, and hence “$\mu$ Pareto dominates $\omega$” is always false. $\diamond$
Appendix A Omitted proofs
Proof of Proposition 2.11
Suppose that $\varphi$ satisfies drop strategy-proofness and the weak endowment lower bound. Let $P^{\prime}_{i}$ be obtained from $P_{i}$ by dropping some subset $X\subseteq O\backslash\omega_{i}$. Suppose that $X=\left\{x_{1},x_{2},\dots,x_{k}\right\}$, where $x_{1}P_{i}x_{2}P_{i}\cdots P_{i}x_{k}$. Then $P^{\prime}_{i}$ is obtained from a sequence of $k$ drop strategies. That is, $P^{\prime}_{i}=P_{i}^{k}$, where $P_{i}^{0}=P_{i}$ and $P_{i}^{1},\dots,P_{i}^{k}$ are such that, for each $\ell\in\left\{1,\dots,k\right\}$, $P_{i}^{\ell}$ is obtained from $P_{i}^{\ell-1}$ by dropping object $x_{\ell}$.
Claim 1.
For each $\ell\in\left\{1,\dots,k\right\}$, $\varphi_{i}\left(P_{i}^{\ell-1},P_{-i}\right)R_{i}\varphi_{i}\left(P_{i}^{\ell},P_{-i}\right)$.
Proof of Claim 1.
The proof is by induction on $\ell$. Clearly, $\varphi_{i}\left(P\right)=\varphi_{i}\left(P_{i}^{0},P_{-i}\right)R_{i}\varphi_{i}\left(P_{i}^{1},P_{-i}\right)$ by drop strategy-proofness. For the inductive step, suppose that $\ell\in\left\{1,\dots,k-1\right\}$ is such that
$\varphi_{i}\left(P\right)=\varphi_{i}\left(P_{i}^{0},P_{-i}\right)R_{i}\varphi_{i}\left(P_{i}^{1},P_{-i}\right)R_{i}\cdots R_{i}\varphi_{i}\left(P_{i}^{\ell},P_{-i}\right).$
It suffices to show that $\varphi_{i}\left(P_{i}^{\ell},P_{-i}\right)R_{i}\varphi_{i}\left(P_{i}^{\ell+1},P_{-i}\right)$. By drop strategy-proofness, we have $\varphi_{i}\left(P_{i}^{\ell},P_{-i}\right)R_{i}^{\ell}\varphi_{i}\left(P_{i}^{\ell+1},P_{-i}\right)$. By the weak endowment lower bound, $\varphi_{i}\left(P_{i}^{\ell},P_{-i}\right)\subseteq O\backslash\left\{x_{1},\dots,x_{\ell}\right\}$ and $\varphi_{i}\left(P_{i}^{\ell+1},P_{-i}\right)\subseteq O\backslash\left\{x_{1},\dots,x_{\ell+1}\right\}\subseteq O\backslash\left\{x_{1},\dots,x_{\ell}\right\}$. Because $P_{i}^{\ell}$ agrees with $P_{i}$ on $O\backslash\left\{x_{1},\dots,x_{\ell}\right\}$, we have $\varphi_{i}\left(P_{i}^{\ell},P_{-i}\right)R_{i}\varphi_{i}\left(P_{i}^{\ell+1},P_{-i}\right)$, as desired. ∎
It follows from Claim 1 that
$\varphi_{i}\left(P\right)=\varphi_{i}\left(P_{i}^{0},P_{-i}\right)R_{i}\varphi_{i}\left(P_{i}^{1},P_{-i}\right)R_{i}\cdots R_{i}\varphi_{i}\left(P_{i}^{k},P_{-i}\right).$
Therefore, $\varphi_{i}\left(P\right)R_{i}\varphi_{i}\left(P_{i}^{k},P_{-i}\right)$. $\blacksquare$
Proof of Proposition 2.12
Suppose that $\varphi$ satisfies the stated properties. Let $P\in\mathcal{P}^{N}$ and $i\in N$. Let $P^{\prime}_{i}\in{\mathcal{T}}\left(P_{i}\right)$ be a truncation strategy such that all non-endowed objects are dropped, i.e., for each $o\in O\setminus\omega_{i}$, $w_{i}(P_{i})\mathbin{P^{\prime}_{i}}o$. By the weak endowment lower bound, $\varphi_{i}\left(P^{\prime}_{i},P_{-i}\right)\subseteq\omega_{i}$. Because $\varphi$ is balanced, we must have $\left|\varphi_{i}\left(P^{\prime}_{i},P_{-i}\right)\right|=\left|\omega_{i}\right|$ and, hence, $\varphi_{i}\left(P^{\prime}_{i},P_{-i}\right)=\omega_{i}$. It follows from truncation-proofness that $\varphi_{i}\left(P\right)R_{i}\varphi_{i}\left(P^{\prime}_{i},P_{-i}\right)=\omega_{i}$. $\blacksquare$
Proof of Theorem 3.2
Toward contradiction, suppose that $\varphi$ satisfies the properties but $\varphi\neq\varphi^{\text{TTC}}$.
For each $P\in\mathcal{P}^{N}$, and each $t\in\mathbb{N}$, let $\mathcal{C}_{t}(P)$ be the set of cycles that obtain at step $t$ of $\text{TTC}\left(P\right)$. Similarly, let $C_{t}(P)\in\mathcal{C}_{t}(P)$ be the cycle that is executed at step $t$ of $\text{TTC}\left(P\right)$.
Denote the size of a profile $P$ by $s\left(P\right)=\sum_{i\in N}\left|\left\{o\in O\mid oR_{i}\bot_{P_{i}}\left(\omega_{i}\right)\right\}\right|$. The similarity $\rho$ is defined exactly as in the proof of Theorem 3.5.
Let $P$ be a profile such that for any other profile $\tilde{P}\neq P$, either
  • (1)
    $\rho(P)<\rho(\tilde{P})$, or
  • (2)
    $\rho(P)=\rho(\tilde{P})$ and $s(P)\leq s(\tilde{P})$.
Let $t\coloneqq\rho(P)$. Since $\rho(P)=t$, $\varphi(P)$ executes the cycles $C_{1}(P),C_{2}(P),\dots,C_{t-1}(P)$ but not $C_{t}(P)$. Let $C\coloneqq C_{t}(P)$ and $O^{t}\coloneqq O\backslash\bigcup_{\tau=1}^{t-1}O\left(C_{\tau}\left(P\right)\right)$. Suppose that
$C=\left(i_{1},o_{i_{2}},i_{2},o_{i_{3}},\dots,i_{k-1},o_{i_{k}},i_{k},o_{i_{k+1}},i_{k+1}=i_{1}\right).$
Because $\varphi(P)$ does not execute $C$, there is an agent $i_{\ell}\in N(C)$ such that, although agent $i_{\ell}$ points to $o_{i_{\ell+1}}$ on $C$, she does not receive $o_{i_{\ell+1}}$ at $\varphi(P)$, i.e., $o_{i_{\ell+1}}\notin\varphi_{i}(P)$. Note that, by the definition of $\text{TTC}(P)$, $o_{i_{\ell+1}}$ is agent $i_{\ell}$’s most-preferred object in $O^{t}$ at $P_{i_{\ell}}$, i.e., $o_{i_{\ell+1}}=\top_{P_{i_{\ell}}}\left(O^{t}\right)$. Without loss of generality, let $i_{\ell}=i_{k}$. Thus, $o_{i_{1}}\notin\varphi_{i_{k}}(P)$.
Let $X_{i_{k}}$ be the set of all objects in $O\backslash O^{t}$ that are assigned to agent $i_{k}$, i.e., $X_{i_{k}}=\varphi_{i_{k}}(P)\backslash O^{t}$. We next show that, apart from the objects in $X_{i}$, agent $i_{k}$ only receives the remainder of her endowment.
Claim 2.
$\varphi_{i_{k}}(P)\cap O^{t}=\omega_{i_{k}}\cap O^{t}$.
Proof of Claim 2.
Suppose otherwise. By balancedness and the fact that $\left|\varphi_{i_{k}}\left(P\right)\backslash O^{t}\right|=\left|\omega_{i_{k}}\backslash O^{t}\right|$, we must have $\left|\varphi_{i_{k}}\left(P\right)\cap O^{t}\right|=\left|\omega_{i_{k}}\cap O^{t}\right|$. Consequently, $\varphi_{i_{k}}\left(P\right)\cap O^{t}\neq\omega_{i_{k}}\cap O^{t}$ implies there is an object $o^{\prime}\in\varphi_{i_{k}}(P)\cap O^{t}$ with $o^{\prime}\not\in\omega_{i_{k}}$.
By the definition of $\text{TTC}(P)$, we know that $o_{i_{1}}$ is agent $i_{k}$’s most-preferred object in $O^{t}$, i.e., for each $o\in O^{t}$, $o_{i_{1}}\mathbin{R_{i_{k}}}o$. Because $o^{\prime}\in\varphi_{i_{k}}(P)\cap O^{t}$ and $o_{i_{1}}\notin\varphi_{i_{k}}(P)$, we have $o_{i_{1}}\mathbin{P_{i_{k}}}o^{\prime}$. By the weak endowment lower bound,
$o^{\prime}\mathbin{P_{i_{k}}}w_{i_{k}}(P_{i_{k}}).$
Let $P^{\prime}_{i_{k}}$ be the truncation of $P_{i_{k}}$ at $o_{i_{1}}$. That is, $P^{\prime}_{i_{k}}$ is obtained from $P_{i_{k}}$ by dropping the set $\{o\in O\backslash\omega_{i_{k}}\mid o_{i_{1}}\mathbin{P_{i_{k}}}o\}$. [^37] If we use drop strategy-proofness, we can consider another preference $P^{\prime}_{i_{k}}$ obtained from $P_{i_{k}}$ by dropping object $o^{\prime}$. Let $P^{\prime}\coloneqq(P^{\prime}_{i_{k}},P_{-i_{k}})$.
By the definition of $\text{TTC}(P)$, we see that until step $t$, all top trading cycles that are executed via TTC at $P$ and $P^{\prime}$ are the same, i.e., $C_{\tau}(P)=C_{\tau}(P^{\prime})$ for $\tau=1,\ldots,t$. By the selection of $P$, we know that $\rho(P)=t\leq\rho(P^{\prime})$. Thus, all cycles in $\{C_{1}(P),\dots,C_{t-1}(P)\}$ are executed at $\varphi(P^{\prime})$. Therefore,
$X_{i_{k}}\subseteq\varphi_{i_{k}}(P^{\prime}).$
Since $\varphi$ is truncation-proof, we know that $i_{k}$ cannot be better off by misreporting $P^{\prime}_{i_{k}}$ at $P$. Together with $X_{i_{k}}\subseteq\varphi_{i_{k}}(P^{\prime})$ and $o_{i_{k}}\notin\varphi_{i_{k}}(P)$, we conclude that $i_{k}$ cannot receive $o_{i_{1}}$ at $\varphi(P^{\prime})$, i.e., $o_{i_{1}}\notin\varphi_{i_{k}}(P^{\prime})$. Therefore, $C$ is not executed at $\varphi(P^{\prime})$, which means that $\rho(P^{\prime})\leq t$. As a result, we find that $\rho(P^{\prime})=t$. However, since $P^{\prime}_{i_{k}}$ is obtained from $P_{i_{k}}$ by dropping the subset $\{o\in O\backslash\omega_{i_{k}}\mid o_{i_{1}}\mathbin{P_{i_{k}}}o\}$, and $o_{i_{1}}\mathbin{P_{i_{k}}}o^{\prime}\mathbin{P_{i_{k}}}w_{i_{k}}(P_{i_{k}})$, we see that $s(P^{\prime})<s(P)$. This contradicts the choice of $P$. ∎
Next, we show that at least two agents are involved in $C$.
Claim 3.
$|N(C)|>1$.
Proof.
Toward contradiction, suppose that $|N(C)|=1$, that is to say, $i_{1}=i_{k}$ and $C=(i_{1},o_{i_{1}},i_{1})$. Then $o_{i_{1}}\in\omega_{i_{k}}$ and Claim 2 implies that $o_{i_{1}}\in\varphi_{i_{k}}(P)$, a contradiction. ∎
By Claim 2 and Claim 3, agent $i_{k-1}$ points to $o_{i_{k}}$ on $C$, yet she does not receive $o_{i_{k}}$. An argument similar to the proof of Claim 2 implies that $\varphi_{i_{k-1}}(P^{\prime})\cap O^{t}=\omega_{i_{k-1}}\cap O^{t}$.
Proceeding by induction, we conclude that for each agent $i_{\ell}$ that is involved in $C$, we have $\varphi_{i_{\ell}}(P)\cap O^{t}=\omega_{i_{\ell}}\cap O^{t}$.
However, this means that $\varphi$ is not ig-efficient (or Pareto efficient), as agents in $N(C)$ can benefit by execution of $C$. This contradiction completes the proof of Theorem 3.2. $\blacksquare$
Proof of Proposition 4.3
Let $P\in\mathcal{R}^{N}$ be any profile of responsive preferences. Let $i\in N$, $x\in O\setminus\omega_{i}$, and let $P^{\prime}_{i}$ be a truncation of $P_{i}$ at $x$. Let $P^{\prime}\coloneqq(P^{\prime}_{i},P_{-i})$. Let $U\coloneqq\{o\in O\mid o\mathbin{R_{i}}x\}$. There are two cases.
Case 1:
Suppose $\varphi_{i}^{\text{TTC}}(P)\subseteq U$. Then, by the definition of TTC, we know that all top trading cycles that are obtained at $P$ and $P^{\prime}$ are exactly the same. As a result, $\varphi_{i}^{\text{TTC}}(P)=\varphi_{i}^{\text{TTC}}(P^{\prime})$. So agent $i$ is not better off by misreporting $P^{\prime}_{i}$.
Case 2:
Suppose $\varphi_{i}^{\text{TTC}}(P)\nsubseteq U$. Then agent $i$ receives at least one object in $O\setminus U$ at $\varphi^{\text{TTC}}(P)$. Assume that at $\text{TTC}(P)$, the earliest step at which agent $i$ receives an object in $O\setminus U$ is $t$.
Let $O_{t}\subseteq O$ be the set of objects that are remaining at step $t$ of $\text{TTC}(P)$. Since $\succ^{P_{i}}$ and $\succ^{P^{\prime}_{i}}$ agree on $U$, [^38] Formally, $\succ^{P_{i}}\cap(U\times U)=\succ^{P^{\prime}_{i}}\cap(U\times U)$. we know that the top trading cycles that are executed at steps $11,\dots,t-$ 1 of $\text{TTC}(P)$ and $\text{TTC}(P^{\prime})$ are the same. Thus, at step $t$ of $\text{TTC}(P^{\prime})$, the set of remaining objects is also $O_{t}$. Moreover, we have
(6) $\varphi^{\text{TTC}}_{i}(P)\setminus O_{t}=\varphi^{\text{TTC}}_{i}(P^{\prime})\setminus O_{t}.$
By the definition of $P^{\prime}_{i}$, we know that all objects in $\{o\in O\setminus\omega_{i}\mid x\mathbin{P_{i}}o\}$ are ranked below $i$’s worst endowed object at $\succ^{P^{\prime}_{i}}$ ($i$’s worst endowed object is the same at $P_{i}$ and $P^{\prime}_{i}$). Thus, by the definition of TTC, from step $t$ of $\text{TTC}(P^{\prime})$, agent $i$ only points to (and hence receives) her endowed objects. That is,
(7) $\varphi^{\text{TTC}}_{i}(P^{\prime})\cap O_{t}=\omega_{i}\cap O_{t}.$
Again, by the definition of TTC, from step $t$ of $\text{TTC}(P)$, agent $i$ only points to (and hence receives) some objects that are weakly better (with respect to $\succ^{P_{i}}$) than her endowed objects. In other words, there is a bijection $\sigma:\omega_{i}\cap O_{t}\to\varphi^{\text{TTC}}_{i}(P)\cap O_{t}$ such that for each $o\in\omega_{i}\cap O_{t}$, $\sigma(o)\mathbin{\succeq^{P_{i}}}o$. Together with (6) and (7), we conclude that there is a bijection $\pi:\varphi_{i}^{\text{TTC}}(P^{\prime})\to\varphi_{i}^{\text{TTC}}(P)$ such that
(8) for each $\text{for each }o\in\varphi^{\text{TTC}}_{i}(P^{\prime}),\pi(o)\mathbin{\succeq^{P_{i}}}o.$
Finally, recall that $i$’s original preference relation $P_{i}$ is responsive. Therefore, we must have $\varphi^{\text{TTC}}_{i}(P)\mathbin{R_{i}}\varphi^{\text{TTC}}_{i}(P^{\prime})$. [^39] Starting from $\varphi^{\text{TTC}}_{i}(P^{\prime})$, replace each object $o\in\varphi^{\text{TTC}}_{i}(P^{\prime})$ with object $\pi(o)$, one at a time, and apply the definition of responsiveness. Thus, agent $i$ is not better off by misreporting $P^{\prime}_{i}$. $\blacksquare$
A.1. Proof of Lemma 4.5
Suppose $\varphi$ is an individual-good-based rule satisfied balancedness and individual rationality. Let $P\in\mathcal{R}^{N}$ be a profile of responsive preferences, and suppose $\mu\coloneqq\varphi(P)$ violates the weak endowment lower bound at $P$. Then there is some agent $i\in N$ such that $\mu_{i}\nsubseteq\{o\in O\mid o\mathbin{R_{i}}\bot_{P_{i}}(\omega_{i})\}$. By balancedness, we must have $|\mu_{i}|=|\omega_{i}|$. Because $\bot_{P_{i}}(\mu_{i})$ is worse than $\bot_{P_{i}}(\omega_{i})$ at $P_{i}$, there is a responsive preference relation $P^{\prime}_{i}\in\mathcal{R}$ such that $\succ^{P^{\prime}_{i}}=\succ^{P_{i}}$ and $\omega_{i}\mathbin{P^{\prime}_{i}}\mu_{i}$. By individual rationality, we must have $\varphi(P^{\prime}_{i},P_{-i})\neq\mu_{i}$. However, $\varphi(P^{\prime}_{i},P_{-i})=\varphi(P)=\mu$ because $\varphi$ is “individual-good-based.” This is a contradiction.
A.2. Proof of Proposition 4.12
Let $P\in\mathcal{CL}^{N}$, $i\in N$, and let $\overline{P}_{i}\in\mathcal{D}\left(P_{i}\right)$. Let $\overline{P}\coloneqq\left(\overline{P}_{i},P_{-i}\right)$ and suppose that $\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)R_{i}\varphi_{i}^{\text{ATTC}}\left(P\right)$. It suffices to show that $\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)=\varphi_{i}^{\text{ATTC}}\left(P\right)$.
Let $\varphi_{i}^{\text{ATTC}}\left(P\right)=\left\{x_{1},\dots,x_{m}\right\}\eqqcolon X$. By relabeling the objects if necessary, we can assume that, for each $k\in\{1,\dots,m\}$, $x_{k}$ is the $k$th object assigned to agent $i$ during $\text{ATTC}\left(P\right)$. Let $\tau_{i}=\tau_{P_{i}}$ denote the LP tree associated with $P_{i}$. For each $k\in\left\{1,\dots,m\right\}$, let $v_{k}$ denote the unique vertex of $\tau_{i}\left(X\right)$ whose label is $x_{k}$. This means that
for $\text{for each }k\in\left\{1,\dots,m\right\},\quad X\cap a\left(v_{k}\right)=\left\{x_{1},\dots,x_{k}\right\}.$
We show by induction that
for $\text{for each }k\in\left\{1,\dots,m\right\},\quad\overline{X}\cap a\left(v_{k}\right)=\left\{x_{1},\dots,x_{k}\right\}.$
Base case: $k=1$.
Consider any object $x\in a\left(v_{1}\right)\backslash\text{$\left\{x_{1}\right\}$}$. That is, $x$ labels one of the (strict) ancestors of $v_{1}$ in $\tau_{i}\left(X\right)$, and $x\notin\varphi_{i}^{\text{ATTC}}\left(P\right)$. This means that, although agent $i$ may point to $x$ during $\text{ATTC}\left(P\right)$, agent $i$ is not on the cycle of $\text{ATTC}\left(P\right)$ that contains object $x$. It follows that agent $i$ does not belong to the cycle of $\text{ATTC}\left(\overline{P}\right)$ that contains object $x$ either. Hence, $x\notin\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)$.
Now consider object $x_{1}$. Because $\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)R_{i}\varphi_{i}^{\text{ATTC}}\left(P\right)$, we must have $x_{1}\in\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)$. Consequently, $\overline{X}\cap a\left(v_{1}\right)=\left\{x_{1}\right\}$.
Inductive step.
Let $k\in\left\{1,\dots,m-1\right\}$ be such that for each $\ell\in\left\{1,\dots,k\right\}$, $\overline{X}\cap a\left(v_{\ell}\right)=\left\{x_{1},\dots,x_{\ell}\right\}$. We show that $\overline{X}\cap a\left(v_{k+1}\right)=\left\{x_{1},\dots,x_{k+1}\right\}$.
Consider any object $x\in\left(a\left(v_{k+1}\right)\backslash\left\{x_{k+1}\right\}\right)\backslash a\left(v_{k}\right)$. That is, $x$ labels a vertex that appears after $v_{k}$ and before $v_{k+1}$ on $\tau_{i}\left(X\right)$, and $x\notin\varphi_{i}^{\text{ATTC}}\left(P\right)$. This means that, although agent $i$ may point to $x$ during $\text{ATTC}\left(P\right)$, agent $i$ is not on the cycle of $\text{ATTC}\left(P\right)$ that contains object $x$. It follows that agent $i$ does not belong to the cycle of $\text{ATTC}\left(\overline{P}\right)$ that contains object $x$ either. Hence, $x\notin\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)$.
Now consider object $x_{k+1}$. Because $\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)R_{i}\varphi_{i}^{\text{ATTC}}\left(P\right)$, we must have $x_{k+1}\in\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)$. Consequently, $\overline{X}\cap a\left(v_{k+1}\right)=\left\{x_{1},\dots,x_{k+1}\right\}$.
By the principle of induction, it then follows that $\overline{X}\cap a\left(v_{m}\right)=\left\{x_{1},\dots,x_{m}\right\}$. Because $\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)$ and $\varphi_{i}^{\text{ATTC}}\left(P\right)$ contain the same number of objects, we must have $\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)=\varphi_{i}^{\text{ATTC}}\left(P\right)$, as we needed to show. $\blacksquare$
A.3. Proof of Theorem 4.13
The proof here is almost identical to the proof of Theorem 3.2. The main difference is that, since we focus on conditionally lexicographic preferences, any statements that pertained to an agent’s most-preferred object in Theorem 3.2 now pertain to an agent’s most-preferred object conditional on receiving some set of objects.
Toward contradiction, suppose that $\varphi$ satisfies the properties but $\varphi\neq\varphi^{ATTC}$.
For each $P\in\mathcal{P}^{N}$, and each $t\in\mathbb{N}$, let $\mathcal{C}_{t}(P)$ be the set of cycles that obtain at step $t$ of $\text{ATTC}\left(P\right)$. Similarly, let $C_{t}(P)\in\mathcal{C}_{t}(P)$ be the cycle that is executed at step $t$ of $\text{ATTC}\left(P\right)$.
Denote the size of a profile $P$ by $\sum_{i\in N}\sum_{X\in 2^{O}}\left|a\left(w_{P_{i}}\left(\omega_{i}\mid X\right)\right)\right|$. The similarity $\rho$ is defined exactly as in the proof of Theorem 3.5.
Let $P$ be a profile such that for any other profile $\tilde{P}\neq P$, either
  • (1)
    $\rho(P)<\rho(\tilde{P})$, or
  • (2)
    $\rho(P)=\rho(\tilde{P})$ and $s(P)\leq s(\tilde{P})$.
Let $t\coloneqq\rho(P)$. Since $\rho(P)=t$, $\varphi(P)$ executes the cycles $C_{1}(P),C_{2}(P),\dots,C_{t-1}(P)$ but not $C_{t}(P)$. Let $C\coloneqq C_{t}(P)$ and $O^{t}\coloneqq O\backslash\bigcup_{\tau=1}^{t-1}O\left(C_{\tau}\left(P\right)\right)$. Suppose that
$C=\left(i_{1},o_{i_{2}},i_{2},o_{i_{3}},\dots,i_{k-1},o_{i_{k}},i_{k},o_{i_{k+1}},i_{k+1}=i_{1}\right).$
Because $\varphi(P)$ does not execute $C$, there is an agent $i_{\ell}\in N(C)$ such that, although agent $i_{\ell}$ points to $o_{i_{\ell+1}}$ on $C$, she does not receive $o_{i_{\ell+1}}$ at $\varphi(P)$, i.e., $o_{i_{\ell+1}}\notin\varphi_{i}(P)$. Note that, by the definition of $\text{ATTC}(P)$, $o_{i_{\ell+1}}$ is agent $i_{\ell}$’s most-preferred object in $O^{t}$ at $P_{i_{\ell}}$ conditional on receiving $\varphi_{i_{\ell}}(P)\setminus O^{t}$, i.e., $o_{i_{\ell+1}}=\top_{P_{i_{\ell}}}(O^{t}\mid\varphi_{i_{\ell}}(P)\setminus O^{t})$. Without loss of generality, let $i_{\ell}=i_{k}$. Thus, $o_{i_{1}}\notin\varphi_{i_{k}}(P)$.
Let $X_{i_{k}}$ be the set of all objects in $O\backslash O^{t}$ that are assigned to agent $i_{k}$, i.e., $X_{i_{k}}=\varphi_{i_{k}}(P)\backslash O^{t}$. We next show that, apart from the objects in $X_{i}$, agent $i_{k}$ only receives the remainder of her endowment.
Claim 4.
$\varphi_{i_{k}}(P)\cap O^{t}=\omega_{i_{k}}\cap O^{t}$.
Proof of Claim 4.
Suppose otherwise. By balancedness and the fact that $\left|\varphi_{i_{k}}\left(P\right)\backslash O^{t}\right|=\left|\omega_{i_{k}}\backslash O^{t}\right|$, we must have $\left|\varphi_{i_{k}}\left(P\right)\cap O^{t}\right|=\left|\omega_{i_{k}}\cap O^{t}\right|$. Consequently, $\varphi_{i_{k}}\left(P\right)\cap O^{t}\neq\omega_{i_{k}}\cap O^{t}$ implies there is an object $o^{\prime}\in\varphi_{i_{k}}(P)\cap O^{t}$ with $o^{\prime}\not\in\omega_{i_{k}}$. By the weak endowment lower bound,
$o^{\prime}\in a\left(w_{P_{i_{k}}}(\omega_{i_{k}}\mid\varphi_{i_{k}}(P))\right).$
By the definition of $\text{ATTC}(P)$, we know that $o_{i_{1}}$ is agent $i_{k}$’s most-preferred object in $O^{t}$ conditional on receiving $X_{i_{k}}$, i.e., $o_{i_{1}}=\top_{P_{i_{k}}}(O^{t}\mid X_{i_{k}})$. Because $o^{\prime}\in\varphi_{i_{k}}(P)\cap O^{t}$ and $o_{i_{1}}\notin\varphi_{i_{k}}(P)$, we have $o_{i_{1}}\neq o^{\prime}$.
Let $P^{\prime}_{i_{k}}$ be the drop strategy for $P_{i_{k}}$ obtained by dropping object $x$. Let $P^{\prime}\coloneqq(P^{\prime}_{i_{k}},P_{-i_{k}})$. By the definition of TTC, we see that until step $t$, all top trading cycles that are executed via TTC at $P$ and $P^{\prime}$ are the same, i.e., $C_{\tau}(P)=C_{\tau}(P^{\prime})$ for $\tau=1,\ldots,t$. By the selection of $P$, we know that $\rho(P)=t\leq\rho(P^{\prime})$. Thus, all cycles in $\{C_{1}(P),\dots,C_{t-1}(P)\}$ are executed at $\varphi(P^{\prime})$. Therefore,
$X_{i_{k}}\subseteq\varphi_{i_{k}}(P^{\prime}).$
Since $\varphi$ is weakly drop strategy-proof, we know that $i_{k}$ cannot be better off by misreporting $P^{\prime}_{i_{k}}$ at $P$. Together with $X_{i_{k}}\subseteq\varphi_{i_{k}}(P^{\prime})$ and $o_{i_{k}}\notin\varphi_{i_{k}}(P)$, we conclude that $i_{k}$ cannot receive $o_{i_{1}}$ at $\varphi(P^{\prime})$, i.e., $o_{i_{1}}\notin\varphi_{i_{k}}(P^{\prime})$. Therefore, $C$ is not executed at $\varphi(P^{\prime})$, which means that $\rho(P^{\prime})\leq t$. As a result, we find that $\rho(P^{\prime})=t$. However, since $P^{\prime}_{i_{k}}$ is obtained from $P_{i_{k}}$ by dropping an object $o^{\prime}\in a\left(w_{P_{i_{k}}}(\omega_{i_{k}}\mid\varphi_{i_{k}}(P))\right)$, we see that $s(P^{\prime})<s(P)$. This contradicts the choice of $P$. ∎
Next, we show that at least two agents are involved in $C$.
Claim 5.
$|N(C)|>1$.
Proof.
Toward contradiction, suppose that $|N(C)|=1$, that is to say, $i_{1}=i_{k}$ and $C=(i_{1},o_{i_{1}},i_{1})$. Then $o_{i_{1}}\in\omega_{i_{k}}$ and Claim 4 implies that $o_{i_{1}}\in\varphi_{i_{k}}(P)$, a contradiction. ∎
By Claim 4 and Claim 5, agent $i_{k-1}$ points to $o_{i_{k}}$ on $C$, yet she does not receive $o_{i_{k}}$. An argument similar to the proof of Claim 4 implies that $\varphi_{i_{k-1}}(P^{\prime})\cap O^{t}=\omega_{i_{k-1}}\cap O^{t}$.
Proceeding by induction, we conclude that for each agent $i_{\ell}$ that is involved in $C$, we have $\varphi_{i_{\ell}}(P)\cap O^{t}=\omega_{i_{\ell}}\cap O^{t}$.
However, this means that $\varphi$ is not ig-efficient, as agents in $N(C)$ can benefit by execution of $C$. This contradiction completes the proof of Theorem 4.13. $\blacksquare$
A.4. Proof of Proposition 5.3
Let $P_{1}\in\mathcal{\overline{P}}\setminus\mathcal{CL}$. By Proposition 4.9, there exist disjoint subsets $X,Y\in 2^{O}$ with $X\neq\emptyset$ such that
for all $\text{for all }x\in X\text{, there exists }Z_{x}\subseteq X\setminus x\text{ such that }(Y\cup Z_{x})\mathbin{P_{1}}(Y\cup x).$
Because $P_{1}$ is monotone, we have
for all $\text{for all }x\in X,\quad[Y\cup(X\setminus x)]\mathbin{P_{1}}(Y\cup x).$
Note also that $|X|\geq 3.$ [^40] If $X$ is a singleton, say $X=\{x\}$, then $Y\mathbin{P_{1}}(Y\cup x)$, a violation of monotonicity. If $X$ contains two objects, say $X=\{x,y\}$, we would have $(Y\cup x)\mathbin{P_{1}}(Y\cup y)\mathbin{P_{1}}(Y\cup x)$, a violation of transitivity.
Let $x^{*}\in X$ be such that
for all $\text{for all }x\in X,\quad(Y\cup x^{*})\mathbin{R_{1}}(Y\cup x).$
Then it follows that
for all $\text{for all }x\in X,\quad[Y\cup(X\setminus x)]\mathbin{P_{1}}(Y\cup x^{*})\mathbin{R_{1}}(Y\cup x).$
Case 1:
Suppose $X\cup Y=O$. Let $N=\{1,2\}$, and let $P_{2}\in\mathcal{L}$ be such that (i) $\top_{P_{2}}(O)=x^{*}$, and (ii) for all $x\in X$, $x\mathbin{P_{2}}Y$. Then, in particular,
$x^{*}\mathbin{P_{2}}[Y\cup(X\setminus x^{*})]\mathbin{R_{2}}(X\setminus x^{*}).$
Consider the allocations
$\mu\coloneqq(Y\cup x^{*},X\backslash x^{*})\quad\text{and}\quad\overline{\mu}\coloneqq(Y\cup\left(X\backslash x^{*}\right),x^{*}).$
Then $\mu$ is not Pareto efficient at $P$ because it is Pareto-dominated by $\overline{\mu}$ at $P$.
We claim that $\mu$ is ig-efficient at $P$. Indeed, consider any single-object exchange identified by the cycle $C=(1,x,2,y,1)$, where $x\in\mu_{2}$ and $y\in\mu_{1}$. Executing the exchange results in the allocation $\mu^{\prime}$ where
$\mu^{\prime}_{1}=\left(Y\cup\left\{x^{*},x\right\}\right)\backslash y\quad\text{and}\quad\mu^{\prime}_{2}=\left(X\backslash\left\{x^{*},x\right\}\right)\cup y.$
If $y\in Y$, then this exchange harms agent 2, i.e., $\mu_{2}\mathbin{P_{2}}\mu^{\prime}_{2}$. On the other hand, if $y=x^{*}$, then this exchange harms agent 1 because $\mu_{1}=(Y\cup x^{*})\mathbin{P_{1}}(Y\cup x)=\mu^{\prime}_{1}$. Hence, $\mu$ is ig-efficient.
Case 2:
Suppose $X\cup Y\subsetneq O$. Let $N=\{1,2,3\}$ and denote $\overline{O}\coloneqq O\setminus(X\cup Y)$. Let $P_{2}\in\mathcal{L}$ be such that (i) $\top_{P_{2}}(O)=x^{*}$, and (ii) for all $x\in X$, $x\mathbin{P_{2}}(O\setminus X)$. Let $P_{3}\in\mathcal{L}$ be such that (i) for all $z\in\overline{O}$, $z\mathbin{P_{3}}(X\cup Y)$.
Consider the allocations
$\mu\coloneqq(Y\cup x^{*},X\backslash x^{*},\overline{O})\quad\text{and}\quad\overline{\mu}\coloneqq(Y\cup\left(X\backslash x^{*}\right),x^{*},\overline{O}).$
Then $\mu$ is not Pareto efficient at $P$ because it is Pareto-dominated by $\overline{\mu}$ at $P$.
We claim that $\mu$ is ig-efficient at $P$. Because $\mu_{3}$ is agent 3’s top-ranked $|\overline{O}|$-subset of $O$, any single-object exchange involving agent 3 must harm agent 3. Thus, any Pareto-improving single-object exchange must involve only agents 1 and 2. However, there is no such exchange by the argument in Case 1. Hence, $\mu$ is ig-efficient. $\blacksquare$
Appendix B Examples
B.1. Independence of Properties in Theorems 3.2 and  4.13
We establish the independence of the properties in Theorem 3.2 by providing examples of rules, different from TTC / ATTC, that violate exactly one of the properties. For each of the examples below we indicate the property that the rule fails (while it satisfies all remaining properties).
Pareto efficiency: the no-trade rule that always selects the initial allocation satisfies all properties except for Pareto efficiency.
The weak endowment lower bound: the serial dictatorships subject to balancedness [^41] That is, each dictator is assigned the same number of objects as her endowment. satisfy all properties except for the weak endowment lower bound.
Balancedness: the serial dictatorships subject to the weak endowment lower bound [^42] That is, each dictator $i$ is assigned her best bundle among all subsets $X$ of the remaining objects that would not violate the weak endowment lower bound (i.e., such that $X\subseteq\{o\in O\mid o\mathbin{R_{i}}\bot_{P_{i}}(\omega_{i})\}$, where $P_{i}$ denotes $i$’s preference relation.) satisfy all properties except for balancedness.
Truncation-proofness / Drop strategy-proofness: the following example is from Altuntaş et al. (2023). Let $N=\{1,2,3,4\}$, and for each $i\in N$, $\omega_{i}=o_{i}$. Let $P^{*}$ be defined as follows (endowment is underlined).
$P^{*}_{1}:o_{2},o_{4},\underline{o_{1}},o_{3};$
$P^{*}_{2}:o_{1},o_{3},\underline{o_{2}},o_{4};$
$P^{*}_{3}:o_{2},o_{4},\underline{o_{3}},o_{1};$
$P^{*}_{4}:o_{1},o_{3},\underline{o_{4}},o_{2}.$
Let $\varphi$ be defined as follows.
$\varphi\left(P\right)=\begin{cases}(o_{4},o_{3},o_{2},o_{1}),&\text{if }P=P^{*}\\ \varphi^{\text{ATTC}}(P),&\text{otherwise.}\end{cases}$
Note that for lexicographic preferences, $\varphi^{\text{ATTC}}=\varphi^{\text{TTC}}$. Altuntaş et al. (2023) show that $\varphi$ satisfies Pareto efficiency and the strong endowment lower bound (hence balancedness and the weak endowment lower bound), but violates drop strategy-proofness / truncation proofness. To see this, note that agent $1$ receives $o_{4}$ at $\varphi\left(P^{*}\right)$, and she can receive her best object $o_{2}$ by misreporting a drop strategy $P^{\prime}_{1}:o_{2},\underline{o_{1}},o_{3},o_{4}$ or a truncation strategy $P^{\prime\prime}_{1}:o_{2},\underline{o_{1}},o_{4},o_{3}$.
B.2. Other examples
Example B.1 (Remark 1).
Suppose $N=\left\{1,2\right\}$ and $\omega=\left(\left\{a,b,c\right\},\left\{x,y,z\right\}\right)$.
Consider the following profile $(P_{1},P_{2})$ ($\omega_{1},\omega_{2}$ are underlined, respectively)
$P_{1}:x,\underline{a},y,\underline{b},z,\underline{c}\quad\text{and}\quad P_{2}:a,\underline{x},b,\underline{y},c,\underline{z}.$
Consider the allocation $\mu=\left(\left\{x,z,c\right\},\left\{a,b,y\right\}\right)$. Clearly, $\mu$ satisfies individual rationality, balancedness, and the weak endowment lower bound. However, $\mu$ violates the strong endowment lower bound because $\mu_{1}$ does not pairwise dominate $\omega_{1}$. In particular, the second-best object in $\omega_{1}$, $b$, is better than the second-best object in $\mu_{1}$, $z$.
Now consider the allocation $\nu=\left(\left\{b,z,c\right\},\left\{a,x,y\right\}\right)$. Then $\nu$ satisfies balancedness and the weak endowment lower bound. However, $\nu$ is not individually rational because $\omega_{1}P_{1}\nu_{1}$.
Finally, consider
$P^{\prime}_{1}:x,\underline{a},\underline{b},\underline{c},y,z.$
We see that at $(P^{\prime}_{1},P_{2})$, $\mu$ satisfies individual rationality, balancedness, but violates the weak endowment lower bound as $z\in\mu_{1}$ is worse than agent $1$’s worst endowed object $c\in\omega_{1}$.  $\diamond$
Example B.2 (A rule that satisfies truncation-proofness, the weak endowment lower bound and individual rationality, but not drop strategy-proofness.).
Suppose $N=\left\{1,2,3\right\}$, and $\omega=\left(o_{1},o_{2},o_{3}\right)$. Let $P^{*}\in{\mathcal{P}}^{N}$ be such that
$P_{1}^{*}:o_{2},o_{3},o_{1},\quad P_{2}^{*}:o_{3},o_{1},o_{2},\quad P_{3}^{*}:o_{1},o_{2},o_{3}.$
Note that for each $i\in N$, $i$ only has two truncation strategies, i.e.,
[…]
Consider the rule $\varphi$ defined for all $P\in{\mathcal{P}}^{N}$ by
$\varphi\left(P\right)=\begin{cases}\omega,&\text{if }P\in{\mathcal{T}}\left(P_{1}^{*}\right)\times{\mathcal{T}}\left(P_{2}^{*}\right)\times{\mathcal{T}}\left(P_{3}^{*}\right)\\ \varphi^{\text{TTC}}\left(P\right),&\text{otherwise}.\end{cases}$
That is, $\varphi\left(P\right)=\omega$ whenever $P$ is such that, for all $i\in N$, $P_{i}$ is a truncation of $P_{i}^{*}$; otherwise, $\varphi\left(P\right)=\varphi^{\text{TTC}}\left(P\right)$. By the definition of $\varphi$, it is easy to see that $\varphi$ is individually rational. It is straightforward to show that $\varphi$ is truncation-proof. To see that $\varphi$ is not drop strategy-proof, consider
$P^{\prime}_{1}:o_{3},o_{1},o_{2}.$
Then $P^{\prime}_{1}$ is obtained from $P_{1}^{*}$ by dropping object $o_{2}\in O\backslash\omega_{1}$. Because
$\varphi_{1}\left(P^{\prime}_{1},P_{-1}^{*}\right)=\varphi_{1}^{\text{TTC}}\left(P^{\prime}_{1},P_{-1}^{*}\right)=o_{3}P_{1}^{*}o_{1}=\varphi_{1}\left(P^{*}\right),$
agent 1 can manipulate at $P^{*}$ by misreporting $P^{\prime}_{1}\in{\mathcal{D}}\left(P_{1}^{*}\right)$. Note that $P^{\prime}_{1}$ is not a truncation of $P_{1}^{*}$.  $\diamond$
Example B.3 (Drop strategy-proofness does not imply truncation-proofness.).
Suppose $N=\left\{1,2,3\right\}$ and $\omega=\left(o_{1},o_{2},o_{3}\right)$. Let $\varphi$ be such that, for all $P\in{\mathcal{P}}^{N}$,
$\varphi\left(P\right)$ =$=\begin{cases}\left(o_{1},o_{2},o_{3}\right),&\text{if }o_{1}\text{ is agent 1's second-ranked object at }P_{1}\\ \left(o_{2},o_{1},o_{3}\right),&\text{if }o_{2}\text{ is agent 1's second-ranked object at }P_{1}\\ \left(o_{3},o_{2},o_{1}\right),&\text{if }o_{3}\text{ is agent 1's second-ranked object at }P_{1}.\end{cases}$
Then $\varphi$ is drop strategy-proof. Indeed, if $P^{\prime}_{1}$ is obtained from $P_{1}$ by dropping any single object, then agent 1 receives her worst object at $\left(P^{\prime}_{1},P_{-1}\right)$, i.e., $\varphi_{1}\left(P^{\prime}_{1},P_{-1}\right)=\bot_{P_{1}}\left(O\right)$.
Consider
$P_{1}:o_{2},o_{3},o_{1}.$
Note that $\varphi_{1}(P)=o_{3}$ for any $P_{-1}$. Next, consider a truncation strategy by dropping $\{o_{2},o_{3}\}$, i.e.,
$P^{\prime}_{1}:o_{1},o_{2},o_{3}.$
We see that $\varphi_{1}\left(P^{\prime}_{1},P_{-1}\right)=o_{2}$. Since $\varphi_{1}\left(P^{\prime}_{1},P_{-1}\right)=o_{2}\mathbin{P_{1}}o_{3}=\varphi_{1}\left(P\right)$, we find that $\varphi$ is not truncation-proof.  $\diamond$