### v4

Abstract
This paper considers reallocation of indivisible objects when agents are endowed with and can consume any bundles. We obtain characterizations of generalized versions of the Top Trading Cycles (TTC) rule on several preference domains. On the lexicographic domain, the TTC rule is uniquely determined by balancedness, Pareto efficiency, the worst endowment lower bound, and either truncation-proofness or drop strategy-proofness. On the more general responsive domain, the TTC rule is the unique individual-good-based rule that satisfies balancedness, individual-good efficiency, truncation-proofness, and either individual rationality or the worst endowment lower bound. On the conditionally lexicographic domain, the augmented TTC rule is characterized by balancedness, Pareto efficiency, the worst endowment lower bound, and drop strategy-proofness. The conditionally lexicographic domain is a maximal domain on which Pareto efficiency coincides with individual-good efficiency. For the housing market introduced by Shapley and Scarf (1974), the TTC rule is characterized by Pareto efficiency, individual rationality, and truncation-proofness.
11footnotetext: University of Melbourne, Australia. Contact: jacob.coreno@unimelb.edu.au.22footnotetext: Dongbei University of Finance and Economics, China. Contact: dfeng@dufe.edu.cn.
Keywords: exchange of indivisible objects; Top Trading Cycles; heuristic manipulation; endowment lower bound.
JEL Classification: C78; D47; D71.
1 Introduction
We consider multi-object reallocation problems without monetary transfers. These problems involve a group of agents, each endowed with a set of heterogeneous and indivisible objects and equipped with strict preferences over bundles of objects. An allocation rule specifies how objects are reassigned based on the agents’ reported preferences. A special case of this framework is when each agent is endowed with a single object, leading to the standard single-object reallocation problem (often called a housing market) introduced by Shapley and Scarf (1974).
Object reallocation problems are ubiquitous. Firms plan shift schedules months in advance, and employees frequently exchange their assigned shifts with one another (Manjunath and Westkamp (2021)). Universities exchange students in programs such as The Tuition Exchange Program in the United States and the Erasmus Student Exchange Program in Europe (Dur and Ünver (2019), Bloch et al. (2020)), and seats in oversubscribed courses are (re)allocated among university students (Bichler et al. (2021), Budish (2011), Budish and Cantillon (2012)). Living-donor organ exchange programs, which facilitate the reallocation of organs (e.g., kidneys and livers) among patient-donor pairs, are an important example featuring single-object exchange (Roth et al. (2004, 2005), Ergin et al. (2020)).
In contrast with single-object reallocation, which admits many positive results, [^1] A central result pertinent to the present study is that Gale’s Top Trading Cycles rule (Shapley and Scarf, 1974) is the unique rule satisfying Pareto efficiency, individual rationality, and strategy-proofness (Ma, 1994). multi-object reallocation presents significant challenges for both practitioners and theorists. One practical challenge is the vast number of feasible bundles, which makes it difficult or impossible for agents to accurately report (or even know) their preferences over bundles. [^2] For instance, in a shift reallocation problem with only 20 shifts, there are (${20\choose 5}=15,504$ bundles composed of five shifts. As Roth (2015, p. 331) explains, “a practical mechanism must simplify the language in which preferences can be reported, and by doing so it will restrict which preferences can be reported.” A prominent example is the National Resident Matching Program (NRMP), which matches doctors to hospitals in the United States. Each hospital reports only its rank-order list over individual doctors, even though it may have rather complex preferences over sets of doctors (see, e.g., Roth and Peranson (1999), Milgrom (2009, 2011)).
Motivated by these considerations, we focus on individual-good-based rules—allocation rules with a simple reporting language that consists of rank-order lists over individual objects. Individual-good-based rules are particularly appealing in environments where objects are substitutable. For instance, if agents have lexicographic preferences, [^3] Preferences are lexicographic if the relative ranking between two bundles depends primarily on the most-preferred objects in each of them. then individual-good-based rules are without loss of generality since an agent’s preferences can be succinctly represented by a rank-order list over individual objects. More generally, when agents have responsive preferences, [^4] Preferences are responsive if the relative ranking between two objects does not depend on the other objects they are obtained with. an agent’s preferences over individual objects capture much—though not all—of the relevant information about her preferences over bundles of objects. [^5] An obvious limitation of individual-good-based rules is that they do not allow agents to express information about complex preferences exhibiting complementarities, as discussed in Budish and Kessler (2022). Recognizing this limitation, we explore an alternative reporting language in Section 6 that accommodates complementarities.
A significant theoretical challenge is the inherent conflict among three criteria of interest: efficiency, individual rationality, and strategic robustness. This conflict manifests in various impossibility results. For example, Sönmez (1999) shows that the ideal properties—Pareto efficiency, individual rationality, and strategy-proofness—are incompatible whenever at least one agent is endowed with more than one object. Much of the literature sidesteps this issue by focusing on the strategy-proof rules that fulfill only one of the remaining two ideals. Loosely speaking, the only strategy-proof allocation rules satisfying Pareto efficiency are dictatorial (Pápai (2001), Klaus and Miyagawa (2002), Ehlers and Klaus (2003), Hatfield (2009)), while the only strategy-proof rules satisfying individual rationality are Segmented Trading Cycles rules (Pápai (2003); see also Pápai (2007)). Such rules are not suitable for our purposes, as they may severely compromise the third criterion.
In this paper, we circumvent the incompatibility by making small compromises relative to the ideal notions of efficiency and strategic robustness. We study the generalized Top Trading Cycles (TTC) rule, an individual-good-based rule that extends Gale’s TTC rule (Shapley and Scarf, 1974) to multi-object problems. Despite the challenges posed by the incompatibility of our ideal desiderata, we show that the TTC rule performs remarkably well according to all three criteria. In particular, we provide axiomatic characterizations of the TTC rule based on individual-good efficiency, individual rationality, and truncation-proofness.
On the lexicographic domain, the TTC rule is uniquely determined by balancedness, Pareto efficiency, the worst endowment lower bound, and either truncation-proofness (Theorem 1) or drop strategy-proofness (Theorem 2). Extending to the responsive preference domain and focusing on individual-good-based rules, we find that although full Pareto efficiency is unattainable, the TTC rule uniquely satisfies balancedness, individual-good efficiency, truncation-proofness, and either the worst endowment lower bound (Theorem 3) or individual rationality (Theorem 4). In the special case of the housing market, where each agent is endowed with a single object, we characterize the TTC rule through Pareto efficiency, individual rationality, and truncation-proofness (Theorem 5), generalizing the classic result of Ma (1994). Our key properties are discussed in detail below.
Balancedness posits that a rule assigns the same number of objects to each agent as the initial allocation. Balancedness is an inviolable constraint in many practical reallocation problems. In a shift reallocation problem, for example, balancedness may be imposed so that staff can meet mandatory training requirements (Manjunath and Westkamp (2021)). It is also a typical requirement in applications such as student exchange programs (Dur and Ünver (2019)) and time banks (Andersson et al. (2021), Biró et al. (2022a)). In the absence of strict constraints, balancedness captures a limited notion of fairness.
Individual rationality and the worst endowment lower bound are participation guarantees that differ in the promises they make to the agents. Individual rationality guarantees that each agent enjoys a bundle at least as good as her endowment, whereas the worst endowment lower bound ensures that no agent ever receives an object worse than the least-preferred object in her endowment. While the two properties coincide in single-object problems, they are distinct in multi-object problems. Intuitively, the worst endowment lower bound allows agents to veto individual objects by ranking them below the worst object in their endowment.
Individual-good efficiency requires that no group of agents can destabilize the allocation by carrying out a simple exchange where each involved agent trades a single object for a better one. Under lexicographic preferences, individual-good efficiency coincides with Pareto efficiency. However, in the more general setting of responsive preferences, Pareto efficiency becomes difficult to achieve and verify in practice. Specifically, verifying full Pareto efficiency requires considering an exponentially large number of possible reallocations, making it computationally impractical as the number of agents and objects grows (e.g., De Keijzer et al. (2009), Aziz et al. (2019)). Adding to this challenge, no individual-good-based rule simultaneously satisfies Pareto efficiency and individual rationality under responsive preferences (Manjunath and Westkamp (2024)). In contrast, individual-good efficiency is more readily attainable and can be verified relatively easily. [^6] More precisely, individual-good efficiency can be verified in polynomial time, whereas verifying Pareto efficiency is co-NP complete (e.g., De Keijzer et al. (2009), Aziz et al. (2019)). This practical advantage makes individual-good efficiency a more realistic goal in settings with complex preferences and limited computational resources.
Truncation-proofness is an incentive compatibility requirement that prevents agents from benefiting by misrepresenting their preferences through truncation strategies. In our multi-object setting, a truncation strategy involves an agent dropping all objects owned by other agents and ranked below a certain cutoff object to the bottom of her rank-order list, possibly to avoid being assigned those objects. Truncation strategies are straightforward for agents to implement, as they only need to identify the cutoff point in their preference ordering. These strategies have been extensively studied in the matching literature due to their simplicity and theoretical appeal (Roth and Vande Vate (1991), Roth and Rothblum (1999), Ehlers (2008), Kojima (2013), Coles and Shorrer (2014)). Empirical evidence shows that agents often employ such strategies in real-world matching markets (Mongell and Roth (1991)) and object allocation problems (Guillen and Hakimov (2018)).
In addition to individual-good-based rules, we explore allocation rules that utilize a richer reporting language capable of expressing complementarities among objects. Specifically, we consider rules defined on the domain of conditionally lexicographic preferences. [^7] Preferences are conditionally lexicographic if, for any bundle $Y$ and any nonempty bundle $X$ disjoint from $Y$, there is an object $o\in X$ that is the “lexicographically best” addition to $Y$. Despite their added flexibility, these preferences can be described succinctly using “lexicographic preference trees.” Furthermore, this domain preserves the desirable equivalence between individual-good efficiency and Pareto efficiency found under lexicographic preferences.
The Augmented Top Trading Cycles (ATTC) rule, introduced by Fujita et al. (2018), extends the TTC rule to the conditionally lexicographic domain. We show that the ATTC rule satisfies generalized versions of our key properties and provide a characterization based on balancedness, Pareto efficiency, the worst endowment lower bound, and drop strategy-proofness (Theorem 6). Finally, we demonstrate that the conditionally lexicographic domain is maximal in the sense that Pareto efficiency and individual-good efficiency coincide only on this domain, highlighting the difficulty in extending our results to broader preference domains.
The remainder of this paper is organized as follows. The next subsection presents an overview of the related literature. In Section 2, we introduce the model details, including the lexicographic and responsive preference domains, key properties of allocation rules, and a formal description of the generalized TTC rule. Our results for the lexicographic domain are presented in Section 3. Section 4 extends our analysis to the more general responsive domain. Our findings for the housing market are discussed in Section 5. In Section 6, we introduce the domain of conditionally lexicographic preferences, discuss the extension of our properties to this domain, and we give a characterization of the ATTC rule. Section 7 concludes. Appendix A contains all proofs omitted from the main text, and we demonstrate the independence of our properties in Appendix B.1. Several other examples are provided in Appendix B.2.
1.1 Related literature
The seminal work of Shapley and Scarf (1974) introduces the housing market and the famous TTC rule, attributed to David Gale. They establish that the TTC rule always selects a core allocation. Roth (1982) demonstrates that the TTC rule is strategy-proof, and Ma (1994) provides a fundamental characterization, showing that it is the unique rule satisfying Pareto efficiency, individual rationality, and strategy-proofness.
Building upon this foundation, the early literature on multi-object reallocation consists mostly of negative results. Sönmez (1999) shows that no rule satisfies Pareto efficiency, individual rationality, and strategy-proofness on the unrestricted preference domain when agents may own multiple objects. Todo et al. (2014) shows that the impossibility holds even on the lexicographic domain, and Konishi et al. (2001) demonstrates that the core can be empty even if agents have additive preferences.
In response to these impossibility results, recent studies have obtained positive findings by restricting to more well-behaved variants of the multi-object reallocation problem. For instance, Pareto efficiency, individual rationality, and strategy-proofness are compatible when agents have “dichotomous” preferences (Andersson et al. (2021)) or “trichotomous” preferences (Manjunath and Westkamp (2021)). Similarly, the ATTC rule always selects an allocation within the core when agents have conditionally lexicographic preferences (Fujita et al. (2018)).
In a related study, Biró et al. (2022a) examine a model where each agent owns multiple copies of a homogeneous, agent-specific object. They characterize the capacity configurations under which Pareto efficiency, individual rationality, and strategy-proofness are compatible. Focusing on individual-good-based rules for responsive preferences, they characterize a variant of the TTC rule using subset drop strategy-proofness. We build upon their approach when extending our characterization from the lexicographic domain to the responsive domain.
Altuntaş et al. (2023) consider the general multi-object reallocation problem but focus on the lexicographic domain. They show that the TTC rule is drop strategy-proof, providing a characterization based on this property. Our Theorem 2 generalizes their result by showing that uniqueness holds under substantially weaker criteria. Unlike the proof in Altuntaş et al. (2023), which is constructive, we proceed by minimal counterexample, borrowing the notions of “size” from Sethuraman (2016) and “similarity” from Ekici (2024). The novelty of our approach lies in simultaneously exploiting both functions to select a profile satisfying a different minimality criterion.
Several papers illustrate the tightness of the incompatibility among Pareto efficiency, individual rationality, and strategy-proofness by characterizing rules that satisfy only two of the three objectives. When agents consume more than one object, combining Pareto efficiency and strategy-proofness with non-bossiness leads to sequential dictatorships (Pápai (2001), Ehlers and Klaus (2003), Hatfield (2009), Monte and Tumennasan (2015)). Other studies prioritize individual rationality and strategy-proofness at the expense of Pareto efficiency. For example, on the domain of responsive preferences, the Segmented Trading Cycles rules are characterized by strategy-proofness, non-bossiness, trade sovereignty, and strong individual rationality (Pápai (2003); see also Pápai (2007), Anno and Kurino (2016)).
In our paper, we resolve the incompatibility by weakening strategy-proofness to truncation-proofness or drop strategy-proofness. This approach is partly justified by several papers emphasizing the difficulty of manipulating the TTC rule (e.g., Fujita et al. (2018), Phan and Purcell (2022)). There are many other relaxed forms of strategy-proofness, such as rank monotonicity (Chen and Zhao (2021)), partial strategy-proofness (Mennle and Seuken (2021)), truncation-invariance (Chen et al. (2024)), weak truncation robustness (Hashimoto et al. (2014)), and convex strategy-proofness (Balbuzanov (2016)).
Our notion of truncation strategies has its roots in the literature on matching theory (Mongell and Roth (1991), Roth and Vande Vate (1991), Roth and Rothblum (1999), Ehlers (2008)), though the version we consider aligns more closely with the definitions provided by Kojima (2013) and Biró et al. (2022a, b) for models of multi-object (re)allocation.
As we move beyond individual-good-based rules and consider more complex reporting languages, (approximate) competitive equilibrium becomes an appropriate solution concept. In this approach, agents trade through a pseudo-market procedure, as first proposed by Hylland and Zeckhauser (1979) for the house allocation model. This method has developed significantly in recent years (e.g., Echenique et al. (2021, 2023), Nguyen et al. (2021), Kornbluth and Kushnir (2023), Nguyen and Vohra (2024)). Although this approach is effective at approximating desirable outcomes, it suffers from onerous reporting requirements and considerable computational complexity. Our study of the conditionally lexicographic domain provides a practical alternative that balances expressiveness with tractability.
2 Model
Let $N=\left\{1,2,\dots,n\right\}$ be a finite set of $n\geq 2$ agents. Let $O$ be a finite set of heterogeneous and indivisible objects such that $|O|\geq n$. A bundle is a subset of $O$. Let $2^{O}$ denote the set of bundles. We denote generic elements of $O$ by lowercase letters (e.g., $x,y,z$), and generic elements of $2^{O}$ by uppercase letters (e.g., $X,Y,Z$). To simplify notation, when there is no risk of confusion, we identify a singleton set $\left\{x\right\}$ with the element $x$ itself. For example, we write $X\cup x$ to denote $X\cup\{x\}$.
An allocation is a function $\mu:N\to 2^{O}$ such that (i) for all $i\in N$, $\mu(i)\neq\emptyset$, (ii) for all $i,j\in N$, $i\neq j$ implies $\mu\left(i\right)\cap\mu\left(j\right)=\emptyset$, and (iii) $\bigcup_{i\in N}\mu\left(i\right)=O$. Thus, an allocation $\mu$ can be represented as a profile $\left(\mu_{i}\right)_{i\in N}$ of nonempty, pairwise disjoint bundles satisfying $\bigcup_{i\in N}\mu_{i}=O$. For each $i\in N$, $\mu_{i}$ is called agent $i$’s assignment at $\mu$. Let $\mathcal{A}$ denote the set of allocations.
The initial allocation, also referred to as the endowment allocation, is denoted by $\omega=\left(\omega_{i}\right)_{i\in N}$. For each $i\in N$, $\omega_{i}$ is called agent $i$’s endowment. The initial owner of object $o$, denoted $\omega^{-1}\left(o\right)$, is the agent $i$ such that $o\in\omega_{i}$.
Each agent $i$ has a (strict) preference relation $P_{i}$ on the set of bundles. We assume that $P_{i}$ belongs to some specified subset $\mathcal{P}_{i}$ of all strict preference relations on $2^{O}$. In subsequent sections, we impose further structure on the sets $\mathcal{P}_{i}$. If agent $i$ prefers bundle $X$ to bundle $Y$, then we write $X\mathbin{P_{i}}Y$. Let $R_{i}$ denote the at least as good as relation associated with $P_{i}$, defined by $X\mathbin{R_{i}}Y$ if and only if ($X\mathbin{P_{i}}Y$ or $X=Y$). [^8] Formally, $R_{i}$ is a linear order (i.e., a complete, transitive, and antisymmetric binary relation) on $2^{O}$ , and $P_{i}$ is the strict (i.e., irreflexive and asymmetric) part of $R_{i}$. Given a nonempty bundle $X\in 2^{O}$, $\max_{P_{i}}\left(X\right)$ denotes the most-preferred object in $X$ at $P_{i}$, i.e., $\max_{P_{i}}\left(X\right)=x$ if $x\in X$ and $x\mathbin{R_{i}}y$ for all $y\in X$. Similarly, $\min_{P_{i}}\left(X\right)$ denotes the least-preferred object in $X$ at $P_{i}$, i.e., $\min_{P_{i}}\left(X\right)=x$ if $x\in X$ and $y\mathbin{R_{i}}x$ for all $y\in X$. A preference profile is an indexed family $P=\left(P_{i}\right)_{i\in N}$ of preference relations. The domain is the set $\mathcal{P}\coloneqq\prod_{i\in N}\mathcal{P}_{i}$, representing all possible preference profiles under consideration.
An object reallocation problem (or simply a problem) is a triple $(N,\omega,P)$. Since $(N,\omega)$ remains fixed throughout, we will identify a problem with its preference profile $P$. Thus, the domain $\mathcal{P}=\prod_{i\in N}\mathcal{P}_{i}$ of preference profiles represents the set of all problems.
A rule (on $\mathcal{P}$) is a function $\varphi:\mathcal{P}\to\mathcal{A}$ that associates with each preference profile $P$ an allocation $\varphi\left(P\right)$. For each $i\in N$, $\varphi_{i}\left(P\right)$ denotes agent $i$’s assignment at $\varphi\left(P\right)$.
Individual-good-based rules
In most of this paper, we focus on rules that can be implemented with a simple reporting language consisting of linear orders over individual objects. Such rules are often desired in practice because their relatively modest informational requirements streamline their implementation and reduce the burden on participants. A landmark example is the NRMP, which matches doctors to hospitals across the United States and has become a model for market-design interventions worldwide (e.g., Roth and Peranson (1999), Milgrom (2009, 2011)). [^9] In the NRMP, each hospital reports only a “rank-order list” over individual doctors together with the number of positions it would like to fill.
To formalize this concept, we need some notation. Given a preference relation $P_{i}$ on $2^{O}$, the induced marginal preference relation over individual objects is the strict linear order $\succ^{P_{i}}$ on $O$ such that, for all $x,y\in O$, $x\succ^{P_{i}}y$ if and only if $x\mathbin{P_{i}}y$ [^10] In other words, $\succ^{P_{i}}$ is the restriction of $P_{i}$ to singleton subsets of $O$. For all $x,y\in O$, $x\succeq^{P_{i}}y$ means that ($x\succ^{P_{i}}y$ or $x=y$). We shall often represent a marginal preference $\succ^{P_{i}}$ as an ordered list of objects; for example, $\succ^{P_{i}}:x_{1},x_{2},\dots,x_{|O|}$ means that $x_{1}\succ^{P_{i}}x_{2}\succ^{P_{i}}\cdots\succ^{P_{i}}x_{|O|}$, and $\succ^{P_{i}}:x_{1},x_{2},\dots,x_{k},\dots$ means that $x_{1}\succ^{P_{i}}x_{2}\succ^{P_{i}}\dots\succ^{P_{i}}x_{k}\succ^{P_{i}}o$ for all $o\in O\backslash\{x_{1},x_{2},\dots,x_{k}\}$.
For a preference profile $P=(P_{i})_{i\in N}$, the induced marginal preference profile over individual objects is the profile $\succ^{P}=(\succ^{P_{i}})_{i\in N}$. A rule is individual-good-based if it depends solely on the agents’ marginal preferences.
Definition 1.
A rule $\varphi$ is individual-good-based if
$\text{for all }P,P^{\prime}\in\mathcal{P},\quad\succ^{P}=\succ^{P^{\prime}}\implies\varphi(P)=\varphi(P^{\prime}).$ (1)
Individual-good-based rules are particularly well suited to environments in which the objects are substitutes, as agents’ preferences over bundles can be effectively summarized by their marginal preferences over individual objects. Moreover, even in settings where complementarities exist, focusing on individual-good-based rules can be practical and effective. As noted by Roth and Peranson (1999) and Roth (2002), the desirable theoretical properties of the NRMP, guaranteed for simple models without complementarities, tend to hold approximately even in more complex real-world markets.
2.1 Preference domains
In the following, we introduce two domains of preferences in which individual-good-based rules are especially effective: the lexicographic and responsive preference domains.
An agent has lexicographic preferences if, when evaluating distinct bundles $X$ and $Y$, she prefers the bundle containing the most-preferred object in $X\cup Y$; if the most-preferred object in $X\cup Y$ is common to $X$ and $Y$, then she prefers the bundle containing the second-most-preferred object in $X\cup Y$, and so on. This decision-making process reflects heuristics that are typical in human behavior. [^11] The “Take The Best” heuristic, where individuals make choices by considering the most important attribute first and proceeding sequentially, performs surprisingly well in real-world inferential tasks (Gigerenzer and Goldstein (1996); Gigerenzer and Todd (1999)). Formally, a preference relation $P_{i}$ on $2^{O}$ is lexicographic if, for any two distinct bundles $X$ and $Y$,
$X\mathbin{P_{i}}Y\iff\max_{P_{i}}\left(X\triangle Y\right)\in X,$ (2)
where $X\triangle Y=(X\backslash Y)\cup(Y\backslash X)$ is the symmetric difference of $X$ and $Y$. For each $i\in N$, let $\mathcal{L}_{i}$ be the set of lexicographic preferences on $2^{O}$. Then $\mathcal{L}\coloneqq\prod_{i\in N}\mathcal{L}_{i}$ is called the lexicographic domain.
Although it is rather restrictive, the lexicographic domain is a natural starting point in our analysis because any rule defined on it is automatically individual-good-based. This is because each lexicographic preference $P_{i}\in\mathcal{L}_{i}$ is uniquely determined by its marginal preference $\succ^{P_{i}}$ over individual objects; that is, for all $P_{i},P^{\prime}_{i}\in\mathcal{L}_{i}$, $\succ^{P_{i}}=\succ^{P^{\prime}_{i}}$ implies $P_{i}=P^{\prime}_{i}$. Consequently, we identify each $P_{i}\in\mathcal{L}_{i}$ with its associated marginal preference $\succ^{P_{i}}$ and write $P_{i}:x_{1},x_{2},\dots,x_{|O|}$ if $x_{1}\mathbin{P_{i}}x_{2}\mathbin{P_{i}}\cdots\mathbin{P_{i}}x_{|O|}$.
We are primarily interested in the domain of responsive preferences, a more general domain first studied by Roth (1985) for many-to-one matching models. An agent has responsive preferences if, for any two bundles that differ in one object, she prefers the bundle containing the more-preferred object. Formally, a preference relation $P_{i}$ on $2^{O}$ is responsive if, for any bundle $X$ and any objects $y,z\in O\setminus X$, $y\mathbin{P_{i}}z$ if and only if $(X\cup y)\mathbin{P_{i}}(X\cup z)$. Intuitively, responsiveness rules out complementarities, as the relative ranking between any two objects is independent of the other objects they are obtained with. For each $i\in N$, let $\mathcal{R}_{i}$ denote the set of responsive preferences on $2^{O}$. Then $\mathcal{R}\coloneqq\prod_{i\in N}\mathcal{R}_{i}$ is called the responsive domain.
Finally, it is convenient to define the domain of monotonic preferences. A preference relation $P_{i}$ on $2^{O}$ is monotonic if, for any bundles $X$ and $Y$, $X\mathbin{R_{i}}Y$ whenever $X\supseteq Y$. For each $i\in N$, let $\mathcal{M}_{i}$ be the set of monotonic preferences on $2^{O}$. Then $\mathcal{M}\coloneqq\prod_{i\in N}\mathcal{M}_{i}$ is called the monotonic domain.
Note that every lexicographic preference relation is both responsive and monotonic, but the converse is false whenever $|O|\geq 3$. Moreover, while the sets of responsive and monotonic preferences overlap, neither set is entirely contained within the other. Specifically, for any $i\in N$, we have $\mathcal{L}_{i}\subseteq\mathcal{R}_{i}\cap\mathcal{M}_{i}$ (with strict inclusion for $|O|\geq 3$), $\mathcal{R}_{i}\nsubseteq\mathcal{M}_{i}$, and $\mathcal{M}_{i}\nsubseteq\mathcal{R}_{i}$.
2.2 Properties of allocation rules
This section introduces several desirable properties of allocation rules. All properties are defined for any arbitrary domain $\mathcal{P}$ in which each agent has strict preferences.
Our first requirement is that exchange be “balanced” in the sense that each agent ends up with the same number of objects as initially endowed. Formally, an allocation $\mu$ is balanced if, for each agent $i\in N$, the cardinality of her assignment equals that of her endowment: $|\mu_{i}|=|\omega_{i}|$.
Definition 2.
A rule $\varphi$ is balanced if, for each $P\in\mathcal{P}$, $\varphi(P)$ is balanced.
Balancedness is often a key consideration—and sometimes a strict constraint—in practical reallocation problems. For example, in shift reallocation, balancedness may be desired to prevent overwork or underemployment. It might also be imposed for training purposes or because employment contracts specify a certain number of shifts per week (Manjunath and Westkamp, 2021). Similarly, in student and tuition exchange programs, balancedness is desired (at least in the long run) to maintain reciprocal relationships and prevent education costs from increasing at popular schools (Andersson et al. (2021); Biró et al. (2022a); Dur and Ünver (2019)). [^12] Indeed, Dur and Ünver (2019) document several cases in which such long-run imbalances led to the failure of an exchange program. In the absence of strict constraints, balancedness captures a limited notion of equity: it ensures the gains from trade are shared fairly among the agents, at least regarding the number of objects exchanged.
2.2.1 Efficiency
An allocation $\overline{\mu}$ Pareto-dominates another allocation $\mu$ at a preference profile $P$ if (i) for all $i\in N$, $\overline{\mu}_{i}\mathbin{R_{i}}\mu_{i}$, and (ii) for some $i\in N$, $\overline{\mu}_{i}\mathbin{P_{i}}\mu_{i}$. An allocation $\mu$ is Pareto efficient at $P$ if it is not Pareto-dominated at $P$ by any other allocation. The strongest efficiency property that we consider is the following.
Definition 3.
A rule $\varphi$ is Pareto efficient if, for each $P\in\mathcal{P}$, $\varphi(P)$ is Pareto efficient at $P$.
If an allocation is not Pareto efficient, then a group of agents could, in principle, destabilize it by carrying out a Pareto-improving exchange. However, such exchanges are generally complex and difficult to coordinate, potentially involving intricate trades of multiple objects among many agents. Indeed, the problem of verifying whether a given allocation is Pareto efficient is computationally intractable: even when agents have “additive” preferences (a subclass of responsive preferences), determining Pareto efficiency is coNP-complete (e.g., De Keijzer et al. (2009), Aziz et al. (2019)).
Given these challenges, we consider a weaker notion of efficiency that is more readily attainable. Individual-good efficiency rules out destabilizing Pareto-improving exchanges that can be easily coordinated due to their relatively simple structure. Toward a formalization, we say that an allocation $\mu$ admits a Pareto-improving single-object exchange at a preference profile $P$ if there is a cycle $C=\left(o_{1},i_{1},o_{2},i_{2},o_{3},\dots,i_{k},o_{k+1}=o_{1}\right)$ of objects and agents such that, for all $\ell\in\left\{1,\dots,k\right\}$,
$i_{\ell}\in N,\quad o_{\ell}\in\mu_{i_{\ell}},\quad\text{and}\quad(\mu_{i_{\ell}}\cup o_{i_{\ell+1}})\backslash o_{i_{\ell}}\mathbin{P_{i_{\ell}}}\mu_{i_{\ell}}.$ (3)
An allocation $\mu$ is called individual-good efficient (ig-efficient) at $P$ if it does not admit a Pareto-improving single-object exchange at $P$ [^13] The terminology “ig-efficiency” is borrowed from Biró et al. (2022a). Similar properties are studied in Aziz et al. (2019), Caspari (2020), and Coreno and Balbuzanov (2022).
Definition 4.
A rule $\varphi$ is individual-good efficient (ig-efficient) if, for each $P\in\mathcal{P}$, $\varphi(P)$ is ig-efficient at $P$.
Clearly, Pareto efficiency implies ig-efficiency on an arbitrary domain of strict preferences. Although the two properties are equivalent under lexicographic preferences (Aziz et al., 2019), [^14] We show that this equivalence extends to “conditionally lexicograhic” preferences (Proposition 5). ig-efficiency is substantially weaker than Pareto efficiency under more general preference domains. However, a practical advantage of ig-efficiency is that it can be verified in polynomial time (e.g., Cechlárová et al. (2014), Aziz et al. (2019)).
2.2.2 Participation guarantees
The following is a standard participation guarantee which ensures that no agent is harmed by the reallocation. An allocation $\mu$ is called individually rational at a preference profile $P$ if, for each $i\in N$, $\mu_{i}\mathbin{R_{i}}\omega_{i}$.
Definition 5.
A rule $\varphi$ is individually rational if, for each $P\in\mathcal{P}$ and each $i\in N$, $\varphi_{i}(P)\mathbin{R_{i}}\omega_{i}$.
We now introduce another participation guarantee that depends exclusively on the agents’ rankings over individual objects. Specifically, it ensures that no agent receives an object that is worse than her least-preferred object in her own endowment. An allocation $\mu$ satisfies the worst endowment lower bound at a preference profile $P$ if, for each agent $i\in N$, and each of her assigned objects $o\in\mu_{i}$, $o\mathbin{R_{i}}\min_{P_{i}}\left(\omega_{i}\right)$.
Definition 6.
A rule $\varphi$ satisfies the worst endowment lower bound if, for each $P\in\mathcal{P}$, $\varphi(P)$ satisfies the worst endowment lower bound at $P$.
Individual rationality and the worst endowment lower bound coincide on the class of single-unit reallocation problems. However, for general multi-object problems, the two properties are independent. Crucially, individual rationality allows for an agent to be assigned any unfavorable object as long as it is part of a desirable bundle, whereas the worst endowment lower bound ensures that agents are not assigned highly unfavorable objects. Thus, a rule satisfying the worst endowment lower bound effectively allows agents to veto certain objects owned by the other agents. This veto right can be viewed as a minimal participation guarantee and is particularly meaningful in certain applications. In shift reallocation, for example, a worker may have prior commitments that prevent her from fulfilling certain shifts. By ranking these shifts below every shift in her endowment, she effectively vetoes them.
Under certain mild conditions, individual rationality is a stronger requirement than the worst endowment lower bound. For example, if we restrict attention to the class of individual-good-based and balanced rules defined on the responsive domain, the main focus of the present paper, then individual rationality implies the worst endowment lower bound. [^15] Alternatively, we could consider a natural domain under which any bundle that violates the worst endowment lower bound is “unacceptable.” Formally, for each $i\in N$ and each $P_{i}\in\mathcal{P}_{i}$, any bundle that intersects $\{o\in O\mid o\mathbin{R_{i}}\min_{P_{i}}(\omega_{i})\}$ is worse than any bundle that does not. Clearly, any rule on $\mathcal{P}^{*}$ that satisfies individual rationality also satisfies the worst endowment lower bound. Furthermore, if we assume preferences $P_{i}$ to be lexicographic (or responsive) when restricted to $\{o\in O\mid o\mathbin{R_{i}}\min_{P_{i}}(\omega_{i})\}$, then our subsequent characterization results would remain valid.
Lemma 1.
On the responsive domain, if an individual-good-based rule is balanced and individually rational, then it satisfies the worst endowment lower bound.
2.2.3 Incentive properties
Let $P=(P_{i})_{i\in N}$ be a preference profile, and let $P^{\prime}_{i}$ be a preference relation for agent $i$. We use the standard notation $\left(P^{\prime}_{i},P_{-i}\right)$ to denote the preference profile in which agent $i$’s preference relation is $P^{\prime}_{i}$ and, for each agent $j\in N\backslash\{i\}$, agent $j$’s preference relation remains $P_{j}$. Given a rule $\varphi$, we say that agent $i$ can manipulate $\varphi$ at $P$ by misreporting $P^{\prime}_{i}$ if $\varphi_{i}(P^{\prime}_{i},P_{-i})\mathbin{P_{i}}\varphi_{i}(P)$. A rule is strategy-proof if no agent can manipulate it by misreporting any preference relation.
Definition 7 (Strategy-proofness).
A rule $\varphi$ is strategy-proof if, for each $P\in\mathcal{P}$, each $i\in N$, and each $P^{\prime}_{i}\in\mathcal{P}_{i}$, $\varphi_{i}(P)\mathbin{R_{i}}\varphi_{i}(P^{\prime}_{i},P_{-i})$.
Even on the relatively narrow lexicographic domain, no allocation rule simultaneously satisfies strategy-proofness, individual-good efficiency, and individual rationality (Todo et al., 2014). Given this incompatibility, we explore relaxations of strategy-proofness by restricting the set of manipulation strategies that agents might employ. In particular, we focus on “truncation strategies” and “drop strategies.” These strategies involve agents misrepresenting their marginal preferences in straightforward ways, making them intuitively appealing and easy to implement.
Truncation strategies and drop strategies are both special kinds of “subset drop strategies” (Biró et al. (2022a, b), Altuntaş et al. (2023)). Loosely speaking, a subset drop strategy is a manipulation whereby an agent drops a subset of the other agents’ endowments to the bottom of her marginal preference list, possibly to avoid being assigned those objects. Truncation strategies are obtained by dropping a “tail subset” of objects, i.e., a subset consisting of the agent’s least-preferred objects, whereas drop strategies are obtained by dropping a singleton subset. Because these strategies involve agents misrepresenting their marginal preferences, they are particularly meaningful when dealing with individual-good-based rules.
Truncation strategies have a well-established history in matching literature, particularly in models where agents are matched to single objects (e.g., Mongell and Roth (1991), Roth and Vande Vate (1991), Roth and Rothblum (1999), Ehlers (2008)). In our multi-object reallocation setting, the notion of truncation strategies aligns more closely with the definitions provided by Kojima (2013) and Biró et al. (2022a, b). Empirical evidence suggests that agents employ truncation strategies in practice (see Mongell and Roth (1991) and Guillen and Hakimov (2018)). Similarly, drop strategies have been studied by Altuntaş et al. (2023) in the context of multi-object reallocation under lexicographic preferences.
Given an agent $i$ and a preference relation $P_{i}\in\mathcal{P}_{i}$, we say that $P^{\prime}_{i}\in\mathcal{P}_{i}$ is a subset drop strategy for $P_{i}$ if there exists $X\subseteq O\backslash\omega_{i}$ such that:
  1. 1.
    for all $x\in X$ and $y\in O\backslash X$, $y\mathbin{P^{\prime}_{i}}x$; and
  2. 2.
    for all $x,y\in X$, $x\mathbin{P^{\prime}_{i}}y$ if and only if $x\mathbin{P_{i}}y$; and
  3. 3.
    for all $x,y\in O\backslash X$, $x\mathbin{P^{\prime}_{i}}y$ if and only if $x\mathbin{P_{i}}y$.
In this case, we say that $P^{\prime}_{i}$ is obtained from $P_{i}$ by dropping $X$ (and that $\succ^{P^{\prime}_{i}}$ is obtained from $\succ^{P_{i}}$ by dropping $X$). Furthermore, $P^{\prime}_{i}$ is a drop strategy for $P_{i}$ if it is obtained by dropping a singleton subset, i.e., $|X|=1$. Finally, $P^{\prime}_{i}$ is a truncation strategy for $P_{i}$ if either $X=O\backslash\omega_{i}$ or $X=\{o\in O\backslash\omega_{i}\mid x\mathbin{P_{i}}o\}$ for some $x\in O$. In the latter case, i.e., when $X=\{o\in O\backslash\omega_{i}\mid x\mathbin{P_{i}}o\}$, $P^{\prime}_{i}$ is called a truncation of $P_{i}$ at $x$ [^16] Equivalently, $P^{\prime}_{i}$ is a truncation of $P_{i}$ at $x$ if (i) $\succ^{P^{\prime}_{i}}$ agrees with $\succ^{P_{i}}$ on $\omega_{i}$; (ii) $\succ^{P^{\prime}_{i}}$ agrees with $\succ^{P_{i}}$ on $O\backslash\omega_{i}$; and (iii) for each $y\in O\backslash\omega_{i}$ with $x\mathbin{P_{i}}y$, $\min_{P^{\prime}_{i}}(\omega_{i})\mathbin{P^{\prime}_{i}}y$. Let $\mathcal{S}_{i}(P_{i})$ $(\subseteq\mathcal{P}_{i})$ denote the set of all subset drop strategies for $P_{i}$. Similarly, $\mathcal{D}_{i}(P_{i})$ and $\mathcal{T}_{i}(P_{i})$ denote, respectively, the sets of drop strategies and truncation strategies for $P_{i}$.
A rule is truncation-proof if it cannot be manipulated through truncation strategies. Drop strategy-proofness and subset drop strategy-proofness are defined analogously.
Definition 8.
A rule $\varphi$ is
  • truncation-proof if, for each $P\in\mathcal{P}$, each $i\in N$, and each $P^{\prime}_{i}\in\mathcal{T}_{i}(P_{i})$, $\varphi_{i}(P)\mathbin{R_{i}}\varphi_{i}(P^{\prime}_{i},P_{-i})$.
  • drop strategy-proof if, for each $P\in\mathcal{P}$, each $i\in N$, and each $P^{\prime}_{i}\in\mathcal{D}_{i}(P_{i})$, $\varphi_{i}(P)\mathbin{R_{i}}\varphi_{i}(P^{\prime}_{i},P_{-i})$.
  • subset drop strategy-proof if, for each $P\in\mathcal{P}$, each $i\in N$, and each $P^{\prime}_{i}\in\mathcal{S}_{i}(P_{i})$, $\varphi_{i}(P)\mathbin{R_{i}}\varphi_{i}(P^{\prime}_{i},P_{-i})$.
Subset drop strategy-proofness entails both truncation-proofness and drop strategy-proofness, as it defends against a larger set of manipulation strategies. However, truncation-proofness and drop strategy-proofness are independent; a rule can satisfy one without satisfying the other.
Before illustrating these definitions with examples (Example 1), we offer two remarks to clarify certain technical points.
Remark 1.
For any preference $P_{i}\in\mathcal{P}_{i}$ and any subset $X\subseteq O\backslash\omega_{i}$, there is a unique marginal preference, say $\succ^{P^{\prime}_{i}}$, obtained from $\succ^{P_{i}}$ by dropping $X$ to the bottom of the preference list. However, there may be multiple preference relations $P^{\prime\prime}_{i}\in\mathcal{P}_{i}$ that share this marginal preference $\succ^{P^{\prime}_{i}}$, and each such $P^{\prime\prime}_{i}$ is a subset drop strategy obtained from $P_{i}$ by dropping $X$. Despite this multiplicity, any individual-good-based rule will choose the same allocation for all such $P^{\prime\prime}_{i}$, so this technical detail does not play a substantive role in our analysis.
Remark 2.
By employing some subset drop strategy for $P_{i}$ (or successively employing drop strategies), an agent can push any object in $\{o\in O\backslash\omega_{i}\mid o\mathbin{P_{i}}\min_{P_{i}}(\omega_{i})\}$ to the top of her marginal preference list. This is not possible with truncation strategies, as every truncation strategy $P^{*}_{i}$ for $P_{i}$ must agree with $P_{i}$ on $O\backslash\omega_{i}$ (see footnote 16).
Example 1.
Suppose $\omega_{i}=\{x,y\}$, and let $P_{i}\in\mathcal{P}_{i}$ be such that $\succ^{P_{i}}:a,b,\underline{x},c,d,\underline{y},e$ (agent $i$’s endowment is underlined for emphasis). Then:
  • any $P^{1}_{i}\in\mathcal{P}_{i}$ with $\succ^{P^{1}_{i}}:b,\underline{x},c,d,\underline{y},e,a$ is a drop strategy obtained from $P_{i}$ by dropping object $a$. Note that any such $P^{1}_{i}$ is not a truncation strategy for $P_{i}$.
  • any $P^{2}_{i}\in\mathcal{P}_{i}$ with $\succ^{P^{2}_{i}}:a,b,\underline{x},c,\underline{y},d,e$ is a truncation strategy obtained from $P_{i}$ from $P_{i}$ by dropping $\{d,e\}$ (or successively dropping $d$ then $e$).
  • any $P^{3}_{i}\in\mathcal{P}_{i}$ with $\succ^{P^{3}_{i}}:a,\underline{x},\underline{y},b,c,d,e,$ is a truncation strategy obtained from $P_{i}$ by dropping $\{b,c,d,e\}$ (or successively dropping $b,c,d$, then $e$). $\diamond$
2.3 Top Trading Cycles
A cycle is a circular sequence
$C=\left(o_{i_{1}},i_{1},o_{i_{2}},i_{2},o_{i_{3}},\dots,i_{k-1},o_{i_{k}},i_{k},o_{i_{k+1}}=o_{i_{1}}\right)$
consisting of $k$ $(\geq 1)$ distinct objects and $k$ distinct agents such that (i) each object on the cycle precedes (or “points to”) its owner, and (ii) each agent on the cycle points to an object. The sets of agents and objects on $C$ are denoted by $N\left(C\right)=\left\{i_{1},i_{2},\dots,i_{k}\right\}$ and $O(C)=\left\{o_{i_{1}},o_{i_{2}},\dots,o_{i_{k}}\right\}$, respectively. An allocation $\mu$ is said to execute the cycle $C$ if it assigns to each agent in $N\left(C\right)$ the object she points to within $C$; that is, for each $i_{\ell}\in N\left(C\right)$, $o_{i_{\ell+1}}\in\mu_{i_{\ell}}$.
We study an extension of Gale’s Top Trading Cycles algorithm (Shapley and Scarf, 1974) from single-object to multi-object reallocation problems. At each step of this procedure, every agent points to her most-preferred unassigned object, and every unassigned object points to its owner. There exists at least one cycle, and each agent involved in a cycle is assigned the object to which she points. All objects involved in a cycle are then removed. If unassigned objects remain, then the procedure continues to the next step; otherwise, it terminates with the corresponding allocation.
For our proofs, we consider a modified version of this procedure that executes only one cycle at each step. Specifically, if multiple cycles arise, then this modified procedure executes only the cycle containing the minimum agent—the agent with the smallest label among those involved in cycles. It is well known that this modified procedure yields the same allocation as the standard procedure, which executes all prevailing cycles at each step.
We now formalize the (generalized) Top Trading Cycles (TTC) rule, which we denote by $\varphi^{\text{TTC}}$. Given a preference profile $P\in\mathcal{P}$, the allocation $\varphi^{\text{TTC}}(P)$ is determined by running the (generalized) TTC algorithm at $P$. We denote this specific instance as $\text{TTC}(P)$. The algorithm $\text{TTC}(P)$ is defined as follows.
Algorithm:
$\text{TTC}\left(P\right)$
Input: A preference profile $P\in\mathcal{P}$.
Output: An allocation $\varphi^{\text{TTC}}\left(P\right)$.
Initialization:
Set $\mu^{0}\coloneqq\left(\emptyset\right)_{i\in N}$ and $O^{1}\coloneqq O$.
Step $t\geq 1$:
  1. 1.
    (Graph construction) Construct a bipartite directed graph with independent vertex sets $N$ and $O^{t}$, and edge sets defined as follows:
    1. (a)
      For each agent $i\in N$, there is a directed edge from $i$ to $\max_{P_{i}}\left(O^{t}\right)$.
    2. (b)
      For each object $o\in O^{t}$, there is a directed edge from $o$ to its owner, $\omega^{-1}\left(o\right)$.
  2. 2.
    (Cycle selection) Because there are finitely many vertices, each with an outgoing edge, there is at least one cycle.
    1. (a)
      Let $\mathcal{C}_{t}(P)$ denote the set of cycles that arise at Step $t$.
    2. (b)
      Let $C_{t}(P)$ be the cycle in $\mathcal{C}_{t}(P)$ containing the minimum agent:
      i.e., $\min N(C_{t}(P))\leq\min N(C)$ for all $C\in\mathcal{C}_{t}(P)$.
  3. 3.
    (Assignment) Assign to each agent $i\in N(C_{t}(P))$ the object $\max_{P_{i}}(O^{t})$. That is, let $\mu^{t}=\left(\mu_{i}^{t}\right)_{i\in N}$ be such that
    1. (a)
      for all $i\in N(C_{t}(P))$, $\mu_{i}^{t}=\mu_{i}^{t-1}\cup\left\{\max_{P_{i}}\left(O^{t}\right)\right\}$, and
    2. (b)
      for all $i\in N\backslash N(C_{t}(P))$, $\mu_{i}^{t}=\mu_{i}^{t-1}$.
  4. 4.
    (Removal) Let $O^{t+1}\coloneqq O^{t}\backslash O(C_{t}(P))$ be the set of objects remaining at Step $t+1$.
    1. (a)
      If $O^{t+1}\neq\emptyset$, then proceed to Step $t+1$.
    2. (b)
      If $O^{t+1}=\emptyset$, then proceed to Termination.
Termination:
Because $O$ is finite and $\left|O^{1}\right|>\left|O^{2}\right|>\cdots>\left|O^{t}\right|$, the algorithm terminates at some step $T$. Return the allocation $\varphi^{\text{TTC}}\left(P\right)\coloneqq\mu^{T}$.
The following example illustrates the TTC algorithm.
Example 2.
Suppose $N=\{1,2,3\}$, $O=\{a,b,c,d\}$, and $\omega=(\{a,b\},\{c\},\{d\})$. Consider a preference profile $P=(P_{1},P_{2},P_{3})$, where $\succ^{P_{1}}:c,a,d,b$, $\succ^{P_{2}}:a,b,c,d$, and $\succ^{P_{3}}:a,c,b,d$. The algorithm $\text{TTC}(P)$ works as follows.
  • Step 1: Each agent points to her most-preferred object, and each object points to its owner. There is a cycle $C_{1}(P)=(1,c,2,a,1)$. We set $\mu^{1}=(\{c\},\{a\},\emptyset)$ and $O^{2}=\{b,d\}$. Since $O^{2}\neq\emptyset$, we proceed to Step 2.
  • Step 2: Each agent points to her most-preferred object among $O^{2}=\{b,d\}$, and each object in $O^{2}$ points to its owner. There is a cycle $C_{2}(P)=(1,d,3,b,1)$. We set $\mu^{2}=(\{c,d\},\{a\},\{b\})$ and $O^{3}=\emptyset$. Since $O^{3}=\emptyset$, we stop and return the allocation $\varphi^{\text{TTC}}(P)=(\{c,d\},\{a\},\{b\})$. $\diamond$
Several key properties of the TTC rule hold universally across any domain of strict preferences. The following proposition outlines these properties.
Proposition 1.
For any domain $\mathcal{P}$, the TTC rule is individual-good-based, balanced, and satisfies the worst endowment lower bound.
Proof.
At each step of the TTC algorithm, each agent points to her most-preferred unassigned object. Since this choice depends only on her marginal preferences over individual objects, the TTC rule is individual-good-based. Balancedness and the worst endowment lower bound follow from the fact that whenever an agent relinquishes an object from her endowment, she receives an object in return that is weakly preferred according to her marginal preferences. ∎
Whether the TTC rule satisfies other properties—such as Pareto efficiency, truncation-proofness, or drop strategy-proofness—depends on the specific domain of preferences being considered.
3 Results for Lexicographic Preferences
Our analysis begins on the domain of lexicographic preferences, a natural starting point because of its relative tractability. As mentioned previously, the lexicographic domain has the following desirable features: (i) any rule defined on it is individual-good-based, and (ii) Pareto efficiency is equivalent to individual-good efficiency, which can be verified in polynomial time (Cechlárová et al. (2014), Aziz et al. (2019)).
Within the lexicographic domain, the TTC rule exhibits several desirable properties. Fujita et al. (2018) demonstrate that it is core selecting, which entails both Pareto efficiency and individual rationality. Furthermore, Altuntaş et al. (2023) establish that the TTC rule is subset drop strategy-proof, implying that it is also truncation-proof and drop strategy-proof. These key properties are formally stated in the following proposition.
Proposition 2.
On the lexicographic domain, the TTC rule satisfies Pareto efficiency, individual rationality, and subset drop strategy-proofness.
We now present our main characterization of the TTC rule for the lexicographic domain. The following theorem states that the TTC rule is uniquely determined by the combination of balancedness, ig-efficiency, the worst endowment lower bound, and truncation-proofness.
Theorem 1.
On the lexicographic domain, a rule satisfies
  • balancedness,
  • ig-efficiency (or Pareto efficiency),
  • the worst endowment lower bound, and
  • truncation-proofness
if and only if it is the TTC rule.
The proof proceeds by minimal counterexample and exploits both a “similarity” function (Ekici, 2024) and a “size” function (Sethuraman, 2016). Our main technical innovation lies in combining both functions to establish the result. Suppose $\varphi$ is a rule that satisfies the properties but differs from $\varphi^{\text{TTC}}$. Call a preference profile $P\in\mathcal{P}$ a conflict profile if $\varphi(P)\neq\varphi^{\text{TTC}}(P)$. For each conflict profile $P$, let $\rho(P)\coloneqq t$ denote the earliest step $t$ of $\text{TTC}(P)$ at which $\varphi(P)$ does not execute $C_{t}(P)$, the cycle executed at step $t$ of $\text{TTC}(P)$. In this way, we define the “similarity” function $\rho$ from conflict profiles to the natural numbers (Ekici, 2024), measuring the earliest point of divergence between $\varphi$ and $\varphi^{\text{TTC}}$. Among all conflict profiles minimizing $\rho$, we select a profile $P$ that further minimizes the “size” function $s(P)\coloneqq\sum_{i\in N}|\{o\in O\mid o\mathbin{R_{i}}\min_{P_{i}}(\omega_{i})\}|$ (Sethuraman, 2016). We then demonstrate that, given our choice of $P$, the allocation $\varphi(P)$ cannot be ig-efficient at $P$, a contradiction.
Theorem 2 allows us to derive an alternative characterization of the TTC rule as a corollary. To see how, we first state a useful lemma. [^17] Example 7 in Appendix B.2 demonstrates that truncation-proofness and the worst endowment lower bound do not jointly imply drop strategy-proofness. Example 8 shows that truncation-proofness is not implied by drop strategy-proofness alone.
Lemma 2.
On the lexicographic domain, if a rule satisfies drop strategy-proofness and the worst endowment lower bound, then it is subset drop strategy-proof.
Suppose $\varphi$ is a rule satisfying balancedness, ig-efficiency, the worst endowment lower bound, and drop strategy-proofness. Such a rule exists by Propositions 1 and 2. Lemma 2 implies that $\varphi$ is subset drop strategy-proof, hence truncation-proof. Therefore, $\varphi$ satisfies the properties in Theorem 1, which means it must equal the TTC rule. This result is stated in the following theorem, which strengthens the characterization provided by Altuntaş et al. (2023, Theorem 1).
Theorem 2.
On the lexicographic domain, a rule satisfies
  • balancedness,
  • ig-efficiency (or Pareto efficiency),
  • the worst endowment lower bound, and
  • drop strategy-proofness
if and only if it is the TTC rule.
We cannot substitute the worst endowment lower bound with individual rationality in Theorems 1 and 2. On the lexicographic domain, there exist rules other than the TTC rule that satisfy balancedness, Pareto efficiency, individual rationality, truncation-proofness, and drop strategy-proofness. The following example illustrates this point.
Example 3.
Suppose that $N=\{1,2,3\}$, $O=\{a,b,c,d\}$, and $\omega=(\{a,b\},\{c\},\{d\})$. Let $\varphi$ be the rule defined as follows. For all $P\in\cal{L}$,
$\varphi(P)=\begin{cases}\varphi^{(1231)}(P),&\text{if }\max_{P_{1}}(O)=c\text{and }\max_{P_{2}}(O)\in\{a,b\}\\ \varphi^{(1321)}(P),&\text{if }\max_{P_{1}}(O)=d\text{ and }\max_{P_{3}}(O)\in\{a,b\}\\ \varphi^{\text{TTC}}(P),&\text{otherwise,}\end{cases}$
where $\varphi^{(ijk\ell)}$ is the sequential priority rule which, at each step, assigns to an agent her most-preferred unassigned object, proceeding in the order $(i,j,k,\ell)$.
One can show that $\varphi$ satisfies balancedness, ig-efficiency, individual rationality, truncation-proofness, and drop strategy-proofness. To see that $\varphi$ violates the worst endowment lower bound, let $P\in\mathcal{L}$ satisfy $P_{1}:c,b,a,d$ and $P_{2}=P_{3}:a,b,c,d$. Then $\varphi(P)=(\{c,d\},\{a\},\{b\})$, which violates the worst endowment lower bound because object $d$ is assigned to agent $1$ although $\min_{P_{1}}(\omega_{1})=a\mathbin{P_{1}}d$. $\diamond$
The issue illustrated in Example 3 arises from a crucial difference between the two participation guarantees. While individual rationality allows an agent to be assigned any object as part of a desirable bundle, the worst endowment lower bound restricts which objects can be included in an agent’s bundle. Interestingly, this issue does not occur on the responsive domain, where individual rationality suffices for the characterization (Theorem 4). [^18] Similarly, this issue does not arise on the domain described in footnote 15, as individual rationality implies the worst endowment lower bound on that domain.
4 Results for Responsive Preferences
In this section, we extend our analysis to the domain of responsive preferences, which presents several challenges compared to the lexicographic domain.
One major hurdle is that achieving Pareto efficiency becomes more difficult, as it generally requires coordinating intricate exchanges involving multiple objects among many agents. As mentioned previously, the problem of verifying whether an allocation is Pareto efficient is coNP-complete (e.g., De Keijzer et al. (2009), Aziz et al. (2019)), making it computationally challenging. Adding to this complexity, Manjunath and Westkamp (2024) show that no individual-good-based rule satisfies both Pareto efficiency and individual rationality. Since the TTC rule is an individual-good-based rule that satisfies individual rationality, it follows that the TTC rule cannot be Pareto efficient in this setting.
To illustrate this incompatibility, consider the following example.
Example 4 (Manjunath and Westkamp (2024)).
Suppose that $N=\{1,2\}$, $O=\{a,b,c,d\}$, and $\omega=(\{a,d\},\{b,c\})$. Toward contradiction, let $\varphi$ be an individual-good-based rule on $\mathcal{R}$ satisfying individual rationality and Pareto efficiency.
Let $P=(P_{1},P_{2})$ be the lexicographic preference profile where both agents rank individual objects in the order $\succ^{P_{1}}=\succ^{P_{2}}:a,b,c,d$. Since $\varphi$ is Pareto efficient and individually rational, it must assign $\varphi(P)=(\{a,d\},\{b,c\})$. Now let $P^{\prime}=(P^{\prime}_{1},P^{\prime}_{2})$ be a responsive preference profile with the same marginal preferences such that each agent prefers the other’s endowment; that is,
$\succ^{P^{\prime}_{1}}=\succ^{P^{\prime}_{2}}:a,b,c,d,\quad\{b,c\}\mathbin{P^{\prime}_{1}}\{a,d\},\quad\text{and}\quad\{a,d\}\mathbin{P^{\prime}_{2}}\{b,c\}.$
Because $\succ^{P^{\prime}}=\succ^{P}$ and $\varphi$ is individual-good-based, we have $\varphi(P^{\prime})=\varphi(P)=(\{a,d\},\{b,c\})$. However, $\varphi(P^{\prime})$ is Pareto-dominated by ($(\{b,c\},\{a,d\})$ at $P^{\prime}$, violating Pareto efficiency.
In this example, the TTC rule assigns the allocation ($(\{a,d\},\{b,c\})$ at both $P$ and $P^{\prime}$, which directly shows that it is not Pareto efficient. $\diamond$
Despite this limitation, the TTC rule remains ig-efficient on the responsive domain. To see this, suppose $\varphi^{\text{TTC}}(P)$ were not ig-efficient at some responsive preference profile $P$. Consider the corresponding lexicographic preference profile $P^{\prime}$ with $\succ^{P^{\prime}}=\succ^{P}$. Since the TTC rule is individual-good-based, we have $\varphi^{\text{TTC}}(P^{\prime})=\varphi^{\text{TTC}}(P)$. However, this would imply that $\varphi^{\text{TTC}}(P^{\prime})$ is not Pareto efficient at $P^{\prime}$, contradicting the Pareto efficiency of the TTC rule on the lexicographic domain.
It turns out that the TTC rule is not drop strategy-proof when there are three or more agents. [^19] If there are only two agents, then the TTC rule is drop strategy-proof as each agent can only drop objects from the other agent’s endowment.
Example 5.
Let $N=\{1,2,3\}$ and $\omega=\left(\left\{a,b\right\},\left\{c,d\right\},\left\{e,f\right\}\right)$. Let $P_{2}$ and $P_{3}$ be lexicographic preference relations such that $P_{2}:e,f,b,\dots$ and $P_{3}:a,c,d,\dots.$
Let $P_{1}$ be a responsive preference relation such that $\succ^{P_{1}}:c,f,e,d,a,b$ and $\{d,f\}\mathbin{P_{1}}\{c,b\}$. Under the TTC rule, the allocation is $\varphi^{\text{TTC}}\left(P\right)=\left(\left\{c,b\right\},\left\{e,f\right\},\left\{a,d\right\}\right)$. Now, suppose agent $1$ employs a drop strategy $P^{\prime}_{1}$ obtained from $P_{1}$ by dropping object $c$, resulting in the marginal preferences $\succ^{P^{\prime}_{1}}:f,e,d,a,b,c$. The TTC rule then yields $\varphi^{\text{TTC}}\left(P^{\prime}_{1},P_{-1}\right)=\left(\left\{d,f\right\},\left\{e,b\right\},\left\{a,c\right\}\right)$. Because $\varphi_{1}^{\text{TTC}}\left(P^{\prime}_{1},P_{-1}\right)=\{d,f\}\mathbin{P_{1}}\{c,b\}=\varphi_{1}^{\text{TTC}}\left(P\right)$, agent $1$ benefits from the drop strategy. Thus, the TTC rule is not drop strategy-proof on $\mathcal{R}$. $\diamond$
Although agents may benefit by employing drop strategies, they cannot benefit by truncating their preferences. Indeed, we find that the TTC rule is truncation-proof on the responsive domain. This distinction highlights the independence between the two types of manipulation heuristics.
The positive properties of the TTC rule on $\mathcal{R}$ are summarized in the following proposition.
Proposition 3.
On the responsive domain, the TTC rule satisfies ig-efficiency, individual rationality, and truncation-proofness.
We extend our characterization of the TTC rule (Theorem 1) from the lexicographic domain to the responsive domain by adapting a standard technique (see, e.g., Biró et al. (2022a)).
Let $\varphi$ be any individual-good-based rule on $\mathcal{R}$ satisfying balancedness, ig-efficiency, the worst endowment lower bound, and truncation-proofness. An argument similar to the proof of Theorem 1 demonstrates that $\varphi$ must coincide with the TTC rule on $\mathcal{L}$. For any $P^{\prime}\in\mathcal{R}$, let $P\in\mathcal{L}$ be the unique lexicographic preference profile such that $\succ^{P}=\succ^{P^{\prime}}$. Since $\varphi$ and the TTC rule are individual-good-based, we have
$\varphi(P^{\prime})=\varphi(P)=\varphi^{\text{TTC}}(P)=\varphi^{\text{TTC}}(P^{\prime}).$
Therefore, the characterization extends to all of $\mathcal{R}$.
Theorem 3.
On the responsive domain, an individual-good-based rule satisfies
  • balancedness,
  • ig-efficiency,
  • the worst endowment lower bound, and
  • truncation-proofness
if and only if it is the TTC rule.
It turns out that the worst endowment lower bound can be replaced with individual rationality in the statement of Theorem 3. The following theorem is an immediate consequence of Theorem 3, Lemma 1, and the fact that the TTC rule is individually rational.
Theorem 4.
On the responsive domain, an individual-good-based rule satisfies
  • balancedness,
  • ig-efficiency,
  • individual rationality, and
  • truncation-proofness
if and only if it is the TTC rule.
Our characterizations establish the TTC rule as the only individual-good-based rule satisfying a set of desirable properties. However, relaxing the individual-good-based assumption allows for alternative rules. There exist rules that are not individual-good-based but still satisfy all the other properties we consider and, moreover, some of these rules can Pareto-dominate the TTC rule; that is, at every preference profile, they either coincide with the TTC rule or provide a Pareto improvement. We present an example of such a rule below.
Example 6 (A non-individual-good-based rule).
Let $N=\{1,2\}$ and $\omega=(\{a,b\},\{c,d\})$. Define the allocation $\mu\coloneqq(\{c,d\},\{a,b\})$. Consider the rule $\varphi$ on $\mathcal{R}$ defined by
$\varphi\left(P\right)=\begin{cases}\mu,&\text{if }\varphi^{\text{TTC}}(P)=\omega\text{, and }\mu\text{ Pareto-dominates }\omega\text{ at }P;\\ \varphi^{\text{TTC}}(P),&\text{otherwise.}\end{cases}$
This rule satisfies balancedness, ig-efficiency, the worst endowment lower bound, and truncation-proofness. Moreover, for any preference profile $P\in\mathcal{R}$, either $\varphi(P)=\varphi^{\text{TTC}}(P)$ or $\varphi(P)$ Pareto-dominates $\varphi^{\text{TTC}}(P)$ at $P$. However, it is not individual-good-based, illustrating that the property is essential for our characterization.
We note that $\varphi$ coincides with $\varphi^{\text{TTC}}$ on the lexicographic domain, since when $P$ is lexicographic with $\varphi^{\text{TTC}}(P)=\omega$, $\mu$ does not Pareto-dominate $\omega$. $\diamond$
The rule $\varphi$ in Example 6, which is not individual-good-based, satisfies all our desired properties and Pareto-dominates the TTC rule. This highlights an important trade-off between efficiency and ease of implementation. While individual-good-based rules are attractive in practice due to their simplicity and transparency, such simplicity may come at the expense of efficiency.
5 Single-object reallocation (The Shapley-Scarf model)
Our characterization of the TTC rule on the lexicographic domain (Theorem 1) leads to a new characterization in the classic single-object reallocation problem (Shapley and Scarf, 1974), where each agent is endowed with a single object. In this section, we formally introduce the single-object reallocation problem and present our main result: the TTC rule is the unique rule satisfying Pareto efficiency, individual rationality, and truncation-proofness in this setting.
Recall that $N=\{1,2,\ldots,n\}$ is the set of agents, and $O$ is the set of objects, with $|O|=n$. Without loss of generality, we assume that $O=\{o_{1},\dots,o_{n}\}$ and that each agent $i\in N$ is endowed with object $\omega_{i}=o_{i}$. Each agent has (strict) preferences $P_{i}$ over individual objects $O$, and $\mathcal{P}=\prod_{i\in N}\mathcal{P}_{i}$ denotes the domain of preference profiles in which each agent $i$ has strict preferences. [^20] In this section, we adopt the standard convention that each agent has preferences $P_{i}$ defined on individual objects $O$. This minor departure plays no role in the analysis. If instead we maintained the assumption that each agent has lexicographic preferences on $2^{O}$, then our subsequent result would be a direct corollary of Theorem 1. An allocation is a bijection $\mu:N\to O$, represented as $\mu=(\mu_{i})_{i\in N}$, where $\mu_{i}$ is the object assigned to agent $i$.
In this setting, all allocations are balanced by definition since each agent receives exactly one object. Moreover, the worst endowment lower bound is equivalent to individual rationality, and Pareto efficiency coincides with individual-good efficiency.
While the definitions of drop strategy-proofness and truncation-proofness remain the same, it is helpful to reformulate the concept of truncation strategies in this context. Given a preference relation $P_{i}\in\mathcal{P}_{i}$, we say that $P^{\prime}_{i}\in\mathcal{P}_{i}$ is a truncation strategy for $P_{i}$ if
  1. 1.
    $\{x\in O\mid x\mathbin{R^{\prime}_{i}}o_{i}\}\subseteq\{x\in O\mid x\mathbin{R_{i}}o_{i}\}$, and
  2. 2.
    for each $x,y\in O\backslash\{o_{i}\}$, $x\mathbin{P^{\prime}_{i}}y$ if and only if $x\mathbin{P_{i}}y$.
Intuitively, a truncation strategy involves agent $i$ promoting her own object in their preference list while preserving the original ordering of other objects. Let $\mathcal{T}_{i}\left(P_{i}\right)$ denote the set of all truncation strategies for $P_{i}$.
We now present our main result for the single-object reallocation problem.
Theorem 5.
In the single-object reallocation problem, a rule satisfies
  • Pareto efficiency,
  • individual rationality, and
  • truncation-proofness
if and only if it is the TTC rule.
This theorem refines several characterizations of the TTC rule for the single-object reallocation problem. For example, Ma (1994) characterizes the TTC rule using Pareto efficiency, individual rationality, and strategy-proofness, while Altuntaş et al. (2023) show that strategy-proofness can be replaced with drop strategy-proofness. In light of Remark 2 and Lemma 2, which carry over to this single-object environment, we generalize these results by establishing uniqueness under even weaker criteria.
Given that the TTC rule also satisfies strategy-proofness in this environment, Theorem 5 is especially relevant in applications where full strategy-proofness is unnecessarily strong. For example, in Paired Kidney Exchange with strict preferences, patients’ preferences over donors’ kidneys are often common knowledge because they are based on publicly verifiable criteria (Ashlagi and Roth (2014), Nicoló and Rodríguez-Álvarez (2012)). In such settings, concerns about strategic manipulation may be limited to specific forms like truncation. A patient might misrepresent her willingness to remain on dialysis (and retain her donor’s kidney), effectively truncating her preference list. Alternatively, a donor might condition her participation on the expected outcome for the patient she is paired with, agreeing to donate only if the patient receives a sufficiently suitable kidney.
Due to the interest in characterizations of the TTC rule for the Shapley-Scarf model, we provide a direct proof of Theorem 5 in Appendix A. The proof is concise and shares key ideas with the proof of Theorem 1, offering insights into the proof of the more general result.
6 Extension: Conditionally Lexicographic Preferences
In this section, we introduce conditionally lexicographic preferences, a generalization of purely lexicographic preferences. Unlike responsive preferences, conditionally lexicographic preferences allow for the relative ranking of two objects to depend on the other objects they are obtained with. For example, an agent might prefer drinking Champagne to Bordeaux when paired with oysters, but Bordeaux to Champagne otherwise. Thus, conditionally lexicographic preferences are flexible enough to accommodate complementarity among objects. Despite this added flexibility, conditionally lexicographic preferences retain some of the appealing features of lexicographic preferences. Notably, they have a compact representation, which makes them appealing from an implementation perspective. Furthermore, as we shall see, Pareto efficiency is equivalent to ig-efficiency on this domain. Conditionally lexicographic preferences have been widely studied in computer science, particularly in artificial intelligence (e.g., Booth et al. (2010), Domshlak et al. (2011), Pigozzi et al. (2016)).
Loosely speaking, an agent has conditionally lexicographic preferences if, for any bundle $Y\subsetneq O$, there is an object $o\in O\backslash Y$ that she considers the “lexicographically best” addition to $Y$. Formally, conditionally lexicographic preferences are represented by “lexicographic preference trees” on the set $O$ of objects.
Definition 9.
A lexicographic preference tree (LP tree) on $O$ is a rooted directed tree $\tau_{i}$ such that
  • each vertex $v$ is labeled with an object $o(v)\in O$.
  • every object appears exactly once on any path from the root to a leaf.
  • every internal (non-leaf) vertex has two outgoing edges:
    • an “in edge” labeled $o(v)$, representing the presence of $o(v)$ in a bundle;
    • a “not-in edge” labeled $\neg o(v)$, representing the absence of $o(v)$ from a bundle.
Intuitively, an LP tree represents the conditional preferences of an agent in a hierarchical manner. For any given bundle $Y$, there is a unique path from the root to a leaf that is consistent with $Y$: at each vertex, we follow the “in edge” if the corresponding object is in $Y$ and the “not-in edge” if it is not. This path specifies the agent’s preference ordering over objects, conditional on receiving $Y$. Figures 2 and 2 provide a graphical representation of two LP trees, $\tau_{i}$ and $\tau^{*}_{i}$, on the set $O=\{a,b,c,d\}$ of objects.
Given an LP tree $\tau_{i}$ and a bundle $X\subseteq O$, let $\tau_{i}\left(X\right)$ be a directed path from the root to a leaf of $\tau_{i}$ containing only edges consistent with $X$, i.e., edges $\left(v,v^{\prime}\right)$ labeled with $o\left(v\right)$ if $o\left(v\right)\in X$ and $\neg o\left(v\right)$ if $o\left(v\right)\notin X$. Note that such a path $\tau_{i}\left(X\right)$ is unique. Given bundles $A$ and $B$ with $A\neq B$, let $\tau_{i}\left(A,B\right)$ denote the first vertex $v$ visited by both $\tau_{i}\left(A\right)$ and $\tau_{i}\left(B\right)$ and such that $o\left(v\right)\in A\triangle B$, i.e., $o\left(v\right)$ belongs to exactly one of $A$ and $B$. Equivalently, $\tau_{i}\left(A,B\right)$ is the last vertex which is common to both $\tau_{i}\left(A\right)$ and $\tau_{i}\left(B\right)$.
Definition 10.
The preference relation $P_{\tau_{i}}$ associated with an LP tree $\tau_{i}$ is defined as follows:
  1. 1.
    for all $A,B\subseteq O$ with $A\neq B$, $\left[AP_{\tau_{i}}B\iff o\left(\tau_{i}\left(A,B\right)\right)\in A\backslash B\right]$.
  2. 2.
    for all $A,B\subseteq O$, $\left[AR_{\tau_{i}}B\iff\left(A=B\text{ or }AP_{\tau_{i}}B\right)\right]$.
Figure 1: An LP tree $\tau_{i}$
Figure 2: An LP tree $\tau^{*}_{i}$
Notes: In both figures, the paths corresponding to the bundles $\{a,b\}$ and $\{a,c\}$ are highlighted in red and blue, respectively. In Figure 2, the last common vertex, $\tau_{i}(\{a,b\},\{a,c\})$, is highlighted in blue. Because $o(\tau_{i}(\{a,b\},\{a,c\})=c$ belongs to $\{a,c\}\backslash\{a,b\}$, we have $\{a,c\}\mathbin{P_{\tau_{i}}}\{a,b\}$. In Figure 2, the last common vertex, $\tau^{*}_{i}(\{a,b\},\{a,c\})$, is highlighted in red. Because $o(\tau^{*}_{i}(\{a,b\},\{a,c\}))=b$ belongs to $\{a,b\}\backslash\{a,c\}$, we have $\{a,b\}\mathbin{P_{\tau^{*}_{i}}}\{a,c\}$. Since all paths in $\tau^{*}_{i}$ use the same object order, $P_{\tau^{*}_{i}}$ is purely lexicographic. $\diamond$
A preference relation $P_{i}$ on $2^{O}$ is called conditionally lexicographic if there exists an LP tree $\tau_{i}$ such that $P_{i}=P_{\tau_{i}}$. For each $i\in N$, let $\mathcal{CL}_{i}$ be the set of conditionally lexicographic preferences on $2^{O}$. Then $\mathcal{CL}\coloneqq\prod_{i\in N}\mathcal{CL}_{i}$ is called the conditionally lexicographic domain. Given the one-to-one correspondence between LP trees and conditionally lexicographic preferences, we denote the unique LP tree associated with $P_{i}$ by $\tau_{P_{i}}$. When context permits, we may refer to a conditionally lexicographic $P_{i}$ and its corresponding LP tree $\tau_{P_{i}}$ interchangeably.
Note that every conditionally lexicographic preference relation is monotonic. While every lexicographic preference relation is conditionally lexicographic, the converse is not true when $|O|\geq 3$. Specifically, a lexicographic preference relation corresponds to an LP tree where all paths from the root to a leaf use the same object order (i.e., all vertices at the same depth are labeled with the same object). Moreover, a preference relation is lexicographic if and only if it is conditionally lexicographic and responsive. In other words, for any $i\in N$, we have $\mathcal{CL}_{i}\subseteq\mathcal{M}_{i}$, $\mathcal{L}_{i}\subseteq\mathcal{CL}_{i}$ (with strict inclusion for $|O|\geq 3$), and $\mathcal{L}_{i}=\mathcal{CL}_{i}\cap\mathcal{R}_{i}$.
The following proposition provides an alternative characterization of conditionally lexicographic preferences. It states that for any set of objects $X$ not yet in agent $i$’s bundle $Y$, there is a unique object in $X$, denoted $\max_{P_{i}}(X\mid Y)$, which is “lexicographically best” among $X$ conditional on receiving $Y$. Its proof is straightforward and it is omitted.
Proposition 4.
A preference relation $P_{i}$ on $2^{O}$ is conditionally lexicographic if and only if the following property holds: for all disjoint bundles $X,Y\in 2^{O}$ with $X\neq\emptyset$, there is a unique object in $x^{*}\in X$, denoted $x^{*}=\max_{P_{i}}(X\mid Y)$, such that
for all $\text{for all }Z\subseteq X\backslash\{x^{*}\},\quad(Y\cup x^{*})\mathbin{P_{i}}(Y\cup Z).$ (4)
That is, $\max_{P_{i}}(X\mid Y)$ is agent $i$’s (lexicographically) most-preferred object in $X$ conditional on already having $Y$ [^21] We note that $\max_{P_{i}}(X\mid Y)=o(\tau_{P_{i}}(Y,Y\cup X))$.
6.1 Properties of allocation rules
In what follows, we drop the requirement that the allocation depends solely on agents’ marginal preferences over individual objects. That is, we consider allocation rules on $\mathcal{CL}$ that are not necessarily individual-good-based.
Most of the properties defined in Section 2.2—namely balancedness, Pareto efficiency, ig-efficiency, and individual rationality—are defined exactly as before. However, the definitions of the worst endowment lower bound and drop strategy-proofness require adaptation to accommodate conditionally lexicographic preferences. [^22] We do not study truncation-proofness in this setting, as the definition becomes rather unwieldy.
To define the worst endowment lower bound, we first need some notation. The idea is to ensure that no agent is assigned a bundle containing an object that she considers conditionally worse than “conditionally worst” object in her own endowment, given the bundle she receives.
Given an LP tree $\tau_{i}\in\mathcal{CL}_{i}$ and a vertex $v$ of $\tau_{i}$, let $a\left(v\right)$ denote the set of objects labeling the ancestors of $v$, including $v$ itself. That is, if $\left(v_{1},v_{2},\dots,v_{k}=v\right)$ is the unique path from the root of $\tau_{i}$ to $v$, then $a\left(v\right)=\left\{o\left(v_{1}\right),o\left(v_{2}\right),\dots,o\left(v_{k}\right)\right\}$. Given a preference relation $P_{i}\in\mathcal{CL}_{i}$ and a bundle $X\in 2^{O}$, define $w_{P_{i}}\left(\omega_{i}\mid X\right)$ as the unique vertex on the path $\tau_{P_{i}}(X)$ that corresponds to the “conditionally worst” object in agent $i$’s endowment $\omega_{i}$, given $X$. Formally, $w_{P_{i}}\left(\omega_{i}\mid X\right)$ is the unique vertex on $\tau_{P_{i}}\left(X\right)$ such that $o\left(w_{P_{i}}\left(\omega_{i}\mid X\right)\right)\in\omega_{i}$ and $\omega_{i}\subseteq a\left(w_{P_{i}}\left(\omega_{i}\mid X\right)\right)$. This means that $w_{P_{i}}(\omega_{i}\mid X))$ is the last vertex on the path $\tau_{P_{i}}(X)$ to be labeled with an object in $\omega_{i}$. We illustrate this construction graphically in figures 4 and 4.
An allocation $\mu$ satisfies the worst endowment lower bound at a preference profile $P\in\mathcal{CL}$ if, for each $i\in N$, $\mu_{i}\subseteq a\left(w_{P_{i}}\left(\omega_{i}\mid\mu_{i}\right)\right)$. This means that every object in $\mu_{i}$ labels a vertex on the path $\tau_{P_{i}}\left(\mu_{i}\right)$ that occurs before or at the vertex $w_{P_{i}}\left(\omega_{i}\mid\mu_{i}\right)$. Intuitively, no agent $i$ is assigned an object that she considers conditionally worse than the “conditionally worst” object in $\omega_{i}$, given the bundle $\mu_{i}$. Note that when agents have purely lexicographic preferences, this definition aligns with the original worst endowment lower bound.
Definition 11.
A rule $\varphi$ satisfies the worst endowment lower bound if, for each $P\in\mathcal{CL}$, $\varphi(P)$ satisfies the worst endowment lower bound at $P$.
Figure 3: The LP tree $\tau_{i}$
Figure 4: The LP tree $\tau^{*}_{i}$
Notes: Figures 4 and 4 illustrate the construction of $w_{P_{i}}(\omega_{i}\mid X)$. Suppose agent $i$’s endowment is $\omega_{i}=\{b,c\}$. In both figures, the vertices corresponding to $w_{P_{i}}(\omega_{i}\mid X)$ for various $X\in 2^{O}$ are highlighted in blue. Notice that on any path from the root to a leaf, there is exactly one such vertex. Under $\tau_{i}$, the worst endowment lower bound requires that agent $i$ is not assigned object $d$ together with any of the bundles ,$\emptyset,\{a\},\{a,b\}$, or $\{c\}$; that is, she does not receive a bundle in $\{\{d\},\{a,d\},\{a,b,d\},\{c,d\}\}$. Under $\tau^{*}_{i}$, it requires that agent $i$ does not receive any bundle containing object $d$. $\diamond$
The concept of drop strategies extends naturally to conditionally lexicographic preferences. An agent employs a drop strategy by dropping an object she does not own to the bottom of her LP tree. Formally, given a preference relation $P_{i}\in\mathcal{CL}_{i}$, we say that a preference relation $P^{\prime}_{i}\in\mathcal{CL}_{i}$ is a drop strategy for $P_{i}$ if there exists $x\in O\backslash\omega_{i}$ such that:
  1. 1.
    for all nonempty bundles $Y\subseteq O\backslash\{x\}$, $Y\mathbin{P^{\prime}_{i}}x$; and
  2. 2.
    for any $Y,Z\subseteq O\backslash\{x\}$, $Y\mathbin{P^{\prime}_{i}}Z$ if and only if $Y\mathbin{P_{i}}Z$.
In this case, we say that $P^{\prime}_{i}$ is obtained from $P_{i}$ by dropping object $x$ [^23] As is the case for purely lexicographic preferences (but not responsive preferences), there is exactly one $P^{\prime}_{i}$ obtained from $P_{i}$ by dropping object $x$. Note that this definition aligns with the original one from Section 2.2 whenever $P_{i}$ is purely lexicographic. Let $\mathcal{D}_{i}(P_{i})$ denote the set of all drop strategies for $P_{i}$.
The drop strategy $P^{\prime}_{i}$ obtained from $P_{i}$ by dropping object $x$ can also be represented via LP trees. Let $\tau_{P_{i}}^{-x}$ denote the LP tree on $O\backslash\{x\}$ representing the restriction of $P_{i}$ to subsets of $O\backslash\{x\}$. [^24] We can construct $\tau_{P_{i}}^{-x}$ as follows. For each vertex $v$ of $\tau_{P_{i}}$, let $T_{v}$ denote the maximal subtree of $\tau_{P_{i}}$ consisting of a vertex $v$ of $\tau_{P_{i}}$ together with all of its successors, and let $v^{\prime}$ be the child of $v$ whose incoming edge $(v,v^{\prime})$ is labeled with $\neg o(v)$. For each vertex $v$ of $\tau_{P_{i}}$ such that $o(v)=x$, simply replace $T_{v}$ with the subtree $T_{v^{\prime}}$ (or the empty tree if $v$ is a leaf of $\tau_{P_{i}}$). To construct $\tau_{P^{\prime}_{i}}$, at each leaf of $\tau_{P_{i}}^{-x}$ append two child nodes labeled with $x$, connected by an “in edge” and a “not-in edge” accordingly. This modification places $x$ and the bottom of the LP tree, ensuring that $x$ is the “lexicographically worst” addition to any bundle. Figures 6 and 6 illustrate this construction.
Figure 5: The LP tree $\tau_{i}$
Figure 6: The LP tree $\tau^{\prime}_{i}$ obtained by dropping $b$
Notes: Figure 6 displays the original LP tree from Figures 2 and 4. To illustrate the drop strategy where agent $i$ drops object $b$, we construct the LP tree $\tau^{\prime}_{i}$ shown in Figure 6. At each internal vertex $v$ labeled $b$, we do the following: (i) remove the subtree consisting of $v$ and all of its descendants, and (ii) append the subtree, shown in blue, consisting of the child of $v$ reached via the “not-in edge” together with all of its descendants. Then, at each leaf of this modified tree, we append two child nodes labeled $b$, connected via an “in edge” and a “not-in edge” accordingly. $\diamond$
Definition 12.
A rule $\varphi$ is drop strategy-proof if, for each $P\in\mathcal{CL}$, each $i\in N$, and each $P^{\prime}_{i}\in\mathcal{D}_{i}(P_{i})$, $\varphi_{i}(P)\mathbin{R_{i}}\varphi_{i}(P^{\prime}_{i},P_{-i})$.
6.2 Augmented Top Trading Cycles
The Augmented Top Trading Cycles (ATTC) rule, introduced by Fujita et al. (2018), is the natural extension of the TTC rule from the lexicographic domain to the conditionally lexicographic domain.
At each step of the ATTC algorithm, every agent points to her most-preferred unassigned object, conditional on the objects already assigned to her, and every unassigned object points to its owner. There exists at least one cycle and, as in the TTC algorithm, we execute only the cycle containing the minimum agent. All objects involved in this cycle are then removed. If unassigned objects remain, then the procedure continues to the next step; otherwise, it terminates with the corresponding allocation.
Formally, given a preference profile $P\in\mathcal{P}$, the Augmented Top Trading Cycles (ATTC) rule selects the allocation $\varphi^{\text{ATTC}}(P)$ determined by the ATTC algorithm at $P$. We denote this algorithm as $\text{ATTC}(P)$.
Algorithm:
$\text{ATTC}(P)$
Input: A preference profile $P\in\mathcal{CL}$.
Output: An allocation $\varphi^{\text{ATTC}}\left(P\right)$.
Initialization:
Set $\mu^{0}\coloneqq\left(\emptyset\right)_{i\in N}$ and $O^{1}\coloneqq O$.
Step $t\geq 1$:
  1. 1.
    (Graph construction) Construct a bipartite directed graph with independent vertex sets $N$ and $O^{t}$, and edge sets defined as follows:
    1. (a)
      For each agent $i\in N$, there is a directed edge from $i$ to $\max_{P_{i}}(O^{t}\mid\mu^{t-1}_{i})$.
    2. (b)
      For each object $o\in O^{t}$, there is a directed edge from $o$ to its owner, $\omega^{-1}\left(o\right)$.
  2. 2.
    (Cycle selection) There is at least one cycle.
    1. (a)
      Let $\mathcal{C}_{t}(P)$ denote the set of cycles that arise at Step $t$.
    2. (b)
      Let $C_{t}(P)$ be the cycle in $\mathcal{C}_{t}(P)$ containing the minimum agent:
      i.e., $\min N(C_{t}(P))\leq\min N(C)$ for all $C\in\mathcal{C}_{t}(P)$.
  3. 3.
    (Assignment) Assign to each agent $i\in N(C_{t}(P))$ the object $\max_{P_{i}}(O^{t}\mid\mu^{t-1}_{i})$. That is, let $\mu^{t}=\left(\mu_{i}^{t}\right)_{i\in N}$ be such that
    1. (a)
      for all $i\in N(C_{t}(P))$, $\mu_{i}^{t}=\mu_{i}^{t-1}\cup\left\{\max_{P_{i}}(O^{t}\mid\mu^{t-1}_{i})\right\}$, and
    2. (b)
      for all $i\in N\backslash N(C_{t}(P))$, $\mu_{i}^{t}=\mu_{i}^{t-1}$.
  4. 4.
    (Removal) Let $O^{t+1}\coloneqq O^{t}\backslash O(C_{t}(P))$ be the set of objects remaining at Step $t+1$.
    1. (a)
      If $O^{t+1}\neq\emptyset$, then proceed to Step $t+1$.
    2. (b)
      If $O^{t+1}=\emptyset$, then proceed to Termination.
Termination:
Because $O$ is finite and $\left|O^{1}\right|>\left|O^{2}\right|>\cdots>\left|O^{t}\right|$, the algorithm terminates at some step $T$. Return the allocation $\varphi^{\text{ATTC}}\left(P\right)\coloneqq\mu^{T}$.
6.3 Results for Conditionally Lexicographic Preferences
Despite its additional flexibility, the conditionally lexicographic domain retains some of the desirable features of the lexicographic domain. In particular, Pareto efficiency and ig-efficiency are equivalent on this domain, making Pareto efficient allocations relatively easy to identify. This equivalence is formalized in the following proposition.
Proposition 5.
On the conditionally lexicographic domain, an allocation is ig-efficient if and only if it is Pareto efficient.
This equivalence implies that myopic procedures like the ATTC algorithm, which greedily execute Pareto-improving single-object exchanges until no such exchange remains, will always result in Pareto efficient allocations on this domain. Moreover, Fujita et al. (2018) establish that the ATTC rule is core selecting, a property stronger than both Pareto efficiency and individual rationality. They also show that it is NP-hard for an agent to manipulate the ATTC rule, and even if a beneficial manipulation is found, the gains from manipulating are bounded. Specifically, no manipulation can yield an object that the agent prefers over her most-preferred object obtained by truth-telling. Complementing the results of Fujita et al. (2018), we establish that the ATTC rule is drop strategy-proof on the conditionally lexicographic domain.
Proposition 6.
On the conditionally lexicographic domain, the ATTC rule is drop strategy-proof.
Proposition 6 extends the work of Altuntaş et al. (2023), who established that the TTC rule is drop strategy-proof on the lexicographic domain. Unlike in their paper, we consider a weak notion of drop strategy-proofness that defends only against manipulations where agents drop objects they do not own. Nonetheless, the ATTC rule actually satisfies the stronger notion, which also defends against manipulations where agents drop objects they do own. However, for our characterization result, the weak version suffices.
Theorem 6.
On the conditionally lexicographic domain, a rule satisfies
  • balancedness,
  • ig-efficiency (or Pareto efficiency),
  • the worst endowment lower bound, and
  • drop strategy-proofness
if and only if it is the ATTC rule.
Theorem 6 states that the ATTC rule is categorically determined by balancedness, Pareto efficiency, the worst endowment lower bound, and drop strategy-proofness, addressing an open question posed by Fujita et al. (2018, p. 531). This effectively extends our previous characterization of the TTC rule (Theorem 2) from the lexicographic domain to the broader conditionally lexicographic domain.
Beyond Conditionally Lexicographic Preferences
We conclude this section by highlighting the difficulty in extending our analysis beyond the conditionally lexicographic domain. On any domain comprising monotonic preferences that is larger than the conditionally lexicographic domain, the equivalence between Pareto efficiency and ig-efficiency breaks down. [^25] It is straightforward to strengthen Proposition 7 if we broaden our definition of allocations to allow agents to be assigned empty bundles. In this case, the conditionally lexicographic domain is a maximal domain among all strict preferences on which the two notions coincide.
Proposition 7.
Within the domain of monotonic preferences, the conditionally lexicographic domain is a maximal domain on which ig-efficiency and Pareto efficiency are equivalent.
(That is, if $\mathcal{CL}\subsetneq\mathcal{P}\subseteq\mathcal{M}$, then there is a set $N$ of agents, a preference profile $P\in\mathcal{P}$, and an allocation $\mu\in\mathcal{A}$ such that $\mu$ is ig-efficient but not Pareto efficient at $P$.)
7 Conclusion
In this paper, we provide several characterizations of generalized versions of Gale’s TTC rule for multi-object reallocation problems. Specifically, we present the first characterizations of the TTC rule on the responsive preference domain and the first characterization of the ATTC rule on the conditionally lexicographic preference domain.
Our analysis sheds light on the trade-offs involved in multi-object (re)allocation with responsive preferences. Given the inherent incompatibility among efficiency, individual rationality, and strategic robustness, other prominent allocation rules typically fulfill only two of the three objectives, often neglecting the third. Although the TTC rule does not meet the most stringent efficiency and incentive requirements, it satisfies ig-efficiency and truncation-proofness, in addition to individual rationality. Furthermore, within the class of balanced and individual-good-based rules, there is no other rule that meets these criteria. Thus, the TTC rule performs remarkably well according to all three objectives.
Our work highlights the potential of the TTC and ATTC rules in settings where agents have complex preferences but practical considerations necessitate simple reporting languages.
Appendix A Omitted proofs
A.1 Proof of Lemma 1
Suppose $\varphi$ is an individual-good-based rule satisfying balancedness and individual rationality.
Let $P\in\mathcal{R}$ be a profile of responsive preferences. Toward contradiction, suppose $\mu\coloneqq\varphi(P)$ violates the worst endowment lower bound at $P$. Then there is some agent $i\in N$ such that $\mu_{i}\nsubseteq\{o\in O\mid o\mathbin{R_{i}}\min_{P_{i}}(\omega_{i})\}$. By balancedness, we must have $|\mu_{i}|=|\omega_{i}|$. Because $\min_{P_{i}}(\mu_{i})$ is worse than $\min_{P_{i}}(\omega_{i})$ at $P_{i}$, there is a responsive preference relation $P^{\prime}_{i}\in\mathcal{R}_{i}$ such that $\succ^{P^{\prime}_{i}}=\succ^{P_{i}}$ and $\omega_{i}\mathbin{P^{\prime}_{i}}\mu_{i}$. By individual rationality, we must have $\varphi_{i}(P^{\prime}_{i},P_{-i})\neq\mu_{i}$. However, $\varphi(P^{\prime}_{i},P_{-i})=\varphi(P)=\mu$ because $\varphi$ is individual-good-based. This is a contradiction. $\blacksquare$
A.2 Proof of Lemma 2
Suppose $\varphi$ satisfies drop strategy-proofness and the worst endowment lower bound. Let $P^{\prime}_{i}$ be a subset drop strategy obtained from $P_{i}$ by dropping some subset $X\subseteq O\backslash\omega_{i}$. Suppose that $X=\left\{x_{1},x_{2},\dots,x_{k}\right\}$, where $x_{1}\mathbin{P_{i}}x_{2}\mathbin{P_{i}}\cdots\mathbin{P_{i}}x_{k}$. Then $P^{\prime}_{i}$ is obtained from $P_{i}$ by successively performing $k$ drop strategies. That is, $P^{\prime}_{i}=P_{i}^{k}$, where $P_{i}^{0}=P_{i}$ and $P_{i}^{1},\dots,P_{i}^{k}$ are such that, for each $\ell\in\left\{1,\dots,k\right\}$, $P_{i}^{\ell}$ is obtained from $P_{i}^{\ell-1}$ by dropping object $x_{\ell}$.
Claim 1.
For each $\ell\in\{1,\dots,k\}$, $\varphi_{i}(P_{i}^{\ell-1},P_{-i})\mathbin{R_{i}}\varphi_{i}(P_{i}^{\ell},P_{-i})$.
Proof of Claim 1.
The proof is by induction on $\ell$. Clearly, $\varphi_{i}(P)=\varphi_{i}(P_{i}^{0},P_{-i})\mathbin{R_{i}}\varphi_{i}(P_{i}^{1},P_{-i})$ by drop strategy-proofness. For the inductive step, suppose that $\ell\in\{1,\dots,k-1\}$ is such that
$\varphi_{i}(P)=\varphi_{i}(P_{i}^{0},P_{-i})\mathbin{R_{i}}\varphi_{i}(P_{i}^{1},P_{-i})\mathbin{R_{i}}\cdots\mathbin{R_{i}}\varphi_{i}(P_{i}^{\ell},P_{-i}).$
It suffices to show that $\varphi_{i}(P_{i}^{\ell},P_{-i})\mathbin{R_{i}}\varphi_{i}(P_{i}^{\ell+1},P_{-i})$. By drop strategy-proofness, we have $\varphi_{i}(P_{i}^{\ell},P_{-i})\mathbin{R_{i}^{\ell}}\varphi_{i}(P_{i}^{\ell+1},P_{-i})$. Moreover, the worst endowment lower bound implies that
$\varphi_{i}(P_{i}^{\ell},P_{-i})\subseteq O\backslash\{x_{1},\dots,x_{\ell}\}\quad\text{and}\quad\varphi_{i}(P_{i}^{\ell+1},P_{-i})\subseteq O\backslash\{x_{1},\dots,x_{\ell+1}\}\subseteq O\backslash\{x_{1},\dots,x_{\ell}\}.$
Because $\succ^{P_{i}^{\ell}}$ agrees with $\succ^{P_{i}}$ on $O\backslash\{x_{1},\dots,x_{\ell}\}$, and each of $P_{i}$ and $P^{\ell}_{i}$ are lexicographic, we have $\varphi_{i}(P_{i}^{\ell},P_{-i})\mathbin{R_{i}}\varphi_{i}(P_{i}^{\ell+1},P_{-i})$, as desired. ∎
It follows from Claim 1 that
$\varphi_{i}(P)=\varphi_{i}(P_{i}^{0},P_{-i})\mathbin{R_{i}}\varphi_{i}(P_{i}^{1},P_{-i})\mathbin{R_{i}}\cdots\mathbin{R_{i}}\varphi_{i}(P_{i}^{k},P_{-i})=\varphi_{i}(P^{\prime}_{i},P_{-i}).$
Therefore, $\varphi_{i}(P)\mathbin{R_{i}}\varphi_{i}(P^{\prime}_{i},P_{-i})$. $\blacksquare$
A.3 Proof of Theorem 1
Toward contradiction, suppose that $\varphi$ satisfies the properties but $\varphi\neq\varphi^{\text{TTC}}$.
For each $P\in\mathcal{P}^{N}$, and each $t\in\mathbb{N}$, let $\mathcal{C}_{t}(P)$ be the set of cycles that obtain at step $t$ of $\text{TTC}\left(P\right)$. Similarly, let $C_{t}(P)\in\mathcal{C}_{t}(P)$ be the cycle that is executed at step $t$ of $\text{TTC}\left(P\right)$.
Denote the size of a profile $P$ by $s\left(P\right)=\sum_{i\in N}\left|\left\{o\in O\mid oR_{i}\min_{P_{i}}\left(\omega_{i}\right)\right\}\right|$.
The similarity between $\varphi$ and $\varphi^{\text{TTC}}$ is a function $\rho:\mathcal{P}^{N}\to\mathbb{N}$ such that, for all $P\in\mathcal{P}$, $\rho(P)$ is defined as follows.
  • Step 1. If $\varphi(P)$ does not execute $C_{1}(P)$, then $\rho(P)=1$. Suppose $\varphi(P)$ executes $C_{1}(P)$. If $\text{TTC}(P)$ terminates at Step 1, then $\rho(P)=|O|+1$; otherwise, proceed to Step $2$.
  • Step $t$ ($\geq 2$). If $\varphi(P)$ does not execute $C_{t}(P)$, then $\rho(P)=t$. Suppose $\varphi(P)$ executes $C_{t}(P)$. If $\text{TTC}(P)$ terminates at Step $t$, then $\rho(P)=|O|+1$; otherwise, proceed to Step $t+1$.
Note that, for all $P\in\mathcal{P}^{N}$, (i) $\rho(P)\leq|O|+1$, and (ii) $\rho(P)=|O|+1$ if and only if $\varphi(P)=\varphi^{\text{TTC}}(P)$.
Let $P$ be a profile such that for any other profile $\tilde{P}\neq P$, either
  • (1)
    $\rho(P)<\rho(\tilde{P})$, or
  • (2)
    $\rho(P)=\rho(\tilde{P})$ and $s(P)\leq s(\tilde{P})$.
Let $t\coloneqq\rho(P)$. Since $\rho(P)=t$, $\varphi(P)$ executes the cycles $C_{1}(P),C_{2}(P),\dots,C_{t-1}(P)$ but not $C_{t}(P)$. Let $C\coloneqq C_{t}(P)$ and $O^{t}\coloneqq O\backslash\bigcup_{\tau=1}^{t-1}O\left(C_{\tau}\left(P\right)\right)$. Suppose that
$C=\left(i_{1},o_{i_{2}},i_{2},o_{i_{3}},\dots,i_{k-1},o_{i_{k}},i_{k},o_{i_{k+1}},i_{k+1}=i_{1}\right).$
Because $\varphi(P)$ does not execute $C$, there is an agent $i_{\ell}\in N(C)$ such that, although agent $i_{\ell}$ points to $o_{i_{\ell+1}}$ on $C$, she does not receive $o_{i_{\ell+1}}$ at $\varphi(P)$, i.e., $o_{i_{\ell+1}}\notin\varphi_{i}(P)$. Note that, by the definition of $\text{TTC}(P)$, $o_{i_{\ell+1}}$ is agent $i_{\ell}$’s most-preferred object in $O^{t}$ at $P_{i_{\ell}}$, i.e., $o_{i_{\ell+1}}=\max_{P_{i_{\ell}}}\left(O^{t}\right)$. Without loss of generality, let $i_{\ell}=i_{k}$. Thus, $o_{i_{1}}\notin\varphi_{i_{k}}(P)$.
Let $X_{i_{k}}$ be the set of all objects in $O\backslash O^{t}$ that are assigned to agent $i_{k}$, i.e., $X_{i_{k}}=\varphi_{i_{k}}(P)\backslash O^{t}$. We next show that, apart from the objects in $X_{i_{k}}$, agent $i_{k}$ only receives the remainder of her endowment.
Claim 2.
$\varphi_{i_{k}}(P)\cap O^{t}=\omega_{i_{k}}\cap O^{t}$.
Proof of Claim 2.
Suppose otherwise. By balancedness and the fact that $\left|\varphi_{i_{k}}\left(P\right)\backslash O^{t}\right|=\left|\omega_{i_{k}}\backslash O^{t}\right|$, we must have $\left|\varphi_{i_{k}}\left(P\right)\cap O^{t}\right|=\left|\omega_{i_{k}}\cap O^{t}\right|$. Consequently, $\varphi_{i_{k}}\left(P\right)\cap O^{t}\neq\omega_{i_{k}}\cap O^{t}$ implies there is an object $o^{\prime}\in\varphi_{i_{k}}(P)\cap O^{t}$ with $o^{\prime}\not\in\omega_{i_{k}}$.
By the definition of $\text{TTC}(P)$, we know that $o_{i_{1}}$ is agent $i_{k}$’s most-preferred object in $O^{t}$, i.e., for each $o\in O^{t}$, $o_{i_{1}}\mathbin{R_{i_{k}}}o$. Because $o^{\prime}\in\varphi_{i_{k}}(P)\cap O^{t}$ and $o_{i_{1}}\notin\varphi_{i_{k}}(P)$, we have $o_{i_{1}}\mathbin{P_{i_{k}}}o^{\prime}$. By the worst endowment lower bound,
$o^{\prime}\mathbin{P_{i_{k}}}w_{i_{k}}(P_{i_{k}}).$
Let $P^{\prime}_{i_{k}}$ be the truncation of $P_{i_{k}}$ at $o_{i_{1}}$. That is, $P^{\prime}_{i_{k}}$ is obtained from $P_{i_{k}}$ by dropping the set $\{o\in O\backslash\omega_{i_{k}}\mid o_{i_{1}}\mathbin{P_{i_{k}}}o\}$. [^26] If we use drop strategy-proofness, we can consider another preference $P^{\prime}_{i_{k}}$ obtained from $P_{i_{k}}$ by dropping object $o^{\prime}$. Let $P^{\prime}\coloneqq(P^{\prime}_{i_{k}},P_{-i_{k}})$.
By the definition of $\text{TTC}(P)$, we see that until step $t$, all top trading cycles that are executed via TTC at $P$ and $P^{\prime}$ are the same, i.e., $C_{\tau}(P)=C_{\tau}(P^{\prime})$ for $\tau=1,\ldots,t$. By the selection of $P$, we know that $\rho(P)=t\leq\rho(P^{\prime})$. Thus, all cycles in $\{C_{1}(P),\dots,C_{t-1}(P)\}$ are executed at $\varphi(P^{\prime})$. Therefore,
$X_{i_{k}}\subseteq\varphi_{i_{k}}(P^{\prime}).$
Since $\varphi$ is truncation-proof, we know that $i_{k}$ cannot be better off by misreporting $P^{\prime}_{i_{k}}$ at $P$. Together with $X_{i_{k}}\subseteq\varphi_{i_{k}}(P^{\prime})$ and $o_{i_{1}}\notin\varphi_{i_{k}}(P)$, we conclude that $i_{k}$ cannot receive $o_{i_{1}}$ at $\varphi(P^{\prime})$, i.e., $o_{i_{1}}\notin\varphi_{i_{k}}(P^{\prime})$. Therefore, $C$ is not executed at $\varphi(P^{\prime})$, which means that $\rho(P^{\prime})\leq t$. As a result, we find that $\rho(P^{\prime})=t$. However, since $P^{\prime}_{i_{k}}$ is obtained from $P_{i_{k}}$ by dropping the subset $\{o\in O\backslash\omega_{i_{k}}\mid o_{i_{1}}\mathbin{P_{i_{k}}}o\}$, and $o_{i_{1}}\mathbin{P_{i_{k}}}o^{\prime}\mathbin{P_{i_{k}}}w_{i_{k}}(P_{i_{k}})$, we see that $s(P^{\prime})<s(P)$. This contradicts the choice of $P$. ∎
Next, we show that at least two agents are involved in $C$.
Claim 3.
$|N(C)|>1$.
Proof.
Toward contradiction, suppose that $|N(C)|=1$, that is to say, $i_{1}=i_{k}$ and $C=(i_{1},o_{i_{1}},i_{1})$. Then $o_{i_{1}}\in\omega_{i_{k}}$ and Claim 2 implies that $o_{i_{1}}\in\varphi_{i_{k}}(P)$, a contradiction. ∎
By Claim 2 and Claim 3, agent $i_{k-1}$ points to $o_{i_{k}}$ on $C$, yet she does not receive $o_{i_{k}}$. An argument similar to the proof of Claim 2 implies that $\varphi_{i_{k-1}}(P^{\prime})\cap O^{t}=\omega_{i_{k-1}}\cap O^{t}$.
Proceeding by induction, we conclude that for each agent $i_{\ell}$ that is involved in $C$, we have $\varphi_{i_{\ell}}(P)\cap O^{t}=\omega_{i_{\ell}}\cap O^{t}$.
However, this means that $\varphi$ is not ig-efficient (or Pareto efficient), as agents in $N(C)$ can benefit by execution of $C$. This contradiction completes the proof of Theorem 1. $\blacksquare$
A.4 Proof of Proposition 3
We only show that the TTC rule is truncation-proof as other properties can be easily verified.
Let $P\in\mathcal{R}$ be any profile of responsive preferences. Let $i\in N$, $x\in O\setminus\omega_{i}$, and let $P^{\prime}_{i}$ be a truncation of $P_{i}$ at $x$. Let $P^{\prime}\coloneqq(P^{\prime}_{i},P_{-i})$. Let $U\coloneqq\{o\in O\mid o\mathbin{R_{i}}x\}$. There are two cases.
Case 1:
Suppose $\varphi_{i}^{\text{TTC}}(P)\subseteq U$. Then, by the definition of TTC, we know that all top trading cycles that are obtained at $P$ and $P^{\prime}$ are exactly the same. As a result, $\varphi_{i}^{\text{TTC}}(P)=\varphi_{i}^{\text{TTC}}(P^{\prime})$. So agent $i$ is not better off by misreporting $P^{\prime}_{i}$.
Case 2:
Suppose $\varphi_{i}^{\text{TTC}}(P)\nsubseteq U$. Then agent $i$ receives at least one object in $O\setminus U$ at $\varphi^{\text{TTC}}(P)$. Assume that at $\text{TTC}(P)$, the earliest step at which agent $i$ points to an object in $O\setminus U$ is $t$.
Let $O_{t}\subseteq O$ be the set of objects that are remaining at step $t$ of $\text{TTC}(P)$. Since $\succ^{P_{i}}$ and $\succ^{P^{\prime}_{i}}$ agree on $U$, [^27] Formally, $\succ^{P_{i}}\cap(U\times U)=\succ^{P^{\prime}_{i}}\cap(U\times U)$. we know that the top trading cycles that are executed at steps $11,\dots,t-$ 1 of $\text{TTC}(P)$ and $\text{TTC}(P^{\prime})$ are the same. Thus, at step $t$ of $\text{TTC}(P^{\prime})$, the set of remaining objects is also $O_{t}$. Moreover, we have
$\varphi^{\text{TTC}}_{i}(P)\setminus O_{t}=\varphi^{\text{TTC}}_{i}(P^{\prime})\setminus O_{t}.$ (5)
By the definition of $P^{\prime}_{i}$, we know that all objects in $O\setminus U=\{o\in O\setminus\omega_{i}\mid x\mathbin{P_{i}}o\}$ are ranked below $i$’s least-preferred object in $\omega_{i}$ at $\succ^{P^{\prime}_{i}}$ ($i$’s least-preferred object in $\omega_{i}$ is the same at $P_{i}$ and $P^{\prime}_{i}$). Thus, by the definition of TTC, from step $t$ of $\text{TTC}(P^{\prime})$, agent $i$ only points to (and hence receives) her endowed objects. That is,
$\varphi^{\text{TTC}}_{i}(P^{\prime})\cap O_{t}=\omega_{i}\cap O_{t}.$ (6)
Again, by the definition of TTC, from step $t$ of $\text{TTC}(P)$, agent $i$ only points to (and hence receives) some objects that are weakly better (with respect to $\succ^{P_{i}}$) than her endowed objects. In other words, there is a bijection $\sigma:\omega_{i}\cap O_{t}\to\varphi^{\text{TTC}}_{i}(P)\cap O_{t}$ such that for each $o\in\omega_{i}\cap O_{t}$, $\sigma(o)\mathbin{\succeq^{P_{i}}}o$. Together with (5) and (6), we conclude that there is a bijection $\pi:\varphi_{i}^{\text{TTC}}(P^{\prime})\to\varphi_{i}^{\text{TTC}}(P)$ such that
for each $\text{for each }o\in\varphi^{\text{TTC}}_{i}(P^{\prime}),\pi(o)\mathbin{\succeq^{P_{i}}}o.$ (7)
Finally, recall that $i$’s original preference relation $P_{i}$ is responsive. Therefore, we must have $\varphi^{\text{TTC}}_{i}(P)\mathbin{R_{i}}\varphi^{\text{TTC}}_{i}(P^{\prime})$. [^28] Starting from $\varphi^{\text{TTC}}_{i}(P^{\prime})$, replace each object $o\in\varphi^{\text{TTC}}_{i}(P^{\prime})$ with object $\pi(o)$, one at a time, and apply the definition of responsiveness. Thus, agent $i$ is not better off by misreporting $P^{\prime}_{i}$. $\blacksquare$
A.5 Proof of Proposition 6
Let $P\in\mathcal{CL}^{N}$, $i\in N$, and let $\overline{P}_{i}\in\mathcal{D}\left(P_{i}\right)$. Let $\overline{P}\coloneqq\left(\overline{P}_{i},P_{-i}\right)$ and suppose that $\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)R_{i}\varphi_{i}^{\text{ATTC}}\left(P\right)$. It suffices to show that $\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)=\varphi_{i}^{\text{ATTC}}\left(P\right)$.
Let $\varphi_{i}^{\text{ATTC}}\left(P\right)=\left\{x_{1},\dots,x_{m}\right\}\eqqcolon X$. By relabeling the objects if necessary, we can assume that, for each $k\in\{1,\dots,m\}$, $x_{k}$ is the $k$th object assigned to agent $i$ during $\text{ATTC}\left(P\right)$. Let $\tau_{i}=\tau_{P_{i}}$ denote the LP tree associated with $P_{i}$. For each $k\in\left\{1,\dots,m\right\}$, let $v_{k}$ denote the unique vertex of $\tau_{i}\left(X\right)$ whose label is $x_{k}$. This means that
for $\text{for each }k\in\left\{1,\dots,m\right\},\quad X\cap a\left(v_{k}\right)=\left\{x_{1},\dots,x_{k}\right\}.$
Let $\overline{X}\coloneqq\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)$. We show by induction that
for $\text{for each }k\in\left\{1,\dots,m\right\},\quad\overline{X}\cap a\left(v_{k}\right)=\left\{x_{1},\dots,x_{k}\right\}.$
Base case: $k=1$.
Consider any object $x\in a\left(v_{1}\right)\backslash\text{$\left\{x_{1}\right\}$}$. That is, $x$ labels one of the (strict) ancestors of $v_{1}$ in $\tau_{i}\left(X\right)$, and $x\notin\varphi_{i}^{\text{ATTC}}\left(P\right)$. This means that, although agent $i$ may point to $x$ during $\text{ATTC}\left(P\right)$, agent $i$ is not on the cycle of $\text{ATTC}\left(P\right)$ that contains object $x$. It follows that agent $i$ does not belong to the cycle of $\text{ATTC}\left(\overline{P}\right)$ that contains object $x$ either. Hence, $x\notin\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)$.
Now consider object $x_{1}$. Because $\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)R_{i}\varphi_{i}^{\text{ATTC}}\left(P\right)$ and $\overline{P}_{i}\in\mathcal{CL}$, we must have $x_{1}\in\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)$. Consequently, $\overline{X}\cap a\left(v_{1}\right)=\left\{x_{1}\right\}$.
Inductive step.
Let $k\in\left\{1,\dots,m-1\right\}$ be such that for each $\ell\in\left\{1,\dots,k\right\}$, $\overline{X}\cap a\left(v_{\ell}\right)=\left\{x_{1},\dots,x_{\ell}\right\}$. We show that $\overline{X}\cap a\left(v_{k+1}\right)=\left\{x_{1},\dots,x_{k+1}\right\}$.
Consider any object $x\in\left(a\left(v_{k+1}\right)\backslash\left\{x_{k+1}\right\}\right)\backslash a\left(v_{k}\right)$. That is, $x$ labels a vertex that appears after $v_{k}$ and before $v_{k+1}$ on $\tau_{i}\left(X\right)$, and $x\notin\varphi_{i}^{\text{ATTC}}\left(P\right)$. This means that, although agent $i$ may point to $x$ during $\text{ATTC}\left(P\right)$, agent $i$ is not on the cycle of $\text{ATTC}\left(P\right)$ that contains object $x$. It follows that agent $i$ does not belong to the cycle of $\text{ATTC}\left(\overline{P}\right)$ that contains object $x$ either. Hence, $x\notin\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)$.
Now consider object $x_{k+1}$. Because $\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)R_{i}\varphi_{i}^{\text{ATTC}}\left(P\right)$, we must have $x_{k+1}\in\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)$. Consequently, $\overline{X}\cap a\left(v_{k+1}\right)=\left\{x_{1},\dots,x_{k+1}\right\}$.
By the principle of induction, it then follows that $\overline{X}\cap a\left(v_{m}\right)=\left\{x_{1},\dots,x_{m}\right\}$. Because $\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)$ and $\varphi_{i}^{\text{ATTC}}\left(P\right)$ contain the same number of objects, we must have $\varphi_{i}^{\text{ATTC}}\left(\overline{P}\right)=\varphi_{i}^{\text{ATTC}}\left(P\right)$, as we needed to show. $\blacksquare$
A.6 Proof of Theorem 5
It is known that TTC satisfies all properties. Therefore, it suffices to prove the uniqueness. Toward contradiction, suppose that $\varphi$ satisfies the properties but $\varphi\neq\varphi^{\text{TTC}}$.
The notions of size $s$ and similarity $\rho$ are defined exactly as in the proof of Theorem 1.
Recall that, for all $P\in\mathcal{P}^{N}$, (i) $\rho(P)\leq|O|+1$, and (ii) $\rho(P)=|O|+1$ if and only if $\varphi(P)=\varphi^{\text{TTC}}(P)$.
Suppose that $\min_{P\in\mathcal{P}^{N}}\rho\left(P\right)=t$. Then $\varphi\neq\varphi^{\text{TTC}}(P)$ implies that $t\leq|O|$. Among all profiles in $\{P\in\mathcal{P}^{N}\mid\rho(P)=t\}$, let $P$ be one whose size is smallest. Hence, for any profile $\tilde{P}\in\mathcal{P}^{N}$, we have
  • (1)
    $\rho(P)<\rho(\tilde{P})$, or
  • (2)
    $\rho(P)=\rho(\tilde{P})$ and $s(P)\leq s(\tilde{P})$.
Since $\rho(P)=t$, $\varphi(P)$ executes the cycles $C_{1}(P),C_{2}(P),\dots,C_{t-1}(P)$ but not $C_{t}(P)$. Let $\mu\coloneqq\varphi^{\text{TTC}}(P)$, $C\coloneqq C_{t}(P)$, and $O^{t}\coloneqq O\backslash\bigcup_{\tau=1}^{t-1}O\left(C_{\tau}\left(P\right)\right)$. Thus, there is an agent $i\in N(C)$ such that, although agent $i$ points to $\mu_{i}$ on $C$, she does not receive $\mu_{i}$ at $\varphi(P)$, i.e., $\mu_{i}\neq\varphi_{i}(P)$. Note that, by the definition of $\text{TTC}(P)$, $\mu_{i}$ is agent $i$’s most-preferred object in $O^{t}$ at $P_{i}$, i.e., $\mu_{i}=\max_{P_{i}}\left(O^{t}\right)$. Thus,
$\mu_{i}\mathbin{P_{i}}\varphi_{i}(P).$ (8)
If $|N(C)|=1$, then $C=(i,o_{i},i)$ and $\mu_{i}=o_{i}\mathbin{P_{i}}\varphi_{i}(P)$, which violates individual rationality. Thus, $|N(C)|\geq 2$.
Next, we show that $\varphi_{i}(P)=o_{i}$. Toward contradiction, suppose that $\varphi_{i}(P)\neq o_{i}$. By (8) and individual rationality, $\mu_{i}\mathbin{P_{i}}\varphi_{i}(P)\mathbin{P_{i}}o_{i}$.
Let $P^{\prime}_{i}$ be the truncation of $P_{i}$ at $\mu_{i}$, i.e., $P^{\prime}_{i}:\dots,\mu_{i},o_{i},\dots$. [^29] A similar argument using drop strategy-proofness applies if $P^{\prime}_{i}$ is obtained from $P_{i}$ by dropping object $\varphi_{i}(P)$. Let $P^{\prime}\coloneqq(P^{\prime}_{i},P_{-i})$. Then $s_{i}(P^{\prime}_{i})<s_{i}(P_{i})$ and $s(P^{\prime})<s(P)$.
By the definition of TTC, for each step $\tau\in\{1,\dots,t\}$, the procedures $\text{TTC}(P)$ and $\text{TTC}(P^{\prime})$ generate and execute the same cycles, i.e., for each $\tau$, $C_{\tau}(P)=C_{\tau}(P^{\prime})$. Moreover, the choice of $P$ implies that $\rho(P^{\prime})\geq\rho(P)=t$. Thus, $\varphi(P^{\prime})$ executes all cycles in $\{C_{1}(P),\dots,C_{t-1}(P)\}$. Thus, $\varphi_{i}(P^{\prime})\in O^{t}$. Moreover, since $\varphi$ is truncation-proof, we know that $i$ cannot be better of by misreporting $P^{\prime}_{i}$ at $P$. Together with (8), we conclude that $\mu_{i}\neq\varphi_{i}(P^{\prime})$. Consequently, $\varphi(P^{\prime})$ does not execute $C_{t}(P^{\prime})$, which means that $\rho(P^{\prime})=t$. But then we have $\rho(P^{\prime})=\rho(P)$ and $s(P^{\prime})<s(P)$, which contradicts the choice of $P$! Thus, we conclude that $\varphi_{i}(P)=o_{i}$.
Let $j$ be the agent who points to $o_{i}$ on $C$, i.e., $j\in N(C)$ is such that $\mu_{j}=o_{i}$. Then $\varphi_{i}(P)=o_{i}=\mu_{j}$ implies that $\varphi_{j}(P)\neq\mu_{j}$. Thus, $\mu_{j}\mathbin{P_{j}}\varphi_{j}(P).$ A similar argument shows that $\varphi_{j}(P)=o_{j}$.
By repeating the same argument for each agent in $N(C)$, we can show that each agent in $N(C)$ is assigned her endowment at $\varphi(P)$, i.e., for each $i^{\prime}\in N(C)$, $\varphi_{i^{\prime}}(P)=o_{i^{\prime}}$. But then $\varphi$ is not Pareto efficient, because the agents in $N(C)$ can benefit by executing $C$. This contradiction completes the proof of Theorem 5.
A.7 Proof of Theorem 6
The proof here is almost identical to the proof of Theorem 1. The main difference is that, since we focus on conditionally lexicographic preferences, any statements that pertained to an agent’s most-preferred object in Theorem 1 now pertain to an agent’s most-preferred object conditional on receiving some set of objects.
Toward contradiction, suppose that $\varphi$ satisfies the properties but $\varphi\neq\varphi^{ATTC}$.
For each $P\in\mathcal{P}^{N}$, and each $t\in\mathbb{N}$, let $\mathcal{C}_{t}(P)$ be the set of cycles that obtain at step $t$ of $\text{ATTC}\left(P\right)$. Similarly, let $C_{t}(P)\in\mathcal{C}_{t}(P)$ be the cycle that is executed at step $t$ of $\text{ATTC}\left(P\right)$.
Denote the size of a profile $P$ by $\sum_{i\in N}\sum_{X\in 2^{O}}\left|a\left(w_{P_{i}}\left(\omega_{i}\mid X\right)\right)\right|$. The similarity $\rho$ is defined exactly as in the proof of Theorem 1.
Let $P$ be a profile such that for any other profile $\tilde{P}\neq P$, either
  • (1)
    $\rho(P)<\rho(\tilde{P})$, or
  • (2)
    $\rho(P)=\rho(\tilde{P})$ and $s(P)\leq s(\tilde{P})$.
Let $t\coloneqq\rho(P)$. Since $\rho(P)=t$, $\varphi(P)$ executes the cycles $C_{1}(P),C_{2}(P),\dots,C_{t-1}(P)$ but not $C_{t}(P)$. Let $C\coloneqq C_{t}(P)$ and $O^{t}\coloneqq O\backslash\bigcup_{\tau=1}^{t-1}O\left(C_{\tau}\left(P\right)\right)$. Suppose that
$C=\left(i_{1},o_{i_{2}},i_{2},o_{i_{3}},\dots,i_{k-1},o_{i_{k}},i_{k},o_{i_{k+1}},i_{k+1}=i_{1}\right).$
Because $\varphi(P)$ does not execute $C$, there is an agent $i_{\ell}\in N(C)$ such that, although agent $i_{\ell}$ points to $o_{i_{\ell+1}}$ on $C$, she does not receive $o_{i_{\ell+1}}$ at $\varphi(P)$, i.e., $o_{i_{\ell+1}}\notin\varphi_{i}(P)$. Note that, by the definition of $\text{ATTC}(P)$, $o_{i_{\ell+1}}$ is agent $i_{\ell}$’s most-preferred object in $O^{t}$ at $P_{i_{\ell}}$ conditional on receiving $\varphi_{i_{\ell}}(P)\setminus O^{t}$, i.e., $o_{i_{\ell+1}}=\max_{P_{i_{\ell}}}(O^{t}\mid\varphi_{i_{\ell}}(P)\setminus O^{t})$. Without loss of generality, let $i_{\ell}=i_{k}$. Thus, $o_{i_{1}}\notin\varphi_{i_{k}}(P)$.
Let $X_{i_{k}}$ be the set of all objects in $O\backslash O^{t}$ that are assigned to agent $i_{k}$, i.e., $X_{i_{k}}=\varphi_{i_{k}}(P)\backslash O^{t}$. We next show that, apart from the objects in $X_{i}$, agent $i_{k}$ only receives the remainder of her endowment.
Claim 4.
$\varphi_{i_{k}}(P)\cap O^{t}=\omega_{i_{k}}\cap O^{t}$.
Proof of Claim 4.
Suppose otherwise. By balancedness and the fact that $\left|\varphi_{i_{k}}\left(P\right)\backslash O^{t}\right|=\left|\omega_{i_{k}}\backslash O^{t}\right|$, we must have $\left|\varphi_{i_{k}}\left(P\right)\cap O^{t}\right|=\left|\omega_{i_{k}}\cap O^{t}\right|$. Consequently, $\varphi_{i_{k}}\left(P\right)\cap O^{t}\neq\omega_{i_{k}}\cap O^{t}$ implies there is an object $o^{\prime}\in\varphi_{i_{k}}(P)\cap O^{t}$ with $o^{\prime}\not\in\omega_{i_{k}}$. By the worst endowment lower bound,
$o^{\prime}\in a\left(w_{P_{i_{k}}}(\omega_{i_{k}}\mid\varphi_{i_{k}}(P))\right).$
By the definition of $\text{ATTC}(P)$, we know that $o_{i_{1}}$ is agent $i_{k}$’s most-preferred object in $O^{t}$ conditional on receiving $X_{i_{k}}$, i.e., $o_{i_{1}}=\max_{P_{i_{k}}}(O^{t}\mid X_{i_{k}})$. Because $o^{\prime}\in\varphi_{i_{k}}(P)\cap O^{t}$ and $o_{i_{1}}\notin\varphi_{i_{k}}(P)$, we have $o_{i_{1}}\neq o^{\prime}$.
Let $P^{\prime}_{i_{k}}$ be the drop strategy for $P_{i_{k}}$ obtained by dropping object $x$. Let $P^{\prime}\coloneqq(P^{\prime}_{i_{k}},P_{-i_{k}})$. By the definition of TTC, we see that until step $t$, all top trading cycles that are executed via TTC at $P$ and $P^{\prime}$ are the same, i.e., $C_{\tau}(P)=C_{\tau}(P^{\prime})$ for $\tau=1,\ldots,t$. By the selection of $P$, we know that $\rho(P)=t\leq\rho(P^{\prime})$. Thus, all cycles in $\{C_{1}(P),\dots,C_{t-1}(P)\}$ are executed at $\varphi(P^{\prime})$. Therefore,
$X_{i_{k}}\subseteq\varphi_{i_{k}}(P^{\prime}).$
Since $\varphi$ is weakly drop strategy-proof, we know that $i_{k}$ cannot be better off by misreporting $P^{\prime}_{i_{k}}$ at $P$. Together with $X_{i_{k}}\subseteq\varphi_{i_{k}}(P^{\prime})$ and $o_{i_{k}}\notin\varphi_{i_{k}}(P)$, we conclude that $i_{k}$ cannot receive $o_{i_{1}}$ at $\varphi(P^{\prime})$, i.e., $o_{i_{1}}\notin\varphi_{i_{k}}(P^{\prime})$. Therefore, $C$ is not executed at $\varphi(P^{\prime})$, which means that $\rho(P^{\prime})\leq t$. As a result, we find that $\rho(P^{\prime})=t$. However, since $P^{\prime}_{i_{k}}$ is obtained from $P_{i_{k}}$ by dropping an object $o^{\prime}\in a\left(w_{P_{i_{k}}}(\omega_{i_{k}}\mid\varphi_{i_{k}}(P))\right)$, we see that $s(P^{\prime})<s(P)$. This contradicts the choice of $P$. ∎
Next, we show that at least two agents are involved in $C$.
Claim 5.
$|N(C)|>1$.
Proof.
Toward contradiction, suppose that $|N(C)|=1$, that is to say, $i_{1}=i_{k}$ and $C=(i_{1},o_{i_{1}},i_{1})$. Then $o_{i_{1}}\in\omega_{i_{k}}$ and Claim 4 implies that $o_{i_{1}}\in\varphi_{i_{k}}(P)$, a contradiction. ∎
By Claim 4 and Claim 5, agent $i_{k-1}$ points to $o_{i_{k}}$ on $C$, yet she does not receive $o_{i_{k}}$. An argument similar to the proof of Claim 4 implies that $\varphi_{i_{k-1}}(P^{\prime})\cap O^{t}=\omega_{i_{k-1}}\cap O^{t}$.
Proceeding by induction, we conclude that for each agent $i_{\ell}$ that is involved in $C$, we have $\varphi_{i_{\ell}}(P)\cap O^{t}=\omega_{i_{\ell}}\cap O^{t}$.
However, this means that $\varphi$ is not ig-efficient, as agents in $N(C)$ can benefit by execution of $C$. This contradiction completes the proof of Theorem 6. $\blacksquare$
A.8 Proof of Proposition 7
Consider an agent with monotone preferences that are not conditionally lexicographic. Without loss of generality, let this agent be agent $1$. That is, $P_{1}\in\mathcal{M}\setminus\mathcal{CL}$. By Proposition 4, there exist disjoint subsets $X,Y\in 2^{O}$ with $X\neq\emptyset$ such that
for all $\text{for all }x\in X\text{, there exists }Z_{x}\subseteq X\setminus x\text{ such that }(Y\cup Z_{x})\mathbin{P_{1}}(Y\cup x).$
Because $P_{1}$ is monotone, $[Y\cup(X\setminus x)]\mathbin{P_{1}}(Y\cup Z_{x})$ as $Z_{x}\subseteq X\setminus x$. Thus, we have
for all $\text{for all }x\in X,\quad[Y\cup(X\setminus x)]\mathbin{P_{1}}(Y\cup x).$
Note also that $|X|\geq 3.$ [^30] If $X$ is a singleton, say $X=\{x\}$, then $Y\mathbin{P_{1}}(Y\cup x)$, a violation of monotonicity. If $X$ contains two objects, say $X=\{x,y\}$, we would have $(Y\cup x)\mathbin{P_{1}}(Y\cup y)\mathbin{P_{1}}(Y\cup x)$, a violation of transitivity.
Let $x^{*}\in X$ be such that
for all $\text{for all }x\in X,\quad(Y\cup x^{*})\mathbin{R_{1}}(Y\cup x).$
Then it follows that
for all $\text{for all }x\in X,\quad[Y\cup(X\setminus x)]\mathbin{P_{1}}(Y\cup x^{*})\mathbin{R_{1}}(Y\cup x).$
Case 1:
Suppose $X\cup Y=O$. Let $N=\{1,2\}$, and let $P_{2}\in\mathcal{L}$ be such that (i) $\max_{P_{2}}(O)=x^{*}$, and (ii) for all $x\in X$, $x\mathbin{P_{2}}Y$. Then, in particular,
$x^{*}\mathbin{P_{2}}[Y\cup(X\setminus x^{*})]\mathbin{R_{2}}(X\setminus x^{*}).$
Consider the allocations
$\mu\coloneqq(Y\cup x^{*},X\backslash x^{*})\quad\text{and}\quad\overline{\mu}\coloneqq(Y\cup\left(X\backslash x^{*}\right),x^{*}).$
Then $\mu$ is not Pareto efficient at $P$ because it is Pareto-dominated by $\overline{\mu}$ at $P$.
We claim that $\mu$ is ig-efficient at $P$. Indeed, consider any single-object exchange identified by the cycle $C=(1,x,2,y,1)$, where $x\in\mu_{2}$ and $y\in\mu_{1}$. Executing the exchange results in the allocation $\mu^{\prime}$ where
$\mu^{\prime}_{1}=\left(Y\cup\left\{x^{*},x\right\}\right)\backslash y\quad\text{and}\quad\mu^{\prime}_{2}=\left(X\backslash\left\{x^{*},x\right\}\right)\cup y.$
If $y\in Y$, then this exchange harms agent 2, i.e., $\mu_{2}\mathbin{P_{2}}\mu^{\prime}_{2}$. On the other hand, if $y=x^{*}$, then this exchange harms agent 1 because $\mu_{1}=(Y\cup x^{*})\mathbin{P_{1}}(Y\cup x)=\mu^{\prime}_{1}$. Hence, $\mu$ is ig-efficient.
Case 2:
Suppose $X\cup Y\subsetneq O$. Let $N=\{1,2,3\}$ and denote $\overline{O}\coloneqq O\setminus(X\cup Y)$. Let $P_{2}\in\mathcal{L}$ be such that (i) $\max_{P_{2}}(O)=x^{*}$, and (ii) for all $x\in X$, $x\mathbin{P_{2}}(O\setminus X)$. Let $P_{3}\in\mathcal{L}$ be such that (i) for all $z\in\overline{O}$, $z\mathbin{P_{3}}(X\cup Y)$.
Consider the allocations
$\mu\coloneqq(Y\cup x^{*},X\backslash x^{*},\overline{O})\quad\text{and}\quad\overline{\mu}\coloneqq(Y\cup\left(X\backslash x^{*}\right),x^{*},\overline{O}).$
Then $\mu$ is not Pareto efficient at $P$ because it is Pareto-dominated by $\overline{\mu}$ at $P$.
We claim that $\mu$ is ig-efficient at $P$. Because $\mu_{3}$ is agent 3’s top-ranked $|\overline{O}|$-subset of $O$, any single-object exchange involving agent 3 must harm agent 3. Thus, any Pareto-improving single-object exchange must involve only agents 1 and 2. However, there is no such exchange by the argument in Case 1. Hence, $\mu$ is ig-efficient. $\blacksquare$
Appendix B Examples
B.1 Independence of Properties in Theorems 1 and  6
We establish the independence of the properties in Theorem 1 by providing examples of rules, different from TTC / ATTC, that violate exactly one of the properties. For each of the examples below we indicate the property that the rule fails (while it satisfies all remaining properties).
Individual-good efficiency: the no-trade rule that always selects the initial allocation satisfies all properties except for individual-good efficiency.
The worst endowment lower bound: the serial dictatorships subject to balancedness [^31] That is, each dictator is assigned the same number of objects as her endowment. satisfy all properties except for the worst endowment lower bound.
Balancedness: the serial dictatorships subject to the worst endowment lower bound [^32] That is, each dictator $i$ is assigned her best bundle among all subsets $X$ of the remaining objects that would not violate the worst endowment lower bound (i.e., such that $X\subseteq\{o\in O\mid o\mathbin{R_{i}}\min_{P_{i}}(\omega_{i})\}$, where $P_{i}$ denotes $i$’s preference relation.) satisfy all properties except for balancedness.
Truncation-proofness / Drop strategy-proofness: the following example is from Altuntaş et al. (2023). Let $N=\{1,2,3,4\}$, and for each $i\in N$, $\omega_{i}=o_{i}$. Let $P^{*}$ be defined as follows (endowment is underlined).
$P^{*}_{1}:o_{2},o_{4},\underline{o_{1}},o_{3};$
$P^{*}_{2}:o_{1},o_{3},\underline{o_{2}},o_{4};$
$P^{*}_{3}:o_{2},o_{4},\underline{o_{3}},o_{1};$
$P^{*}_{4}:o_{1},o_{3},\underline{o_{4}},o_{2}.$
Let $\varphi$ be defined as follows.
$\varphi\left(P\right)=\begin{cases}(o_{4},o_{3},o_{2},o_{1}),&\text{if }P=P^{*}\\ \varphi^{\text{ATTC}}(P),&\text{otherwise.}\end{cases}$
Note that for lexicographic preferences, $\varphi^{\text{ATTC}}=\varphi^{\text{TTC}}$. Altuntaş et al. (2023) show that $\varphi$ satisfies Pareto efficiency and the strong endowment lower bound (hence balancedness and the worst endowment lower bound), but violates drop strategy-proofness / truncation-proofness. To see this, note that agent $1$ receives $o_{4}$ at $\varphi\left(P^{*}\right)$, and she can receive her best object $o_{2}$ by misreporting a drop strategy $P^{\prime}_{1}:o_{2},\underline{o_{1}},o_{3},o_{4}$ or a truncation strategy $P^{\prime\prime}_{1}:o_{2},\underline{o_{1}},o_{4},o_{3}$.
B.2 Other examples
Example 7 (A rule on $\mathcal{L}$ that satisfies truncation-proofness, the worst endowment lower bound and individual rationality, but not drop strategy-proofness.).
Suppose $N=\left\{1,2,3\right\}$, and $\omega=\left(o_{1},o_{2},o_{3}\right)$. Let $P^{*}\in\mathcal{P}$ be such that
$P_{1}^{*}:o_{2},o_{3},o_{1},\quad P_{2}^{*}:o_{3},o_{1},o_{2},\quad P_{3}^{*}:o_{1},o_{2},o_{3}.$
Note that each agent $i\in N$ only has two truncation strategies, i.e.,
[…]
Consider the rule $\varphi$ defined for all $P\in\mathcal{P}$ by
$\varphi\left(P\right)=\begin{cases}\omega,&\text{if }P\in\mathcal{T}_{1}\left(P_{1}^{*}\right)\times\mathcal{T}_{2}\left(P_{2}^{*}\right)\times\mathcal{T}_{3}\left(P_{3}^{*}\right)\\ \varphi^{\text{TTC}}\left(P\right),&\text{otherwise}.\end{cases}$
That is, $\varphi\left(P\right)=\omega$ whenever $P$ is such that, for all $i\in N$, $P_{i}$ is a truncation of $P_{i}^{*}$; otherwise, $\varphi\left(P\right)=\varphi^{\text{TTC}}\left(P\right)$. By the definition of $\varphi$, it is easy to see that $\varphi$ is individually rational. It is straightforward to show that $\varphi$ is truncation-proof. To see that $\varphi$ is not drop strategy-proof, consider
$P^{\prime}_{1}:o_{3},o_{1},o_{2}.$
Then $P^{\prime}_{1}$ is obtained from $P_{1}^{*}$ by dropping object $o_{2}\in O\backslash\omega_{1}$. Because
$\varphi_{1}\left(P^{\prime}_{1},P_{-1}^{*}\right)=\varphi_{1}^{\text{TTC}}\left(P^{\prime}_{1},P_{-1}^{*}\right)=o_{3}\mathbin{P_{1}^{*}}o_{1}=\varphi_{1}\left(P^{*}\right),$
agent 1 can manipulate at $P^{*}$ by misreporting $P^{\prime}_{1}\in\mathcal{D}_{1}\left(P_{1}^{*}\right)$. Note that $P^{\prime}_{1}$ is not a truncation of $P_{1}^{*}$.  $\diamond$
Example 8 (Drop strategy-proofness does not imply truncation-proofness.).
Suppose $N=\left\{1,2,3\right\}$ and $\omega=\left(o_{1},o_{2},o_{3}\right)$. Let $\mathcal{P}$ be any domain of preferences containing $\mathcal{L}$. Let $\varphi$ be a rule on $\mathcal{P}$ such that, for all $P\in\mathcal{P}$,
$\varphi\left(P\right)$ =$=\begin{cases}\left(o_{1},o_{2},o_{3}\right),&\text{if }o_{1}\text{ is agent 1's second-ranked object at }P_{1}\\ \left(o_{2},o_{1},o_{3}\right),&\text{if }o_{2}\text{ is agent 1's second-ranked object at }P_{1}\\ \left(o_{3},o_{2},o_{1}\right),&\text{if }o_{3}\text{ is agent 1's second-ranked object at }P_{1}.\end{cases}$
Then $\varphi$ is drop strategy-proof. Indeed, if $P^{\prime}_{1}$ $(\neq P_{1})$ is a drop strategy obtained from $P_{1}$ by dropping any object, then agent 1 receives her least-preferred object at $\left(P^{\prime}_{1},P_{-1}\right)$, i.e., $\varphi_{1}\left(P^{\prime}_{1},P_{-1}\right)=\min_{P_{1}}\left(O\right)$.
Consider $P_{1}\in\mathcal{L}$ such that $\succ^{P_{1}}:o_{2},o_{3},o_{1}.$ Note that $\varphi_{1}(P_{1},P_{-1})=o_{3}$ for each profile $P_{-1}$ of the other agents. Next, consider a truncation strategy obtained from $P_{1}$ by dropping $\{o_{2},o_{3}\}$, i.e., such that $\succ^{P^{\prime}_{1}}:o_{1},o_{2},o_{3}.$ We see that $\varphi_{1}\left(P^{\prime}_{1},P_{-1}\right)=o_{2}$. Since $\varphi_{1}\left(P^{\prime}_{1},P_{-1}\right)=o_{2}\mathbin{P_{1}}o_{3}=\varphi_{1}\left(P\right)$, we find that $\varphi$ is not truncation-proof.  $\diamond$